### 04 Apr 2026 v1

# Magnitude homology of real hyperplane arrangements

Abstract.
We initiate the study of magnitude theory of real hyperplane arrangements. Magnitude is a cardinality-like invariant of metric spaces or enriched categories measuring the effective size. Its categorification, the magnitude homology, is a more powerful invariant. For a real hyperplane arrangement, or more generally, an oriented matroid, the tope graph encapsulates considerable amount of information. Since tope graphs are equipped with the shortest path metric, we feed them to the magnitude and magnitude homology machinery to derive new invariants of real hyperplane arrangements. We prove some structural results of the magnitude of arrangements, including reciprocity, palindromic numerator and denominator. For magnitude homology of arrangements, we give combinatorial descriptions in small length and prove that tope graphs are diagonal if and only if the arrangement is Boolean. We present a face decomposition of magnitude homology, using which we obtain a combinatorial formula of diagonal magnitude Betti numbers. Many open problems are posted for future study. In particular, we conjecture that magnitude and magnitude homology of arrangements are determined by the intersection lattice.
Key words and phrases:
hyperplane arrangement, tope graph, magnitude, magnitude homology
2020 Mathematics Subject Classification:
Primary 52C35; Secondary 55N35
1. Introduction
Let $\mathcal{A}$ be an arrangement of hyperplanes in $\mathbb{R}^{\ell}$ all passing through the origin. The hyperplanes in $\mathcal{A}$ cut $\mathbb{R}^{\ell}$ into a disjoint union of chambers. Let $\mathrm{Ch}(\mathcal{A})$ denote the set of chambers. For two chambers $C,C^{\prime}\in\mathrm{Ch}(\mathcal{A})$, we write $S(C,C^{\prime})$ as the set of hyperplanes in $\mathcal{A}$ that separate $C$ and $C^{\prime}$.
The tope graph $\mathcal{T}(\mathcal{A})$ of $\mathcal{A}$ is the graph with vertex set $\mathrm{Ch}(\mathcal{A})$, and two vertices $C$ and $C^{\prime}$ are connected by an edge if $S(C,C^{\prime})$ consists of a single hyperplane. The graph metric $d$ on $\mathcal{T}(\mathcal{A})$ satisfies $d(C,C^{\prime})=\#S(C,C^{\prime})$. Note that $\mathcal{T}(\mathcal{A})$ is a partial cube, that is, isometric to a subgraph of the hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}$ equipped with the Hamming distance. For an oriented matroid $\mathcal{M}$, the tope graph $\mathcal{T}(\mathcal{M})$ can be defined similarly.
Although seemingly elementary, the tope graphs turn out to be quite informative.
Theorem 1.1 (Theorem 4.2.14 of [4]).
A simple oriented matroid is uniquely determined (upto reorientation) by its unlabeled tope graph.
We refer to Figure 1 of [32] for a detailed diagram of implications. In particular, if two arrangements have isomorphic tope graphs, then they have homoemorphic complexified complements, isomorphic intersection lattices and isomorphic filtered Varchenko–Gelfand algebras. Hence we may study $\mathcal{A}$ via $\mathcal{T}(\mathcal{A})$.
In this paper, we feed tope graphs into the emerging machinery of magnitude theory. Originally introduced by Leinster for enriched categories and metric spaces [24, 21], magnitude acts as a cardinality-like invariant measuring the “effective size” of a space. For finite graphs [25], it is a rational function defined via the inverse of the $q$-distance matrix, which is intimately related to the classical Varchenko matrix in the context of arrangements. The categorification of this concept, magnitude homology, developed by Hepworth and Willerton for graphs [16], and by Leinster and Shulman for enriched categories and metric spaces [23], provides a powerful bigraded homology theory that detects profound metric features.
We initiate the study of magnitude and magnitude homology specifically tailored to the metric geometry of real hyperplane arrangements. Our motivation is two-fold. The first is that we introduce a new rational function invariant and a new bigraded homology invariant of real arrangements. As a special class of graphs, tope graphs of arrangements (or oriented matroids) possess many unique features, such as antipodal symmetry and organized face decomposition, that enable us to contribute to the theory of magnitude with new beautiful examples.
At the decategorified level, we demonstrate that the magnitude of an arrangement, $\mathrm{Mag}(\mathcal{A},q)$, exhibits remarkable structural results. We prove a reciprocity formula relating $\mathrm{Mag}(\mathcal{A},q)$ and $\mathrm{Mag}(\mathcal{A},q^{-1})$, and establish that its reduced rational form consists of palindromic numerator and denominator whose poles are roots of unity. Supported by extensive computational evidence generated via SageMath on the Grünbaum catalogue [14], we further propose a conjectural Face Decomposition formula, which serves as a discrete $q$-analogue to the classical decomposition of the Euler characteristic of a zonotope.
At the homological level, by exploiting the metric structure of gated subgraphs within partial cubes, which naturally arise from the faces of the arrangement, we systematically dissect the magnitude homology of $\mathcal{T}(\mathcal{A})$. After providing explicit combinatorial descriptions of the homology groups for small lengths, we establish a strict topological dichotomy for arrangements. We prove that the tope graph is diagonal, meaning its off-diagonal magnitude homology strictly vanishes, if and only if the arrangement is Boolean. We also prove a face decomposition of magnitude homology of arrangements, and as an application, we derive a formula of diagonal magnitude Betti numbers in terms of the intersection lattice.
The paper is organized as follows. Section 2 collects necessary recollections from metric graph theory and the theory of hyperplane arrangements, establishing the foundational link between gated subgraphs and covector composition. Section 3 investigates the basic properties and structural rigidity of the magnitude rational function. Section 4 explores magnitude homology, presenting the Boolean diagonality theorem and face decomposition theorem. Finally, Section 5 posts several open problems and conjectures for future study. Results of many computer experiments with SageMath form the appendix.
Convention: When we talk about a topological property/construction of a poset, we mean that of its order complex.
2. Recollections
2.1. Metric graph theory
In this paper, all graphs we mention are supposed to be finite simple undirected graphs unless otherwise specified. For a graph $G=(V,E)$, a walk (or gallery) of length $k$ from vertex $x\in V$ to vertex $y\in V$ is a sequence of vertices $(x_{0},x_{1},\ldots,x_{k})$ such that $x_{0}=x$, $x_{k}=y$ and $\{x_{i},x_{i+1}\}\in E$ for $i=0,1,\ldots,k-1$. We define a distance function $d:V\times V\to\mathbb{Z}\cup\{\infty\}$ by declaring $d(x,y)$ is the minimal length of walks from $x$ to $y$. If there is no walk from $x$ to $y$, we set $d(x,y)=\infty$. Then $(G,d)$ is a generalized metric space. In particular, when $G$ is connected, that is there is a walk between any pair of vertices, $(G,d)$ is a metric space. We call the maximal value of $d$ the diameter of $G$, denoted by $\mathrm{diam}(G)$.
A subgraph $H$ of $G$ can be equipped with a metric in two ways. One is to restrict $d_{G}$ to $H$ and the other is $d_{H}$ on $H$ as a graph itself. We say $H$ is an isometric subgraph of $G$ if the two metrics coincide $d_{H}(x,y)=d_{G}(x,y)$ for $x,y\in V(H)$.
In a graph $G$, a walk from $x$ to $y$ of minimal length $d(x,y)$ is called a minimal walk. Given $x,y\in V$, let us define the closed interval $x,y]_{G}$ as the set of all vertices traversed by a minimal walk from $x$ to $y$. It is a poset equipped with partial order $u\prec_{G}v$ if there is a minimal walk from $x$ to $y$ traversing $u$ before traversing $v$. Note that $x,y]_{G}$ has a unique minimal element $x$ and a unique maximal element $y$. A vertex $u$ belongs to $x,y]_{G}$ if and only if $d(x,y)=d(x,u)+d(u,y)$. More generally, vertices $u_{1},\ldots,u_{k}$ form a strict linearly ordered chain in $x,y]_{G}$, that is $x\prec_{G}u_{1}\prec_{G}\cdots\prec_{G}u_{k}\prec_{G}y$ if and only if
$d(x,y)=d(x,u_{1})+d(u_{1},u_{2})+\cdots+d(u_{k-1},u_{k})+d(u_{k},y).$
We define the open interval $(x,y)_{G}$ as the subposet of $x,y]_{G}$ excluding $x$ and $y$.
We say a subgraph $H$ of $G$ is convex in $G$ if for any $x,y\in V(H)$, any minimal walk from $x$ to $y$ traverses only vertices in $H$, equivalently $x,y]_{G}\subseteq V(H)$ for $x,y\in V(H)$. Apparently, any convex subgraph is an isometric subgraph, but the converse is false.
Definition 2.1.
Let $G$ be a connected graph, $H$ an isometric subgraph of $G$ and $x\in V(G)$. We say $H$ is gated with respect to $x$ if there is $u\in V(H)$ such that for any $v\in V(H)$,
$d(x,v)=d(x,u)+d(u,v).$
Such $u$ must be unique since the above equation says $u$ is the closest vertex to $x$ in $H$. We may call $u$ the gate of $H$ with respect to $x$. Intuitively, any minimal walk from $x$ to a vertex of $H$ enters $H$ at the gate. We say $H$ is a gated subgraph of $G$ is it is gated with respect to all vertices of $G$. In this case, we call the map $p:G\to H$ taking any vertex of $G$ to its gate in $H$ the gate projection. Note that the restriction of $p$ to $H$ is the identity map $\text{id}_{H}$.
Proposition 2.2 (Proposition A.1 of [2]).
A gated subgraph is necessarily convex.
Proof.
Suppose $H$ is gated in $G$. Choose any minimal walk in $G$ from $x$ to $y$ for $x,y\in V(H)$ and let $z$ be an arbitrary vertex in this walk, we need to show $z\in V(H)$. Since $H$ is gated, we have
d(z,x)=d(z,p(z))+d(p(z),x),d(z,y)=d(z,p(z))+d(p(z),y).d(z,x)=d(z,p(z))+d(p(z),x),~d(z,y)=d(z,p(z))+d(p(z),y).
Since $z$ belongs to $x,y]_{G}$, adding these two equations yields
$d(x,y)=d(x,z)+d(z,y)=d(x,p(z))+d(p(z),y)+2d(z,p(z))\geq d(x,y)+2d(z,p(z)).$
This forces $d(z,p(z))=0$, i.e. $z=p(z)\in V(H)$. ∎
In the scope of this paper, the homomorphisms between graphs we shall use are distance non-increasing maps (1-Lipschitz maps) $f:G\to H$, which are maps $f:V(G)\to V(H)$ such that $d_{H}(f(x),f(y))\leq d_{G}(x,y)$ for $x,y\in V(G)$. Note that gate projections are distance non-increasing.
2.2. Graph magnitude
See Leinster [25] for details. For a finite simple graph $G=(V,E)$, define the Zeta matrix $Z_{G}=Z_{G}(q)=[q^{d(x,y)}]_{x,y\in V}$, where $q^{\infty}=0$ by convention. Note that $Z_{G}(0)$ is the identity matrix and hence $\det Z_{G}(q)\in\mathbb{Z}[q]$ is a polynomial with constant term $1$. The rational function $1/\det Z_{G}(q)\in\mathbb{Q}(q)$ then can be expressed as a power series in $\mathbb{Z}[\![q]\!]$. In particular, $Z_{G}(q)$ is invertible over $\mathbb{Q}(q)$, with inverse $Z_{G}(q)^{-1}=\frac{1}{\det Z_{G}(q)}\mathrm{adj}~Z_{G}(q)$, whose entries can be expressed as power series in $\mathbb{Z}[\![q]\!]$.
The magnitude of $G$ is defined by the sum of all entries of $Z_{G}(q)^{-1}$.
$\mathrm{Mag}(G)=\mathrm{Mag}(G,q):=\sum_{x,y\in V}Z_{G}(q)^{-1}_{x,y}\in\mathbb{Q}(q)~(\mathrm{or}~\mathbb{Z}[\![q]\!]).$
In general, the sum of all entries of a matrix $A$ is equal to $\mathbf{1}^{T}A\mathbf{1}$, where $\mathbf{1}$ is the column vector of appropriate size with all entries equal to $1$. If $B$ is invertible, then the sum of all entries of $B^{-1}$ can be computed from $\mathbf{1}^{T}B^{-1}\mathbf{1}=\mathbf{1}^{T}\mathbf{x}$, where $\mathbf{x}$ is the unique solution of $B\mathbf{x}=\mathbf{1}$. For graph $G$, define the weight vector $w_{G}:=Z_{G}^{-1}\mathbf{1}$ as the unique solution to $Z_{G}\mathbf{x}=\mathbf{1}$, in other words, as a function $w_{G}:V\to\mathbb{Q}(q)$, $w_{G}(x)=\sum_{y\in V}Z_{G}(q)^{-1}_{x,y}$ is the $x$-row sum of $Z_{G}^{-1}$. Then the magnitude
$\mathrm{Mag}(G)=\mathbf{1}^{T}w_{G}=\sum_{x\in V}w_{G}(x).$
There is a simple case that we can compute magnitude easily.
Proposition 2.3 (Lemma 3.2 of [25]).
If $G$ is vertex-transitive (i.e. $\mathrm{Aut}(G)$ acts transitively on $V$), then
$\mathrm{Mag}(G)=\frac{\#V}{\sum_{x\in V}q^{d(g,x)}},$
for any $g\in V$.
Proof.
The denominator $s(g)=\sum_{x\in V}q^{d(g,x)}$ is independent of $g\in V$ by the assumption that $G$ is vertex-transitive, and thus denoted by $s=s(g)$. Then it is easily seen that $w_{G}=[1/s\cdots 1/s]^{T}$ is the solution of $Z_{G}\mathbf{x}=\mathbf{1}$. We conclude $\mathrm{Mag}(G)=\#V/s$. ∎
Example 2.4.
Let $K_{n}$ be the complete graph on $n$ vertices and $C_{n}$ be the cycle graph on $n$ vertices. They are vertex-transitive. Then
$\mathrm{Mag}(K_{n})$ $=\frac{n}{1+(n-1)q},$
$\mathrm{Mag}(C_{2n})$ $=\frac{2n(q-1)}{(q^{n}-1)(q+1)}=\frac{2n}{(1+q)(1+q+q^{2}+\cdots+q^{n-1})},$
$\mathrm{Mag}(C_{2n-1})$ $=\frac{(2n-1)(q-1)}{2q^{n}-q-1}=\frac{2n-1}{1+2q+2q^{2}+\cdots+2q^{n-1}}.$
Magnitude behaves as cardinality in the following sense.
Proposition 2.5 (Lemmas 3.5 and 3.6 of [25]).
Let $G,H$ be graphs. Write $G\sqcup H$ and $G\square H$ the disjoint union and cartesian product of $G$ and $H$ respectively. Then
$\mathrm{Mag}(G\sqcup H)=\mathrm{Mag}(G)+\mathrm{Mag}(H),~\mathrm{Mag}(G\square H)=\mathrm{Mag}(G)\mathrm{Mag}(H).$
Under some conditions, magnitude satisfies inclusion-exclusion.
Theorem 2.6 (Theorem 4.9 of [25]).
If a graph $X$ has a decomposition into subgraphs $X=G\cup H$ such that $G\cap H$ is convex in $X$ and gated in $H$, then
$\mathrm{Mag}(X)=\mathrm{Mag}(G)+\mathrm{Mag}(H)-\mathrm{Mag}(G\cap H).$
2.3. Magnitude homology of graphs
For a graph $G$, we call a $(k+1)$-tuple $\vec{x}=(x_{0},x_{1},\ldots,x_{k})\in V^{k+1}$ of vertices of $G$ a $k$-chain of $G$ and define its length by $\ell(\vec{x})=d(x_{0},x_{1})+d(x_{1},x_{2})+\cdots+d(x_{k-1},x_{k})$. We say a $k$-chain $\vec{x}$ is proper if adjacent nodes are different $x_{0}\neq x_{1}\neq\cdots\neq x_{k}$. Let $P_{k}(G)$ be the set of proper $k$-chains of $G$, and $P_{k,\ell}(G)$ the set of proper $k$-chains of length $\ell$. The following result of Leinster is the starting point of the categorification of graph magnitude introduced by Hepworth–Willerton [16].
Proposition 2.7 (Proposition 3.9 of [25]).
For a graph $G$,
$\mathrm{Mag}(G)=\sum_{k=0}^{\infty}(-1)^{k}\sum_{\vec{x}\in P_{k}(G)}q^{\ell(\vec{x})}\in\mathbb{Z}[\![q]\!].$
In other words, if $\mathrm{Mag}(G)=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell}\in\mathbb{Z}[\![q]\!]$, then
$c_{\ell}=\sum_{k=0}^{\ell}(-1)^{k}\#P_{k,\ell}(G).$
In particular, $V=c_{0}$ and $E=\frac{-c_{1}}{2}$.
Let $MC_{k,\ell}(G)$ be the free abelian group generated by $P_{k,\ell}(G)$. We define the boundary operator $\partial:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ as $\partial=\sum_{i=1}^{k-1}(-1)^{i-1}\partial_{i}$, where $\partial_{i}:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ is defined by
$\partial_{i}(x_{0},\ldots,x_{k})=\begin{cases}(x_{0},\ldots,\widehat{x_{i}},\ldots,x_{k}),&\text{ if }d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1});\\ 0,&\text{ otherwise},\end{cases}$
where $\widehat{x_{i}}$ means deletion of $x_{i}$. One can verify $\partial^{2}=0$. Therefore we call the homology of the chain complex $(MC_{*,\ell}(G),\partial)$ the length $\ell$ magnitude homology of $G$, denoted by $MH_{*,\ell}(G)$. Since the boundary map preserves endpoints, we have a decomposition
$MC_{k,\ell}(G)=\bigoplus_{u,v\in V(G)}MC_{k,\ell}(G;u,v),$
where $MC_{k,\ell}(G;u,v)$ is the subgroup generated by chains $P_{k,\ell}(G;u,v)=\{\vec{x}\in P_{k,\ell}(G)\mid x_{0}=u,x_{k}=v\}$, forming subcomplexes and inducing decomposition of homology groups
(2.1) $MH_{k,\ell}(G)=\bigoplus_{u,v\in V(G)}MH_{k,\ell}(G;u,v).$
Let us denote the magnitude Betti number $\beta_{k,\ell}(G)=\mathrm{rank}MH_{k,\ell}(G)$ and the length $\ell$ magnitude Euler characteristic $\chi_{\ell}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}(G)$.
Proposition 2.8 (Propositions 2.9, 2.10 of [16]).
Let $G=(V,E)$ be a graph.
  1. (1)
    $MH_{0,0}(G)$ is free abelian of rank $\beta_{0,0}(G)=\#V$.
  2. (2)
    $MH_{1,1}(G)$ is free abelian of rank $\beta_{1,1}(G)=2\#E$.
  3. (3)
    If $MH_{k,\ell}(G)\neq 0$, then $\ell/\mathrm{diam}(G)\leq k\leq\ell$ and the inequality $\ell/\mathrm{diam}(G)<k$ is strict if $\mathrm{diam}(G)>1$ and $\ell>0$.
The following proposition is straightforward from definition, which says magnitude can be recovered from magnitude homology.
Proposition 2.9 (Theorem 2.8 of [16]).
Let $G$ be a graph, and $\mathrm{Mag}(G)=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell}\in\mathbb{Z}[\![q]\!]$. Then
$c_{\ell}=\chi_{\ell}(G)=\sum_{k=0}^{\infty}(-1)^{k}\beta_{k,\ell}(G)=\sum_{k=\ell/\mathrm{diam}(G)}^{\ell}(-1)^{k}\beta_{k,\ell}(G).$
Hepworth–Willerton proved that magnitude homology is a functor.
Proposition 2.10 (Proposition 3.3 of [16]).
Magnitude homology is a functor from the category of finite graphs with distance non-increasing maps to the category of bigraded abelian groups.
A direct application of the functorality is as follow. Let $G$ be a connected graph and $H$ be a gated subgraph of $G$ with gate projection $p:G\to H$. Let $\iota:H\to G$ be the inclusion. Both $p$ and $\iota$ are distance non-increasing. The composition $p\circ\iota=\text{id}_{H}$ induces $\text{id}=p_{*}\circ\iota_{*}:MH_{k,\ell}(H)\to MH_{k,\ell}(G)\to MH_{k,\ell}(H)$. This gives a decomposition of $MH_{k,\ell}(G)$.
Proposition 2.11.
If $p:G\to H$ is a gate projection, then
$MH_{k,\ell}(G)\cong MH_{k,\ell}(H)\oplus\ker(p_{*}).$
There is a Künneth formula for magnitude homology of graphs.
Theorem 2.12 (Theorem 5.3 of [16]).
There is a natural short exact sequence which non-naturally splits
$0\to MH_{*,*}(G)\otimes MH_{*,*}(H)\xrightarrow{\square}MH_{*,*}(G\square H)\to\mathrm{Tor}(MH_{*-1,*}(G),MH_{*,*}(H))\to 0,$
where $\square$ is the exterior product defined in Definition 5.2 of [16].
Example 2.13.
For the complete graph $K_{n}$,
$MH_{k,\ell}(K_{n})=\begin{cases}\bigoplus_{\vec{x}\in P_{\ell}(K_{n})}\mathbb{Z}\vec{x},&\text{ if }k=l;\\ 0,&\text{ otherwise}.\end{cases}$
The Betti number $\beta_{\ell,\ell}(K_{n})=n(n-1)^{\ell}$. One can verify $\mathrm{Mag}(K_{n},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(K_{n})q^{\ell}=\frac{n}{1+(n-1)q}$.
Example 2.14.
For the hypercube graph $Q_{d}=K_{2}^{\square d}$, we use the Künneth formula (Theorem 2.12) to obtain
$MH_{k,\ell}(Q_{d})=\bigoplus_{\sum k_{i}=k,\sum\ell_{i}=\ell}MH_{k_{1},\ell_{1}}(K_{2})\otimes\cdots\otimes MH_{k_{d},\ell_{d}}(K_{2}),$
and use the result of $K_{2}$, we conclude
$MH_{k,\ell}(Q_{d})=\begin{cases}\mathbb{Z}^{\beta_{\ell,\ell}(Q_{d})},&\text{ if }k=l;\\ 0,&\text{ otherwise},\end{cases}$
where $\beta_{\ell,\ell}(Q_{d})=2^{d}\binom{l+d-1}{d-1}$. One can verify $\mathrm{Mag}(Q_{d},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(Q_{d})q^{\ell}=\frac{2^{d}}{(1+q)^{d}}$.
Definition 2.15.
We say $G$ is diagonal if $MH_{k,\ell}(G)=0$ for $k\neq\ell$.
Example 2.16.
We have seen that $K_{n}$ and $Q_{d}$ are diagonal. In particular the cycle graph $C_{4}=Q_{2}$ is diagonal. However Gu [15] computed $MH_{k,\ell}(C_{n})$ for all $n\geq 5$ and showed that $C_{n}(n\geq 5)$ are not diagonal.
Let us recall results from [19] in the context of graphs. Let $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$. We say $\vec{x}$ is geodesic if $\ell(\vec{x})=d(x_{0},x_{k})$. Any sub-chain of a geodesic proper chain is again geodesic proper. Any proper $1$-chain is geodesic. A proper $2$-chain $(x_{0},x_{1},x_{2})$ is geodesic if and only if $d(x_{0},x_{2})=d(x_{0},x_{1})+d(x_{1},x_{2})$.
Consider $u,v\in V(G)$, such that $\ell=d(u,v)$, then any chain in $P_{k,\ell}(G;u,v)$ must be geodesic.
Definition 2.17.
Let us define the length $\ell$ geodesic magnitude complex
$MC_{*,\ell}^{geod}(G):=\bigoplus_{u,v\in V(G):d(u,v)=\ell}MC_{*,\ell}(G;u,v).$
The homology is called the length $\ell$ geodesic magnitude homology, which admits the decomposition
$MH_{k,\ell}^{geod}(G)=\bigoplus_{u,v\in V(G):d(u,v)=\ell}MH_{k,\ell}(G;u,v).$
Note that $MC_{*,\ell}^{geod}(G)$ and $MH_{k,\ell}^{geod}(G)$ are trivial if $\ell>\mathrm{diam}(G)$.
Let us denote the geodesic magnitude Betti number $\beta_{k,\ell}^{geo}(G)=\mathrm{rank}MH_{k,\ell}^{geod}(G)$ and the length $\ell$ geodesic magnitude Euler characteristic $\chi_{\ell}^{geod}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}^{geod}(G)$. Notice that $\beta_{k,\ell}(G)\geq\beta_{k,\ell}^{geod}(G)$.
Kaneta–Yoshinaga [19] generalized the above geodesic magnitude chains.
Definition 2.18.
A proper $3$-chain $(x_{0},x_{1},x_{2},x_{3})\in P_{3}(G)$ is called a $4$-cut of $G$ if $(x_{0},x_{1},x_{2})$ and $(x_{1},x_{2},x_{3})$ are geodesic but $(x_{0},x_{1},x_{2},x_{3})$ is not geodesic. Write $m_{G}$ as the minimum length of $4$-cuts of $G$. We say $m_{G}=\infty$ if $G$ has no $4$-cuts.
For a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$, an internal node $x_{i}(i=1,2,\ldots,k-1)$ is called a smooth node of $\vec{x}$ if $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$. A node that is not smooth is called a singular node. Note that endpoints are singular. We write $\varphi(\vec{x})$ as the sub-chain of $\vec{x}$ consisting of all singular nodes, called the frame of $\vec{x}$. Note that $\varphi(\vec{x})$ is not necessarily proper.
Definition 2.19.
A proper chain $\vec{x}$ is geodesically simple if $\ell(\varphi(\vec{x}))=\ell(\vec{x})$.
Note that proper $1$-chains and proper $2$-chains are all geodesically simple, and proper chains with length smaller than $m_{G}$ are geodesically simple. Kaneta–Yoshinaga proved that the subgroups $MC_{*,\ell}^{gs}(G)$ of $MC_{*,\ell}(G)$ generated by geodesically simple chains is a subcomplex, whose homology is called the geodesically simple magnitude homology of $G$, denoted by $MH_{*,\ell}^{gs}(G)$, and the homology group $MH_{k,\ell}^{gs}(G)$ factors as a direct sum of framed magnitude homology groups as explained below.
Let $\vec{a}\in P_{m,\ell}(G)$ and $k\geq m$. We define $P_{k,\vec{a}}(G):=\{\vec{x}\in P_{k,\ell}(G)\mid\varphi(\vec{x})=\vec{a}\}$. Note that $P_{k,\vec{a}}(G)$ consists of geodesically simple chains and would be empty if $\varphi(\vec{a})\neq\vec{a}$. We further define $MC_{k,\vec{a}}(G)$ as the free abelian group generated by $P_{k,\vec{a}}(G)$, which is a subgroup of $MC_{k,\ell}^{gs}(G)$. They form a subcomplex $MC_{*,\vec{a}}(G)$ whose homology $MH_{*,\vec{a}}(G)$ is called the $\vec{a}$-framed magnitude homology of $G$. Kaneta–Yoshinaga proved the chain complex decomposes
(2.2) $MC_{k,\ell}^{gs}(G)=\bigoplus_{m\leq k,\vec{a}\in P_{m,\ell}(G)}MC_{k,\vec{a}}(G),$
and so does the homology group.
Theorem 2.20 (Theorem 3.12 of [19]).
Let $G$ be a graph and $k,\ell>0$. We have
$MH_{k,\ell}^{gs}(G)\cong\bigoplus_{m\leq k,\vec{a}\in P_{m,\ell}(G)}MH_{k,\vec{a}}(G).$
Furthermore, if $k=1$ or $\ell<m_{G}$, then $MH_{k,\ell}^{gs}(G)\cong MH_{k,\ell}(G)$.
Kaneta–Yoshinaga also related framed magnitude homology with open intervals determined by the frame. Given $a,b\in V(G)$, consider the open interval $(a,b)_{G}$ introduced in Section 2.1. Also write $C_{*}(a,b)_{G}$ the reduced chain complex of the order complex of $(a,b)_{G}$.
Proposition 2.21 (Proposition 2.3 of [13]).
For $a,b\in G$ with $d(a,b)=\ell$, there is an isomorphism of chain complexes
$MC_{*,\ell}(G;a,b)\to C_{*-2}(a,b)_{G},$
inducing an isomorphism of homology groups
$MH_{k,\ell}(G;a,b)\xrightarrow{\cong}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
Proof.
For a proper chain $\vec{x}=(a,x_{1},\ldots,x_{k-1},b)\in P_{k,\ell}(G;a,b)$, it must be geodesic since $\ell=d(a,b)$. Then mapping $\vec{x}$ to the linearly order chain $x_{1}\prec_{G}\cdots\prec_{G}x_{k-1}$ in $C_{k-2}(a,b)_{G}$ defines the desired isomorphism. ∎
Corollary 2.22.
The geodesic magnitude homology has the decomposition
$MH_{k,\ell}^{geod}(G)\cong\bigoplus_{a,b\in V(G):d(a,b)=\ell}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
For a frame $\vec{a}=(a_{0},a_{1},\ldots,a_{m})\in P_{m,\ell}(G)$, we denote by $C_{*}(\vec{a})_{G}$ the total complex of $C_{*}(a_{0},a_{1})_{G}\otimes\cdots\otimes C_{*}(a_{m-1},a_{m})_{G}$. A more generalized version of Proposition 2.21 is as follow.
Theorem 2.23 (Theorem 4.4 and Corollary 4.5 of [19]).
Suppose a frame $\vec{a}$ has length $\ell(\vec{a})<m_{G}$, then there is an isomorphism of chain complexes
$MC_{*,\vec{a}}(G)\to C_{*-2m}(\vec{a})_{G}$
grouping smooth nodes in order between $a_{i}$ and $a_{i+1}$, which induces isomorphisms of homology groups
$MH_{k,\vec{a}}(G)\cong\widetilde{H}_{k-2m}(C_{*}(\vec{a})_{G}).$
2.4. Hyperplane arrangements
In this paper, all arrangements are real central hyperplane arrangements unless otherwise specified. Standard references are [27, 30]. Let $\mathcal{A}$ be an arrangement in $\mathbb{R}^{\ell}$. We denote by $L(\mathcal{A})=\{\cap_{H\in\mathcal{B}}H\mid\mathcal{B}\subseteq\mathcal{A}\}$ the intersection lattice of $\mathcal{A}$, ordered by reverse inclusions. An element of $L(\mathcal{A})$ is called a flat. Note that the maximal flat is $\hat{1}=\cap_{H\in\mathcal{A}}H=:\cap\mathcal{A}$, called the center of $\mathcal{A}$, and the minimal flat $\hat{0}$ is the ambient space $\mathbb{R}^{\ell}$. The arrangement is called essential if the center is the origin. The rank of a flat $X\in L(\mathcal{A})$ is defined as the codimension $\mathrm{rank}X=\mathrm{codim}X$. The rank of $\mathcal{A}$ is defined as $\mathrm{rank}\mathcal{A}=\mathrm{rank}(\cap\mathcal{A})$. We write $L_{k}(\mathcal{A})$ as the set of flats of rank $k$ and $L_{\leq k}(\mathcal{A})$ the set of flats of rank at most $k$. Using the Möbius function $\mu:L(\mathcal{A})\times L(\mathcal{A})\to\mathbb{Z}$, we define the characteristic polynomial of $\mathcal{A}$ by $\chi(\mathcal{A},t)=\sum_{X\in L(\mathcal{A})}\mu(\hat{0},X)t^{\dim X}$. It is classically known $\mathrm{Ch}(\mathcal{A})=|\chi(\mathcal{A},1)|$ [33].
The arrangement $\mathcal{A}$ determines a stratification of $\mathbb{R}^{\ell}$, the strata are called faces (covectors). Let $\mathcal{F}(\mathcal{A})$ be the poset of faces, ordered by closure inclusions. The unique minimal face is the center $\cap\mathcal{A}=\cap_{H\in\mathcal{A}}H$, and maximal faces are the chambers $\mathrm{Ch}(\mathcal{A})$. For faces $F,G\in\mathcal{F}(\mathcal{A})$, their meet $F\wedge G$ always exists, while their join $F\vee G$ may not exist. There is an order-reversing surjective map $s:\mathcal{F}(\mathcal{A})\to L(\mathcal{A})$ taking a face $F$ to its support $s(F)$, that is the minimal subspace of $\mathbb{R}^{\ell}$ containing $F$. We define the dimension of a face as the dimension of its support.
For $F\in\mathcal{F}(\mathcal{A})$, let $\mathcal{A}_{F}=\{H\in\mathcal{A}\mid F\subseteq H\}$. Also, for a flat $X\in L(\mathcal{A})$, let $\mathcal{A}_{X}=\{H\in\mathcal{A}\mid X\subseteq H\}$ and $\mathcal{A}^{X}=\{H\cap X\mid H\in\mathcal{A}\setminus\mathcal{A}_{X}\}$. Note that if $F,G\in\mathcal{F}(\mathcal{A})$ have the same support $X\in L(\mathcal{A})$, then $\mathcal{A}_{F}=\mathcal{A}_{G}=\mathcal{A}_{X}$.
Definition 2.24.
For a flat $X\in L(A)$, we denote by
$c^{X}:=\#s^{-1}(X)$
the number of faces whose support is $X$, which is also equal to $\mathrm{Ch}(\mathcal{A}^{X})$. If $X$ is not the ambient space $\hat{0}$$\mathbb{R}^{\ell}$, we also define the beta invariant of $X$ by
$\beta_{X}:=\beta(\mathcal{A}_{X}).$
Recall that the beta invariant of an arrangement is defined as
$\beta(\mathcal{A})=(-1)^{\mathrm{rank}\mathcal{A}-1}\chi^{\prime}(\mathcal{A},1)=|\chi^{\prime}(\mathcal{A},1)|.$
Remark 2.25.
The beta invariant $\beta_{X}$ counts the bounded chambers in the deconing of $\mathcal{A}_{X}$. In other words, let $F$ be a face whose support is $X$ and pick $H\in\mathcal{A}$ such that $F\subseteq H$, then $\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid\overline{C}\cap H=\overline{F}\}$ (See [33, 2, 30]). Hence the product $c^{X}\beta_{X}$ is also a counting function
$c^{X}\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid s(\overline{C}\cap H)=X\}.$
For arrangement $\mathcal{A}$ in $V=\mathbb{R}^{\ell}$, we fix a defining form $\alpha_{H}\in V^{*}$ for each $H\in\mathcal{A}$ and consider the line segment $L_{H}=[-\alpha_{H},\alpha_{H}]$ in $V^{*}$. We write $Z(\mathcal{A})=\sum_{H\in\mathcal{A}}L_{H}$ as the Minkowski sum of the line segments and call it the zonotope of $\mathcal{A}$. The poset of nonempty faces of this polytope is anti-isomorphic to $\mathcal{F}(\mathcal{A})$. For a face $F\in\mathcal{F}(A)$, we say the corresponding face $z_{F}$ of $Z(\mathcal{A})$ dual to $F$. In particular, 0-faces (vertices) of $Z(\mathcal{A})$ are dual to chambers of $\mathcal{A}$.
With these $\alpha_{H}$ chosen, each face $F\in\mathcal{F}(\mathcal{A})$ can be represented by a sign vector $F=(F_{H})_{H\in\mathcal{A}}\subseteq\{+,-,0\}^{\mathcal{A}}$, where $F_{H}=\mathrm{sgn}(\alpha_{H}(F))$. Declare $0<\pm$ and this determines a partial order on sign vectors coordinate-wisely, which coincides with the face order of $\mathcal{F}(\mathcal{A})$. We define the Tits product on $\mathcal{F}(\mathcal{A})$ using sign vectors as follow. For $F,G\in\mathcal{F}(\mathcal{A})$, let $FG\in\mathcal{F}(A)$ be the sign vector
$(FG)_{H}=\begin{cases}F_{H},\text{ if }F_{H}\neq 0;\\ G_{H},\text{ if }F_{H}=0.\end{cases}$
This product does not depend on the choices of defining forms, since geometrically, $FG$ is the face we land in when we start from a relative interior point of $F$, move out a small enough distance towards a relative interior point of $G$ (see Figure 1).
[…]
Figure 1. A geometric visualization of the non-commutative Tits product in the $A_{2}$ arrangement. Let $F$ be a ray (blue) and $G$ be a chamber (red). (a) The product FGFG represents taking a small step from $F$ towards $G$, which immediately lands in the adjacent chamber. (b) The product GFGF represents taking a small step from $G$ towards $F$. Because $G$ is an open chamber, a sufficiently small step remains entirely within $G$, so $GF=G$.
2.5. Tope graphs and graph-theoretic properties
For an arrangement $\mathcal{A}$ in $\mathbb{R}^{\ell}$, notice that the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the $1$-skeleton of $Z(\mathcal{A})$. For a face $F\in\mathcal{F}(A)$, we let $\mathcal{T}(\mathcal{A})_{F}$ be the induced subgraph of $\mathcal{T}(\mathcal{A})$ with vertices $\mathrm{Ch}(\mathcal{A})_{F}=\{C\in\mathrm{Ch}(\mathcal{A})\mid F\leq C\}$, or equivalently, $\mathcal{T}(\mathcal{A})_{F}$ is the $1$-skeleton of the face $z_{F}$ of $Z(\mathcal{A})$. We call $\mathcal{T}(\mathcal{A})_{F}$ the $F$-subgraph of $\mathcal{T}(\mathcal{A})$. We first prove a useful lemma.
Lemma 2.26.
For $F\in\mathcal{F}(\mathcal{A})$, $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph of $\mathcal{T}(\mathcal{A})$.
Proof.
We prove that the map $p_{F}:\mathcal{T}(\mathcal{A})\to\mathcal{T}(\mathcal{A})_{F}$ defined by $p_{F}(D)=FD$ is the gate projection, that is, we prove the equation
(2.3) $d(D,C)=d(D,FD)+d(FD,C)$
holds for $D\in\mathrm{Ch}(\mathcal{A})$ and $C\in\mathrm{Ch}(\mathcal{A})_{F}$. In terms of sign vectors, the distance $d(X,Y)=\sum_{H\in\mathcal{A}}d(X,Y)_{H}$ is the Hamming distance (number of non-equal signs), where
$d(X,Y)_{H}=\begin{cases}1,\text{ if }X_{H}\neq Y_{H};\\ 0,\text{ if }X_{H}=Y_{H}.\end{cases}$
Let us prove both sides of equation (2.3) have the same evaluation on each $H\in\mathcal{A}$. If $H\notin\mathcal{A}_{F}$, then $F_{H}\neq 0$ and $(FD)_{H}=F_{H}$, also since $C\in\mathrm{Ch}(\mathcal{A})_{F}$, we have $C_{H}=F_{H}$. Therefore (2.3) evaluated at $H$ is $d(D,C)_{H}=d(D,FD)_{H}+0$, which is trivially true since both sides equal $1$ if $D_{H}\neq F_{H}$ and 0 if $D_{H}=F_{H}$. Next suppose $H\in\mathcal{A}_{F}$, we have $F_{H}=0$. Then $d(D,FD)_{H}=0$. The equation (2.3) evaluated at $H$ is $d(D,C)_{H}=0+d(FD,C)_{H}$, which is triviall true since $D_{H}=(FD)_{H}$. ∎
Here we record some important examples of tope graphs.
Example 2.27.
Let $\mathcal{A}=Bl_{d}$ be the $d$-th Boolean arrangement, that is, the arrangement of $d$ coordinate hyperplanes of $\mathbb{R}^{d}$. It is clear that the tope graph $\mathcal{T}(\mathcal{A})$ is the whole hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{d}$.
Example 2.28.
Let $r\geq 2$. An arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$ is said to be generic if any subarrangement of $r$ hyperplanes has rank $r$. Boolean arrangements are generic. Generic arrangements are realizations of uniform matroids $U_{r,n}$. By abuse of notations, we use $U_{r,n}$ to denote a generic arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$. For example, consider a generic arrangement $\mathcal{A}=U_{r,r+1}$. The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is isomorphic to the induced subgraph of $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{r+1}$ with all vertices (sign vectors) except the all ++ vector and the all - vector.
Example 2.29.
Consider a finite Coxeter system $(W,S)$ and the associated Coxeter arrangement $\mathcal{A}_{W}$. The zonotope $Z(\mathcal{A}_{W})$ is the $W$-permutohedron, and the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the Cayley graph of $(W,S)$.
Tope graphs have the following graph-theoretic properties. Recall that a chamber $C$ is simplicial if its intersection with the unit sphere of the ambient space is combinatorially a simplex. An arrangement is simplicial if all chambers all simplicial.
Proposition 2.30 (Propositions 4.2.15 and 4.4.8 of [4]).
Let $\mathcal{A}$ be an arrangement of $\mathrm{rank}\mathcal{A}=r$. The tope graph $\mathcal{T}=\mathcal{T}(\mathcal{A})$ has the following properties as an unlabeled graph.
  1. (1)
    For each vertex $C$, there is a unique vertex C-C such that $d(C,-C)=\mathrm{diam}(\mathcal{T})$.
  2. (2)
    The mapping $C\mapsto-C$ is a fixed point free automorphism of $\mathcal{T}$.
  3. (3)
    $\mathcal{T}$ is bipartite.
  4. (4)
    For each vertex $C$, $\deg(C)\geq r$, where the equality holds iff $C$ is simplicial.
  5. (5)
    $\mathcal{T}$ is $r$-connected.
Recall that a graph is called $k$-regular if every vertex has degree $k$.
Corollary 2.31.
$\mathcal{A}$ is simplicial of rank $r$ if and only if $\mathcal{T}(\mathcal{A})$ is $r$-regular.
A matching of a graph is a collection of edges where no two edges share a common incident vertex. We say a matching is perfect if every vertex of the graph is incident to exactly one edge in the matching. The marriage theorem (see Theorem 2.1.2 of [9]) has a corollary (see Corollary 2.1.3 of [9]) asserting that every $k$-regular ($k\geq 1$) bipartite graph has a perfect matching.
Corollary 2.32.
If $\mathcal{A}$ is simplicial, then $\mathcal{T}(\mathcal{A})$ has a perfect matching.
Very recently, Hamiltonian cycles in tope graphs of simplicial and supersolvable arrangements are studied in [5, 20].
2.6. Varchenko matrix
Varchenko introduced a bilinear form and computed the determinant [31]. Generalization to oriented matroids can be found in [17]. See [8] and Sections 8.4-8.5 of [2] for details.
Here we only consider the equal weight case. For a real hyperplane arrangement $\mathcal{A}$, we define the $q$-Varchenko matrix
$\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]_{C,D\in\mathrm{Ch}(\mathcal{A})}.$
In other words, $\mathcal{V}_{q}(\mathcal{A})$ is the Zeta matrix of the tope graph $\mathcal{T}(\mathcal{A})$ (see Section 2.2). It can also be defined for oriented matroids.
Theorem 2.33 ([31], see also Theorem 8.23 of [2]).
The determinant of the $q$-Varchenko matrix is
$\det\mathcal{V}_{q}(\mathcal{A})=\prod_{X\in L(\mathcal{A})\setminus\{\hat{0}\}}(1-q^{2|\mathcal{A}_{X}|})^{c^{X}\beta_{X}}$
2.7. Tope poset and generating function
Fix a base chamber $B\in\mathrm{Ch}(\mathcal{A})$ and define a partial order on $\mathrm{Ch}(\mathcal{A})$ by $C\preceq C^{\prime}$ if $S(B,C)\subseteq S(B,C^{\prime})$. This poset is called the tope poset of $\mathcal{A}$ based at $B$, denoted by $\mathcal{T}(\mathcal{A},B)$. The poset $\mathcal{T}(\mathcal{A},B)$ has bottom element $B$ and top element B-B, graded by rank function $\rho_{B}(C)=d(B,C)$. In terms of notations in Section 2.1, $\mathcal{T}(\mathcal{A},B)=[B,-B]_{\mathcal{T}(\mathcal{A})}$.
Proposition 2.34 (Theorem 2.2 of [10], see also Theorem 4.4.2 of [4]).
An open interval $(C_{1},C_{2})$ inside $\mathcal{T}(\mathcal{A},B)$ is homotopy equivalent to a sphere $S^{\mathrm{rank}\mathcal{A}-\dim F-2}$, if the closed interval $C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$; contractible otherwise.
Corollary 2.35 (Theorem 1.11 of [11], see also Corollary 4.4.3 of [4]).
The Möbius function on $\mathcal{T}(\mathcal{A},B)$ is
$\mu(C_{1},C_{2})=\begin{cases}(-1)^{\mathrm{rank}\mathcal{A}-\dim F},&\text{ if }[C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F},\exists F\in\mathcal{F}(\mathcal{A});\\ 0,&\text{ otherwise}.\end{cases}$
Proposition 2.36 (Theorem 3.4 of [3]).
If $\mathcal{A}$ is simplicial, then $\mathcal{T}(\mathcal{A},B)$ is a lattice for any $B\in\mathrm{Ch}(\mathcal{A})$.
Let $D_{\mathcal{A},B}(q)=\sum_{C\in\mathrm{Ch(\mathcal{A})}}q^{d(B,C)}$ be the rank generation function of $\mathcal{T}(\mathcal{A},B)$. Note that it is the sum of entries of the $B$-row/column of the $q$-Varchenko matrix $\mathcal{V}_{q}(\mathcal{A})$.
Proposition 2.37 ([29]).
If $\mathcal{A}$ is a Coxeter arrangement with exponents $e_{1},\ldots,e_{\ell}$, then
$D_{\mathcal{A},B}(q)=\prod_{i=1}^{\ell}[1+e_{i}]_{q}$
for every $B\in\mathrm{Ch}(\mathcal{A})$, where $n]_{q}=1+q+q^{2}+\cdots+q^{n-1}$ is the $q$-analog.
The generating function decomposition formula is generalized to supersolvable arrangements [3], inductively factored arrangements [18]. See also [1, 26, 12].
3. Magnitude of tope graphs
For a real hyperplane arrangement $\mathcal{A}$, we define the magnitude $\mathrm{Mag}(\mathcal{A})=\mathrm{Mag}(\mathcal{A},q)$ of $\mathcal{A}$ as the magnitude of $\mathcal{T}(\mathcal{A})$.
3.1. Basic properties
We first prove a reciprocity result.
Proposition 3.1.
Let $\mathcal{A}$ be a central hyperplane arrangement with $N=\#\mathcal{A}$ hyperplanes. Then
$\mathrm{Mag}(\mathcal{A},q^{-1})=q^{N}\mathrm{Mag}(\mathcal{A},q).$
Proof.
Recall that there is an involution $\tau:\mathrm{Ch}(\mathcal{A})\to\mathrm{Ch}(\mathcal{A})$ sending $D$ to D-D. For any two chambers $C,D\in\mathrm{Ch}(\mathcal{A})$, we have d(C,D)+d(C,D)=Nd(C,D)+d(C,-D)=N. Then the $q$-Varchenko matrix $\mathcal{V}_{q}=\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]$ satisfies
$q^{N}\mathcal{V}_{q^{-1}}=[q^{N-d(C,D)}]=[q^{d(C,-D)}]=\mathcal{V}_{q}P_{\tau},$
where $P_{\tau}$ is the permutation matrix induced by $\tau$. By definition of magnitude,
$\mathrm{Mag}(\mathcal{A},q^{-1})=\mathbf{1}^{T}(\mathcal{V}_{q^{-1}})^{-1}\mathbf{1}=\mathbf{1}^{T}(q^{-N}\mathcal{V}_{q}P_{\tau})^{-1}\mathbf{1}=q^{N}\mathbf{1}^{T}P_{\tau}^{-1}\mathcal{V}_{q}^{-1}\mathbf{1}=q^{N}\mathbf{1}^{T}\mathcal{V}_{q}^{-1}\mathbf{1}=q^{N}\mathrm{Mag}(\mathcal{A},q),$
where we used $\mathbf{1}^{T}P_{\tau}^{-1}=\mathbf{1}^{T}$. ∎
Since tope graphs are partial cubes, the magnitude $\mathrm{Mag}(\mathcal{A})$ has the one-point property by [22].
Proposition 3.2.
Let $\mathcal{A}$ be a central hyperplane arrangement, then $\mathrm{Mag}(\mathcal{A},1)=1$.
Next we prove some main structural results of $\mathrm{Mag}(\mathcal{A})$.
Theorem 3.3.
For central arrangement $\mathcal{A}$, let $\mathrm{Mag}(\mathcal{A},q)=P(q)/Q(q)$ be the reduced form, with the leading coefficient of $Q(q)$ positive. Then
  1. (1)
    $\mathcal{A}=\deg Q-\deg P$.
  2. (2)
    $P,Q$ are palindromic.
  3. (3)
    Roots of $Q$ are all roots of unity, excluding $1$.
Proof.
Let $P(q)=a_{m}q^{m}+\cdots+a_{0}$ and $Q(q)=b_{n}q^{n}+\cdots+b_{0}$ where $a_{m}\neq 0$ and $b_{n}\neq 0$. Then $m=\deg P$ and $n=\deg Q$. Since $\mathrm{Mag}(\mathcal{A},0)=a_{0}/b_{0}=\#\mathrm{Ch}(\mathcal{A})$ (Proposition 2.7), we have $a_{0}\neq 0$ and $b_{0}\neq 0$. By the reciprocity formula (Proposition 3.1),
$q^{N}=\frac{\mathrm{Mag}(\mathcal{A},q^{-1})}{\mathrm{Mag}(\mathcal{A},q)}=\frac{P(q^{-1})Q(q)}{P(q)Q(q^{-1})}.$
Multiplying both sides by $q^{m-n}$, we have
$q^{N+m-n}=\frac{q^{m}P(q^{-1})Q(q)}{P(q)q^{n}Q(q^{-1})}=\frac{(a_{m}+\cdots+a_{0}q^{m})(b_{n}q^{n}+\cdots+b_{0})}{(a_{m}q^{m}+\cdots+a_{0})(b_{n}+\cdots+b_{0}q^{n})}.$
Comparing degrees, we conclude N+mn=0N+m-n=0 and (1) is proved.
Let us write $P^{*}(q)=q^{m}P(q^{-1})=a_{m}+\cdots+a_{0}q^{m}$ and $Q^{*}(q)=q^{n}Q(q^{-1})=b_{n}+\cdots+b_{0}q^{n}$ as the reverse polynomials. We have proved
$\mathrm{Mag}(\mathcal{A},q)=\frac{P(q)}{Q(q)}=\frac{P^{*}(q)}{Q^{*}(q)}.$
Note that the fraction $P^{*}(q)/Q^{*}(q)$ is reduced since if $P^{*}$ and $Q^{*}$ share a common root $\alpha\neq 0$, $P$ and $Q$ would share a common root $1/\alpha$, contradicting to the assumption $P/Q$ is reduced. Now we must have $P^{*}(q)=rP(q)$ and $Q^{*}(q)=rQ(q)$ for some constant $r\neq 0$. Substituting $q=1$ into the equations $Q^{*}(q)=q^{n}Q(1/q)$ and $Q^{*}(q)=rQ(q)$, we obtain
$Q^{*}(1)=Q(1)=rQ(1).$
Since $\mathrm{Mag}(\mathcal{A},1)=P(1)/Q(1)=1$ (Proposition 3.2) and $P/Q$ is reduced, $Q(1)\neq 0$, we conclude $r=1$. Then $P^{*}=P$ and $Q^{*}=Q$, so (2) is proved.
Since $\mathcal{V}_{q}^{-1}=\frac{1}{\det\mathcal{V}_{q}}\mathrm{adj~}\mathcal{V}_{q}$, roots of $Q(q)$ are all roots of $\det\mathcal{V}_{q}$, which are roots of unity by Theorem 2.33. We have already seen $Q(1)\neq 0$. So (3) is proved. ∎
If an arrangement $\mathcal{A}$ decomposes as a direct sum $\mathcal{A}_{1}\oplus\mathcal{A}_{2}$, then the tope graph decomposes as cartesian product $\mathcal{T}(\mathcal{A})=\mathcal{T}(\mathcal{A}_{1})\square\mathcal{T}(\mathcal{A}_{2})$. So by Proposition 2.5, we have a decomposition of magnitude.
Proposition 3.4.
$\mathrm{Mag}(\mathcal{A}_{1}\oplus\mathcal{A}_{2})=\mathrm{Mag}(\mathcal{A}_{1})\mathrm{Mag}(\mathcal{A}_{2})$.
Coxeter arrangements have the natural action of the Coxeter groups on the chambers, which makes the magnitude easy to compute.
Proposition 3.5.
For Coxeter arrangement $\mathcal{A}_{W}$ with exponents $\{e_{1},\ldots,e_{\ell}\}$, we have
$\mathrm{Mag}(\mathcal{A}_{W})=\frac{\#W}{\prod_{i}[1+e_{i}]_{q}}.$
Proof.
The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is the $1$-skeleton of the $W$-permutohedron, or equivalently the Cayley graph of $W$ with respect to the standard generating set (Example 2.29), hence is vertex-transitive. The result follows from Proposition 2.3 and Proposition 2.37. ∎
3.2. Face decomposition conjecture
For the purpose of reciprocity, let us define the following.
Definition 3.6.
We define the interior magnitude of $\mathcal{A}$ by
$\mathrm{Mag}^{\circ}(\mathcal{A},q):=(-1)^{\mathrm{rank}\mathcal{A}}\mathrm{Mag}(\mathcal{A},q^{-1})=(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}\mathrm{Mag}(\mathcal{A},q),$
where the latter equality is Proposition 3.1.
Let us propose a main conjecture of the magnitude of arrangements.
Conjecture 3.7.
We have a face decomposition of the magnitude.
$\mathrm{Mag}(\mathcal{A})$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\mathrm{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})$
$=\sum_{X\in L(\mathcal{A})}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})=\sum_{X\in L(\mathcal{A})}(-1)^{\mathrm{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X}),$
where $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$. Equivalently,
$\mathrm{Mag}(\mathcal{A})$ $=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\mathrm{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}$
$=\frac{\sum_{X\in L(\mathcal{A})\setminus\{\cap\mathcal{A}\}}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{X\in L(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\mathrm{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}.$
Remark 3.8.
This formula is the discrete $q$-analogue of the decomposition of the Euler characteristic of a zonotope into its open faces. Conjecture 3.7 is verified by computer for all examples listed in Section 3.3 and the appendix, and by easy computations for Boolean arrangements and all rank $2$ arrangements.
Corollary 3.9.
Conjecture 3.7 implies that $\mathrm{Mag}(\mathcal{A})$ is determined by $L(\mathcal{A})$.
3.3. Examples
We use the notion $\mathcal{A}(n,k)$ from Grünbaum’s catalogue [14].
Example 3.10.
The near-pencil $\mathcal{A}(n,0)$ has tope graph $C_{2(n-1)}\Box K_{2}$ and hence
$\mathrm{Mag}(\mathcal{A}(n,0))=\frac{4(n-1)}{[2]_{q}^{2}[n-1]_{q}}.$
Example 3.11.
Note that $\mathcal{A}(6,1)$ is the Coxeter arrangement of type $A_{3}$ or $D_{3}$, and $\mathcal{A}(9,1)$ is the Coxeter arrangement of type $B_{3}$, which can be computed from Proposition 3.5. In what follows, $\Phi_{k}(q)\in\mathbb{Z}[q]$ is the $k$-th cyclotomic polynomial.
$\mathrm{Mag}(\mathcal{A}(6,1))$ $=\frac{24}{[4]_{q}!}=\frac{24}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)}$
$=24-72q+96q^{2}-72q^{3}+48q^{4}-72q^{5}+120q^{6}-144q^{7}+144q^{8}-144q^{9}+144q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(9,1))$ $=\frac{48}{[2]_{q}[4]_{q}[6]_{q}}=\frac{48}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{6}(q)}$
$=48-144q+192q^{2}-192q^{3}+240q^{4}-336q^{5}+432q^{6}-528q^{7}+624q^{8}-720q^{9}+816q^{10}+\cdots$
More examples are in the appendix.
4. Magnitude homology of tope graphs
We continue denoting invariants of $\mathcal{T}(\mathcal{A})$ simply by invariants of $\mathcal{A}$, such as $MH_{k,\ell}(\mathcal{A}):=MH_{k,\ell}(\mathcal{T}(\mathcal{A}))$, $\beta_{k,\ell}(\mathcal{A}):=\mathrm{rank}MH_{k,\ell}(\mathcal{A})$ and $\chi_{\ell}(\mathcal{A})=\sum_{k}(-1)^{k}\beta_{k,\ell}(\mathcal{A})$. Also recall the notation $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$ for $X\in L(\mathcal{A})$.
4.1. Small length
Let us start our study of $MH_{k,\ell}(\mathcal{A})$ from investigating small length $\ell\leq 2$.
Proposition 4.1.
Let $\mathcal{A}$ be a real central hyperplane arrangement. Then the only nontrivial $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are as follows.
  1. (1)
    $MH_{0,0}(\mathcal{A})$ is free abelian of rank $\beta_{0,0}(\mathcal{A})=\#\mathrm{Ch}(\mathcal{A})$.
  2. (2)
    $MH_{1,1}(\mathcal{A})$ is free abelian of rank $\beta_{1,1}(\mathcal{A})=2\sum_{H\in\mathcal{A}}c^{H}$.
  3. (3)
    $MH_{2,2}(\mathcal{A})$ is free abelian of rank $\beta_{2,2}(\mathcal{A})=\beta_{1,1}(\mathcal{A})+4\sum_{X\in L_{2}(\mathcal{A}),\#\mathcal{A}_{X}=2}c^{X}$.
Proof.
(1) and (2) follow from (1) and (2) of Proposition 2.8 and noting that an edge in $\mathcal{T}(\mathcal{A})$ corresponds to a chamber of restriction $\mathcal{A}^{H}$ for some $H\in\mathcal{A}$.
For (3), since $m_{\mathcal{T}(\mathcal{A})}\geq 3$, we may apply Theorem 2.20 for $\ell=2<m_{\mathcal{T}(\mathcal{A})}$.
$MH_{2,2}(\mathcal{A})\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})\cup P_{2,2}(\mathcal{A})}MH_{2,\vec{a}}(\mathcal{A}).$
By Theorem 2.23, the framed magnitude homology of $\mathcal{A}$ can be computed from interval posets. For $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$, where $a_{0},a_{1}$ are chambers at distance $2$, we may suppose $S(a_{0},a_{1})=\{H_{1},H_{2}\}$ and let $X=H_{1}\cap H_{2}\in L_{2}(\mathcal{A})$. If $\mathcal{A}_{X}=2$, then the interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ consists of two incomparable chambers between $a_{0}$ and $a_{1}$. If $\mathcal{A}_{X}>2$, then $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ is a singleton. Since $MH_{2,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{0}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})$, taking reduced homology, only the first type $(a_{0},a_{1})$ with $\mathcal{A}_{X}=2$ contributes to the magnitude homology a copy of $\mathbb{Z}$. Then the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{1,2}(\mathcal{A})$ is $4\sum_{X\in L_{2}(\mathcal{A}),\#\mathcal{A}_{X}=2}c^{X}$ copies of $\mathbb{Z}$.
Similarly, for $\vec{a}=(a_{0},a_{1},a_{2})\in P_{2,2}(\mathcal{A})$, if $a_{0}\neq a_{2}$, then $\vec{a}$ is geodesic and hence $\varphi(\vec{a})\neq\vec{a}$, making the $\vec{a}$-framed magnitude homology trivial. So we only need to consider $\vec{a}=(a_{0},a_{1},a_{0})\in P_{2,2}(\mathcal{A})$, where $a_{0}$ and $a_{1}$ are adjacent. The interval posets $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ and $(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})}$ are empty. Then $MH_{2,(a_{0},a_{1},a_{0})}(\mathcal{A})\cong\widetilde{H}_{-2}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}\otimes C(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})})\cong\mathbb{Z}$. Each such $(a_{0},a_{1},a_{0})$ corresponds to a directed edge of $\mathcal{T}(\mathcal{A})$. Then the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{2,2}(\mathcal{A})$ is $\beta_{1,1}(\mathcal{A})$ copies of $\mathbb{Z}$. Now (3) is proved.
To finish the proof, we need to show other $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are trivial. By Proposition 2.8 (3), we only need to prove $MH_{1,2}(\mathcal{A})=0$. We use Theorem 2.20 again
$MH_{1,2}(A)\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})}MH_{1,\vec{a}}(\mathcal{A}).$
Consider $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$, the interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ cannot be empty. Hence $MH_{1,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{-1}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})=0$. ∎
4.2. Geodesic magnitude homology and diagonal tope graphs
Let us determine the geodesic magnitude homology of arrangements.
Theorem 4.2.
Let $\mathcal{A}$ be an arrangement. Then its geodesic magnitude homology is as follow
$MH_{k,\ell}^{geod}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A}):\mathrm{rank}\mathcal{A}_{F}=k,\#\mathcal{A}_{F}=\ell}\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A}_{F})}$
Proof.
By Corollary 2.22,
$MH_{k,\ell}^{geod}(\mathcal{A})\cong\bigoplus_{a,b\in\mathrm{Ch}(\mathcal{A}):d(a,b)=\ell}\widetilde{H}_{k-2}(C_{*}(a,b)_{\mathcal{T}(\mathcal{A})}).$
By Proposition 2.34, the interval $(a,b)_{\mathcal{T}(\mathcal{A})}$ is homotopy equivalent to $S^{\mathrm{rank}\mathcal{A}_{F}-2}$ if $a,b]_{\mathcal{T}(\mathcal{A})}=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$ and contractible otherwise. Therefore each $F\in\mathcal{F}(\mathcal{A})$ such that $k=\mathrm{rank}\mathcal{A}_{F}$ and $\ell=\#\mathcal{A}_{F}$ contribute to $MH_{k,\ell}^{geod}(\mathcal{A})$ a copy of $\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A})_{F}}$ and there is no other summand. ∎
Recall that $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$ and let $c_{X}=\#\mathrm{Ch}(\mathcal{A}_{X})$. We can give a combinatorial description of geodesic magnitude Betti number.
Corollary 4.3.
The geodesic magnitude Betti number
$\beta_{k,\ell}^{geod}(\mathcal{A})=\sum_{X\in L_{k}(\mathcal{A}):\#\mathcal{A}_{X}=\ell}c^{X}c_{X}.$
In particular, geodesic magnitude homology is determined by $L(\mathcal{A})$.
As an application, we give a characterization of diagonal tope graphs.
Corollary 4.4.
Let $\mathcal{A}$ be an essential arrangement. Then $\mathcal{T}(\mathcal{A})$ is diagonal if and only if $\mathcal{A}$ is Boolean.
Proof.
We know that the Boolean arrangement $Bl_{d}$ has tope graph $\mathcal{T}(Bl_{d})=Q_{d}$, which is diagonal (see Example 2.14). Suppose $\mathcal{A}$ is not Boolean, then $n=\#\mathcal{A}>r=\mathrm{rank}\mathcal{A}$. Therefore
$\beta_{r,n}(\mathcal{A})\geq\beta_{r,n}^{geod}(\mathcal{A})=\#\mathrm{Ch}(\mathcal{A}),$
where the equation is from Corollary 4.3. This means $\mathcal{A}$ is not diagonal. ∎
4.3. Magnitude homology with support
Consider a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A})$, let us define $S(\vec{x})=S(x_{0},x_{1})\cup\cdots\cup S(x_{k-1},x_{k})$ as the set of hyperplanes crossed by $\vec{x}$. Notice that $S(\vec{x})\leq\ell(\vec{x})$. It is not difficult to verify that $x_{i}$ is smooth in $\vec{x}$, i.e. $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$, is equivalent to $S(x_{i-1},x_{i+1})=S(x_{i-1},x_{i})\sqcup S(x_{i},x_{i+1})$. So if a proper chain $\vec{y}$ appears in the boundary of $\vec{x}$, then $S(\vec{x})=S(\vec{y})$. In other words, the boundary operator $\partial:MC_{k,\ell}(\mathcal{A})\to MC_{k-1,\ell}(\mathcal{A})$ preserves crossed hyperplanes. This motivated us to define the following.
Definition 4.5.
For arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$, we define the $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A})$ generated by $P_{k,\ell}(\mathcal{A};\mathcal{B})=\{\vec{x}\in P_{k,\ell}(\mathcal{A})\mid S(\vec{x})=\mathcal{B}\}$. The above discussion shows that $MC_{*,\ell}(\mathcal{A};\mathcal{B})$ forms a subcomplex of $MC_{*,\ell}(\mathcal{A})$, called the $\mathcal{B}$-supported magnitude complex of $\mathcal{A}$. Its homology $MH_{*,\ell}(\mathcal{A};\mathcal{B})$ is called the $\mathcal{B}$-supported magnitude homology of $\mathcal{A}$.
Definition 4.6.
When $\mathcal{B}=\mathcal{A}$, we call $MC_{*,\ell}^{\circ}(\mathcal{A}):=MC_{*,\ell}(\mathcal{A};\mathcal{A})$ the length $\ell$ interior magnitude complex of $\mathcal{A}$ and $MH_{*,\ell}^{\circ}(\mathcal{A}):=MH_{*,\ell}(\mathcal{A};\mathcal{A})$ the length $\ell$ interior magnitude homology of $\mathcal{A}$. Let us also define the interior magnitude Betti number $\beta_{k,\ell}^{\circ}(\mathcal{A}):=\mathrm{rank}MH_{k,\ell}^{\circ}(\mathcal{A})$ and the length $\ell$ interior magnitude Euler characteristic $\chi_{\ell}^{\circ}(\mathcal{A}):=\sum_{k}(-1)^{k}\beta_{k,\ell}^{\circ}(\mathcal{A})$.
The following result is a direct consequence of the above observation that the magnitude boundary operator preserves crossed hyperplanes.
Proposition 4.7.
For arrangement $\mathcal{A}$, we have the decomposition of magnitude chain groups
$MC_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MC_{k,\ell}(\mathcal{A};\mathcal{B}),$
inducing a decomposition of homology groups
$MH_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MH_{k,\ell}(\mathcal{A};\mathcal{B}).$
Furthermore $MC_{k,\ell}(\mathcal{A};\mathcal{B})=0$ and $MH_{k,\ell}(\mathcal{A};\mathcal{B})=0$ if $\ell<\#\mathcal{B}$.
4.4. Face decomposition of magnitude homology
We would like to further decompose the magnitdue complex/homology with support. To this end, we introduce another notation.
Definition 4.8.
For arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$. Consider an induced subgraph $G$ of $\mathcal{T}(\mathcal{A})$. We define the localized $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(G;\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ generated by chains entirely contained in $G$,
$P_{k,\ell}(G;\mathcal{B})=\{(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A};\mathcal{B})\mid x_{0},\ldots,x_{k}\in V(G)\}.$
Lemma 4.9.
For $F\in\mathcal{F}(\mathcal{A})$, the chain groups $MC_{*,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})$ form a subcomplex of $MC_{*,\ell}(\mathcal{A};\mathcal{B})$.
Proof.
We must show that the boundary operator $\partial$ stabilizes the subgroup. By Proposition 4.7, $\partial$ strictly preserves the set $\mathcal{B}$ of crossed hyperplanes. Furthermore, by Lemma 2.26, the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$ is gated, and therefore convex (Proposition 2.2). Because the boundary operator $\partial_{i}$ merely deletes the chamber $x_{i}$ from the sequence, all remaining chambers in the chain $\partial_{i}\vec{x}$ are still contained within the vertex set of $\mathcal{T}(\mathcal{A})_{F}$. Thus, $\partial(MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B}))\subseteq MC_{k-1,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})$. ∎
Now we are ready to state the main structural result of the magnitude homology of an arrangement.
Theorem 4.10.
For an arrangement $\mathcal{A}$, we have the face decomposition of its magnitude complex
$MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC_{k,\ell}^{\circ}(\mathcal{A}_{F}),$
inducing the face decomposition of magnitude homology
$MH_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MH_{k,\ell}^{\circ}(\mathcal{A}_{F}).$
Proof.
By Proposition 4.7, the global magnitude chain complex decomposes into a direct sum over all subarrangements $\mathcal{B}\subseteq\mathcal{A}$,
$MC_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MC_{k,\ell}(\mathcal{A};\mathcal{B}).$
For a fixed subarrangement $\mathcal{B}$, consider a chain $\vec{x}\in P_{k,\ell}(\mathcal{A};\mathcal{B})$. By Definition 4.5, $S(\vec{x})=\mathcal{B}$, meaning the chain never crosses any hyperplane in the complement $\mathcal{A}\setminus\mathcal{B}$. Therefore, every chamber in $\vec{x}$ shares a strictly constant, non-zero sign on all hyperplanes in $\mathcal{A}\setminus\mathcal{B}$. This constant sign pattern uniquely defines a face $F\in\mathcal{F}(\mathcal{A})$ such that $F_{H}=0$ for $H\in\mathcal{B}$, and $F_{H}\in\{+,-\}$ for $H\notin\mathcal{B}$. By definition, the subarrangement associated with this face is exactly $\mathcal{A}_{F}=\mathcal{B}$. Because the sign pattern on $\mathcal{A}\setminus\mathcal{B}$ is uniquely determined for each chain, the subgroup $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ further decomposes into a direct sum over the faces $F$ that satisfy $\mathcal{A}_{F}=\mathcal{B}$,
$MC_{k,\ell}(\mathcal{A};\mathcal{B})=\bigoplus_{F\in\mathcal{F}(\mathcal{A}):\mathcal{A}_{F}=\mathcal{B}}MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B}).$
By Lemma 2.26, the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph isometrically isomorphic to the tope graph of the localized subarrangement $\mathcal{A}_{F}$. Under this isomorphism, the $\mathcal{B}$-supported complex maps identically to the full interior complex of $\mathcal{A}_{F}$. Therefore, by Definition 4.6,
$MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})\cong MC_{k,\ell}(\mathcal{A}_{F};\mathcal{A}_{F})=MC^{\circ}_{k,\ell}(\mathcal{A}_{F}).$
Substituting this isomorphism back into the original sum over subarrangements $\mathcal{B}$ yields the result
$MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}\left(\bigoplus_{F:\mathcal{A}_{F}=\mathcal{B}}MC^{\circ}_{k,\ell}(\mathcal{A}_{F})\right)=\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC^{\circ}_{k,\ell}(\mathcal{A}_{F}).$
Since this is a decomposition into subcomplexes, we have the face decomposition of magnitude homology
$MH_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MH_{k,\ell}^{\circ}(\mathcal{A}_{F}).$
For an arrangement $\mathcal{A}$, let us write
$\chi^{\circ}(\mathcal{A},q):=\sum_{\ell=0}^{\infty}\chi_{\ell}^{\circ}(\mathcal{A})q^{\ell}\in\mathbb{Z}[\![q]\!].$
Then we have a reformulation of Conjecture 3.7.
Corollary 4.11.
Let $r=\mathrm{rank}\mathcal{A}$ and $n=\#\mathcal{A}$. The following are equivalent
  1. (1)
    $\mathrm{Mag}(\mathcal{A},q)=\sum_{F}\mathrm{Mag}^{\circ}(\mathcal{A}_{F},q)$;
  2. (2)
    $\chi^{\circ}(\mathcal{A},q)=\mathrm{Mag}^{\circ}(\mathcal{A},q)$;
  3. (3)
    $\chi_{\ell+n}^{\circ}(\mathcal{A})=(-1)^{r}\chi_{\ell}(\mathcal{A})$ for $\ell\geq 0$.
Proof.
Since Euler characteristic is additive over direct sums, the face decomposition of magnitude homology (Theorem 4.10) yields
$\chi_{\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi_{\ell}^{\circ}(\mathcal{A}_{F}).$
Multiplying by $q^{\ell}$ and summing over $\ell$, we obtain
$\mathrm{Mag}(\mathcal{A},q)=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi^{\circ}(\mathcal{A}_{F},q),$
By Möbius inversion on $\mathcal{F}(\mathcal{A})$, we see that (1) is equivalent to (2). The equivalence between (2) and (3) is straightforward by definition. ∎
Remark 4.12.
Note that if there is a (quasi-)isomorphism of chain complexes
$MC_{*,\ell}(\mathcal{A})\to MC_{*+r,\ell+n}^{\circ}(\mathcal{A})=MC_{*+r,\ell+n}(\mathcal{A};\mathcal{A})$
for all $\ell\geq 0$, then the conditions in Corollary 4.11 hold.
4.5. Magnitude homology on the diagonal
In this subsection, we give a combinatorial formula of the magnitude homology of arrangements on the diagonal $MH_{\ell,\ell}(\mathcal{A})$.
Using Theorem 4.10, we can determine the interior magnitude homology of Boolean arrangements.
Corollary 4.13.
The interior magnitude homology of Boolean arrangement $MH_{k,\ell}^{\circ}(Bl_{d})$ concentrates on the diagonal $k=\ell$ and $MH_{\ell,\ell}^{\circ}(Bl_{d})$ is free abelian of rank
$\beta_{\ell,\ell}^{\circ}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$.
Proof.
We know from Corollary 4.4 and Example 2.14 that $\mathcal{T}(Bl_{d})$ is diagonal and $MH_{\ell,\ell}(Bl_{d})$ is free abelian of rank $\beta_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell+d-1}{d-1}$. By Theorem 4.10, the interior magnitude homology $MH_{k,\ell}^{\circ}(Bl_{d})$ is a direct summand of $MH_{k,\ell}(Bl_{d})$, and hence concentrates on the diagonal $k=\ell$. Theorem 4.10 also gives
$\beta_{\ell,\ell}(Bl_{d})=\sum_{F\in\mathcal{F}(Bl_{d})}\beta_{\ell,\ell}^{\circ}((Bl_{d})_{F}).$
Notice that if $F\in\mathcal{F}(Bl_{d})$ has $\mathrm{codim}F=j$, the localization $(Bl_{d})_{F}$ has essentialization $Bl_{j}$. Then applying Möbius inversion on $\mathcal{F}(\mathcal{A})$ yields the desired formula. ∎
Let us prove the following vanishing result of diagonal interior magnitude homology for non-Boolean arrangements.
Theorem 4.14.
Let $\mathcal{A}$ be an arrangement of $\mathrm{rank}\mathcal{A}=r$ and $\mathcal{A}=n$. Suppose $\mathcal{A}$ is not Boolean $r<n$, then the diagonal interior magnitude homology of $\mathcal{A}$ vanishes for all length
$MH_{\ell,\ell}^{\circ}(\mathcal{A})=0,\forall\ell\geq 0.$
Proof.
Choose an isometric embedding $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{n}=\{+,-\}^{n}$ of $\mathcal{T}(\mathcal{A})$ into the hypercube $Q_{n}$, which is not surjective since $r<n$. Since $\phi$ is distance preserving, it induces injective chain maps $\phi_{\#}:MC_{*,\ell}(\mathcal{A})\hookrightarrow MC_{*,\ell}(Q_{n})$. Let us take $c\in MC_{\ell,\ell}^{\circ}(\mathcal{A})\cap\ker\partial\subseteq MC_{\ell,\ell}(\mathcal{A})$. Since $\partial c=0$, $\phi_{\#}(c)\in MC_{\ell,\ell}(Q_{n})$ is a cycle. For fixed $\ell$, $MC_{\ell+1,\ell}(Q_{n})=0$, by abuse of notation, we say $\phi_{\#}(c)$ is a homology class in $MH_{\ell,\ell}(Q_{n})$. Recall the isomorphism given in Example 2.14
$\nabla:\bigoplus_{\ell_{1}+\cdots+\ell_{n}=\ell}MH_{\ell_{1},\ell_{1}}(Q_{1})\otimes\cdots\otimes MH_{\ell_{n},\ell_{n}}(Q_{1})\xrightarrow{\cong}MH_{\ell,\ell}(Q_{n}),$
is induced by the Eilenberg–Zilber chain map [16]. We may assume there are basis elements $b^{\lambda}$ of the left hand side, one for each summand, indexed by a composition $\lambda=(\ell_{1},\ldots,\ell_{n})$ of $\ell$, such that $\phi_{\#}(c)=\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})$ for some scalars $d_{\lambda}$. We must show that $d_{\lambda}=0$.
Since $c\in MC_{\ell,\ell}^{\circ}(\mathcal{A})$, it is a linear combination of walks that cross all hyperplanes in $\mathcal{A}$. Its image $\phi_{\#}(c)$ in $MC_{\ell,\ell}(Q_{n})$ is a linear combination of walks in $Q_{n}$ along which every coordinate gets flipped at least once. This means the basis elements $b^{\lambda}$ involved are indexed by compositions $\lambda=(\ell_{1},\ldots,\ell_{n})$ with all $\ell_{i}\geq 1$. We write such $b^{\lambda}=b_{1}^{\lambda}\otimes\cdots\otimes b_{n}^{\lambda}$, where $b_{i}^{\lambda}\in MH_{\ell_{i},\ell_{i}}(Q_{1})$ is an $\ell_{i}$-step walk in the $i$-th $Q_{1}$. So $b^{\lambda}$ consists of totally $\ell$ steps in these $n$ dimensions with distribution $\lambda$.
By definition of the Eilenberg-Zilber map, $\nabla(b^{\lambda})$ is the signed sum of all possible shuffles of all these $\ell$ steps. It is a linear combination of walks in $Q_{n}$ from the vertex $(s(b_{1}^{\lambda}),\ldots,s(b_{n}^{\lambda}))$ to the vertex $(t(b_{1}^{\lambda}),\ldots,t(b_{n}^{\lambda}))$, where $s(b_{i}^{\lambda})$ and $t(b_{i}^{\lambda})$ represent the starting vertex and terminating vertex in $Q_{1}$ respectively. Since all $\ell_{i}\geq 1$, any vertex in $Q_{n}$ must be visited by at least one walk in $\nabla(b^{\lambda})$. The next key observation is that if $\lambda\neq\mu$, then $\nabla(b^{\lambda})$ and $\nabla(b^{\mu})$ share no common walk due to distribution difference.
Now recall that $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{n}$ is not surjective, there is a vertex $v_{0}\in Q_{n}$ that is not a vertex of $\mathcal{T}(\mathcal{A})$. We observed that any $\nabla(b^{\lambda})$ contains a walk passing through $v_{0}$ (see Figure 2), but $c\in MC_{\ell,\ell}(\mathcal{A})$ contains no walk passing through $v_{0}$. We must have $d_{\lambda}=0$ in the sum $\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})=\phi_{\#}(c)$. Therefore we conclude $c=0$, and $MH_{\ell,\ell}^{\circ}(\mathcal{A})=0$.
$G^{+}(+,+,+)$$G^{-}(-,-,-)$(+,,)(+,-,-)Start: (+,+,)(+,+,-)(,+,)(-,+,-)(,+,+)(-,+,+)(,,+)(-,-,+)End: (+,,+)(+,-,+)Step 1 ($Z_{1}$)Step 2 ($X_{1}$)Step 3 ($X_{2}$)Step 4 ($Y_{1}$)
Figure 2. The tope graph of the $A_{2}$ arrangement embedded as a $C_{6}$ hexagon (blue/bold) inside the Boolean $Q_{3}$ cube. Suppose $b^{\lambda}=(+\xrightarrow{X_{1}}-\xrightarrow{X_{2}}+)\otimes(+\xrightarrow{Y_{1}}-)\otimes(-\xrightarrow{Z_{1}}+)$, where $\lambda=(2,1,1)$. A 4-step shuffle walk $Z_{1}X_{1}X_{2}Y_{1}$ (orange/dotted) in $\nabla(b^{\lambda})$ is shown hitting the forbidden vertex $G^{+}$. Also notice that this walk cannot appear in other $\nabla(b^{\mu})$.
Remark 4.15.
By Corollary 4.13 and Theorem 4.14, we see that $\mathcal{A}$ is Boolean if and only if $\beta_{\ell,\ell}^{\circ}(\mathcal{A})\neq 0$ for some $\ell\geq 1$.
Corollary 4.16.
Let $\mathcal{A}$ be an arrangement. The diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian of rank $\beta_{\ell,\ell}(\mathcal{A})$, which is determined by the Boolean flats of the intersection lattice $L(\mathcal{A})$. Precisely, for $\ell\geq 1$,
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \mathrm{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\mathrm{rank}X}\binom{\ell-1}{\mathrm{rank}X-1}$.
Consequently, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{max}-1$, where $d_{max}=\max\{\mathrm{rank}X\mid X\in L(\mathcal{A}),\mathrm{rank}X=\#\mathcal{A}_{X}\}$.
Proof.
Since $MC_{\ell+1,\ell}(\mathcal{A})=0$, the diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian. By Theorem 4.10, the global magnitude homology admits a face decomposition. By restricting this decomposition to the diagonal ($k=\ell$) and taking Betti numbers, we obtain
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\beta^{\circ}_{\ell,\ell}(\mathcal{A}_{F}).$
By Theorem 4.14, the interior Betti number $\beta^{\circ}_{\ell,\ell}(\mathcal{A}_{F})$ vanishes unless the localization $\mathcal{A}_{F}$ is Boolean, i.e. $\mathcal{A}_{F}=\mathrm{rank}\mathcal{A}_{F}$. Therefore, the sum collapses to the faces whose support flat $X=s(F)$ satisfies $\mathrm{rank}X=\#\mathcal{A}_{X}$. For these faces, $\mathcal{A}_{F}\cong Bl_{\mathrm{rank}X}$. We can group the non-zero terms of the summation by their support flats $X\in L(\mathcal{A})$. For each valid flat $X$, the number of faces in $\mathcal{A}$ with support $X$ is exactly $c^{X}$. Substituting the known formula for the interior diagonal Betti numbers of the Boolean arrangement (Corollary 4.13), $\beta^{\circ}_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$ where $d=\mathrm{rank}X$, we obtain
$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \mathrm{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\mathrm{rank}X}\binom{\ell-1}{\operatorname{rank}X-1}$.
For $\ell\geq\mathrm{rank}X$, the binomial $\binom{\ell-1}{\mathrm{rank}X-1}$ is a polynomial in $\ell$ of degree $\mathrm{rank}X-1$. Therefore for $\ell$ sufficiently large, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{max}-1$. ∎
Remark 4.17.
Example 2.14 and Proposition 4.1 (2)(3) are special cases of the formula in Corollary 4.16.
5. Open problems
Besides the face decomposition conjecture (Conjecture 3.7, see also Corollary 4.11 and Remark 4.12), we propose the following conjectures based on observations of examples in the appendix.
Conjecture 5.1.
Let $\mathcal{A}$ be an arrangement of rank $r$ and $n=\#\mathcal{A}$. Let $\mathrm{Mag}(\mathcal{A},q)=P(q)/Q(q)$ be the reduced form. Then
  1. (1)
    Suppose $r\geq 3$ is odd. Then $\mathcal{T}(\mathcal{A})$ is not vertex-transitive if and only if $\Phi_{2n}(q)$ divides $Q(q)$.
  2. (2)
    Suppose $r\geq 4$ is even. Then $\mathcal{T}(\mathcal{A})$ is not vertex-transitive if and only if $\Phi_{n}(q)$ divides $Q(q)$.
  3. (3)
    The coefficients of the power series expression of $\mathrm{Mag}(\mathcal{A},q)$ have alternating signs (equivalently $(-1)^{\ell}\chi_{\ell}(\mathcal{A})\geq 0,\forall\ell\geq 0$).
  4. (4)
    $MH_{k,\ell}(\mathcal{A})$ is determined by $L(\mathcal{A})$.
  5. (5)
    $MH_{k,\ell}(\mathcal{A})$ is torsion free.
  6. (6)
    If $\mathcal{A}$ is not Boolean, then $\beta_{0,0}(\mathcal{A})=\beta_{r,n}(\mathcal{A})$.
Acknowledgement
The author thanks Weili Guo and Masahiko Yoshinaga for comments on the preliminary draft of this paper.
Appendix A Computer experiments
The following examples are computed by using SageMath [28].
Example A.1.
We list the results of magnitudes of simplicial arrangements $\mathcal{A}(n,k)$ in Grünbaum’s catalogue [14] for $n\leq 13$. Note that $\mathcal{A}(n,0),\mathcal{A}(6,1),\mathcal{A}(9,1)$ are recorded in Examples 3.10 and 3.11. We use the database of [6] to pull the list of normal vectors.
$\mathrm{Mag}(\mathcal{A}(7,1))$ $=\frac{32q^{4}-40q^{2}+32}{(q^{4}+3q^{3}+4q^{2}+3q+1)(q^{7}+1)}$
$=\frac{8(4q^{4}-5q^{2}+4)}{[2]_{q}^{2}[3]_{q}(q^{7}+1)}$
$=\frac{8(4q^{4}-5q^{2}+4)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{14}(q)}$
$=32-96q+120q^{2}-72q^{3}+24q^{4}-48q^{5}+144q^{6}-272q^{7}+360q^{8}-336q^{9}+240q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(8,1))$ $=\frac{40q^{6}-8q^{4}-16q^{3}-8q^{2}+40}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{8}+1)}$
$=\frac{8(5q^{6}-q^{4}-2q^{3}-q^{2}+5)}{[4]_{q}!(q^{8}+1)}$
$=\frac{8(5q^{6}-q^{4}-2q^{3}-q^{2}+5)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{16}(q)}$
$=40-120q+152q^{2}-112q^{3}+88q^{4}-136q^{5}+240q^{6}-344q^{7}+352q^{8}-248q^{9}+176q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(10,1))$ $=\frac{60q^{8}+60q^{7}-20q^{6}-40q^{5}-40q^{3}-20q^{2}+60q+60}{(q^{8}+4q^{7}+8q^{6}+11q^{5}+12q^{4}+11q^{3}+8q^{2}+4q+1)(q^{10}+1)}$
$=\frac{20(3q^{8}+3q^{7}-q^{6}-2q^{5}-2q^{3}-q^{2}+3q+3)}{[2]_{q}^{2}[3]_{q}[5]_{q}(q^{10}+1)}$
$=\frac{20(3q^{8}+3q^{7}-q^{6}-2q^{5}-2q^{3}-q^{2}+3q+3)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{5}(q)\Phi_{20}(q)}$
$=60-180q+220q^{2}-140q^{3}+60q^{4}-80q^{5}+220q^{6}-380q^{7}+420q^{8}-340q^{9}+220q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(10,2))$ $=\mathrm{Mag}(\mathcal{A}(10,3))=\frac{60q^{6}-12q^{4}-48q^{3}-12q^{2}+60}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{10}+1)}$
$=\frac{12(5q^{6}-2q^{4}-4q^{3}-2q^{2}+5)}{[4]_{q}!(q^{10}+1)}$
$=\frac{12(5q^{6}-2q^{4}-4q^{3}-2q^{2}+5)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)^{2}\Phi_{20}(q)}$
$=60-180q+228q^{2}-192q^{3}+204q^{4}-300q^{5}+432q^{6}-564q^{7}+660q^{8}-672q^{9}+576q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(11,1))$ $=\frac{72q^{6}-16q^{4}-64q^{3}-16q^{2}+72}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{11}+1)}$
$=\frac{8(9q^{6}-2q^{4}-8q^{3}-2q^{2}+9)}{[4]_{q}!(q^{11}+1)}$
$=\frac{8(9q^{6}-2q^{4}-8q^{3}-2q^{2}+9)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{22}(q)}$
$=72-216q+272q^{2}-232q^{3}+256q^{4}-376q^{5}+528q^{6}-680q^{7}+800q^{8}-824q^{9}+784q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(12,1))$ $=\frac{84q^{6}-84q^{5}-36q^{4}+96q^{3}-36q^{2}-84q+84}{(q^{6}+2q^{5}+2q^{4}+2q^{3}+2q^{2}+2q+1)(q^{12}+1)}$
$=\frac{12(7q^{6}-7q^{5}-3q^{4}+8q^{3}-3q^{2}-7q+7)}{[2]_{q}[6]_{q}(q^{12}+1)}$
$=\frac{12(7q^{6}-7q^{5}-3q^{4}+8q^{3}-3q^{2}-7q+7)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{6}(q)\Phi_{8}(q)\Phi_{24}(q)}$
$=84-252q+300q^{2}-168q^{3}+36q^{4}-84q^{5}+336q^{6}-588q^{7}+636q^{8}-504q^{9}+372q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(12,2))$ $=\frac{84q^{10}+84q^{9}+64q^{8}-4q^{7}-44q^{6}-128q^{5}-44q^{4}-4q^{3}+64q^{2}+84q+84}{(q^{10}+4q^{9}+9q^{8}+15q^{7}+20q^{6}+22q^{5}+20q^{4}+15q^{3}+9q^{2}+4q+1)(q^{12}+1)}$
$=\frac{4(21q^{10}+21q^{9}+16q^{8}-q^{7}-11q^{6}-32q^{5}-11q^{4}-q^{3}+16q^{2}+21q+21)}{[5]_{q}!(q^{12}+1)}$
$=\frac{4(21q^{10}+21q^{9}+16q^{8}-q^{7}-11q^{6}-32q^{5}-11q^{4}-q^{3}+16q^{2}+21q+21)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{5}(q)\Phi_{8}(q)\Phi_{24}(q)}$
$=84-252q+316q^{2}-260q^{3}+252q^{4}-344q^{5}+508q^{6}-692q^{7}+804q^{8}-796q^{9}+760q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(12,3))$ $=\frac{84q^{6}-12q^{4}-96q^{3}-12q^{2}+84}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{12}+1)}$
$=\frac{12(7q^{6}-q^{4}-8q^{3}-q^{2}+7)}{[4]_{q}!(q^{12}+1)}$
$=\frac{12(7q^{6}-q^{4}-8q^{3}-q^{2}+7)}{\Phi_{2}(q)^{2}\Phi_{3}(q)\Phi_{4}(q)\Phi_{8}(q)\Phi_{24}(q)}$
$=84-252q+324q^{2}-312q^{3}+396q^{4}-564q^{5}+720q^{6}-876q^{7}+1044q^{8}-1128q^{9}+1116q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,1))$ $=\frac{96q^{8}-96q^{7}+72q^{6}-48q^{5}-48q^{3}+72q^{2}-96q+96}{(q^{8}+2q^{7}+3q^{6}+4q^{5}+4q^{4}+4q^{3}+3q^{2}+2q^{1}+1)(q^{13}+1)}$
$=\frac{24(4q^{8}-4q^{7}+3q^{6}-2q^{5}-2q^{3}+3q^{2}-4q+4)}{[2]_{q}^{2}(1+2q^{2}+2q^{4}+q^{6})(q^{13}+1)}$
$=\frac{24(4q^{8}-4q^{7}+3q^{6}-2q^{5}-2q^{3}+3q^{2}-4q+4)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{6}(q)\Phi_{26}(q)}$
$=96-288q+360q^{2}-288q^{3}+264q^{4}-384q^{5}+624q^{6}-864q^{7}+984q^{8}-960q^{9}+888q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,2))$ $=\frac{96q^{6}-144q^{3}+96}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{13}+1)}$
$=\frac{48(2q^{6}-3q^{3}+2)}{[4]_{q}!(q^{13}+1)}$
$=\frac{48(2q^{6}-3q^{3}+2)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{26}(q)}$
$=96-288q+384q^{2}-432q^{3}+624q^{4}-864q^{5}+1008q^{6}-1152q^{7}+1392q^{8}-1584q^{9}+1632q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,3))$ $=\frac{96q^{10}+96q^{9}+80q^{8}-8q^{7}-64q^{6}-160q^{5}-64q^{4}-8q^{3}+80q^{2}+96q+96}{(q^{10}+4q^{9}+9q^{8}+15q^{7}+20q^{6}+22q^{5}+20q^{4}+15q^{3}+9q^{2}+4q+1)(q^{13}+1)}$
$=\frac{8(12q^{10}+12q^{9}+10q^{8}-q^{7}-8q^{6}-20q^{5}-8q^{4}-q^{3}+10q^{2}+12q+12)}{[5]_{q}!(q^{13}+1)}$
$=\frac{8(12q^{10}+12q^{9}+10q^{8}-q^{7}-8q^{6}-20q^{5}-8q^{4}-q^{3}+10q^{2}+12q+12)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{5}(q)\Phi_{26}(q)}$
$=96-288q+368q^{2}-328q^{3}+336q^{4}-424q^{5}+584q^{6}-784q^{7}+912q^{8}-920q^{9}+920q^{10}+\cdots$
$\mathrm{Mag}(\mathcal{A}(13,4))$ $=\frac{104q^{6}-56q^{4}-48q^{3}-56q^{2}+104}{(q^{6}+3q^{5}+5q^{4}+6q^{3}+5q^{2}+3q+1)(q^{13}+1)}$
$=\frac{8(13q^{6}-7q^{4}-6q^{3}-7q^{2}+13)}{[4]_{q}!(q^{13}+1)}$
$=\frac{8(13q^{6}-7q^{4}-6q^{3}-7q^{2}+13)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{26}(q)}$
$=104-312q+360q^{2}-192q^{3}+72q^{4}-168q^{5}+432q^{6}-696q^{7}+792q^{8}-672q^{9}+504q^{10}+\cdots$
Example A.2.
Consider the rank $3$ generic arrangement $\mathcal{A}=U_{3,4}$ defined by $xyz(x+y+z)$.
$\mathrm{Mag}(\mathcal{A})$ $=\frac{14q^{2}-20q+14}{(q^{2}+2q+1)(q^{4}+1)}$
$=\frac{2(7q^{2}-10q+7)}{[2]_{q}^{2}(q^{4}+1)}$
$=\frac{2(7q^{2}-10q+7)}{\Phi_{2}(q)^{2}\Phi_{8}(q)}$
$=14-48q+96q^{2}-144q^{3}+178q^{4}-192q^{5}+192q^{6}-192q^{7}+206q^{8}-240q^{9}+288q^{10}+\cdots$
Example A.3.
Consider the non-simplicial arrangement $\mathcal{A}$ defined by $xyz(x+z)(y+z)$, which is also the graphic arrangement $\mathcal{A}$ of the graph $K_{4}\setminus e$. Note that $K_{4}\setminus e$ is chordal.
$\mathrm{Mag}(\mathcal{A})$ $=\frac{18q^{4}-2q^{3}-8q^{2}-2q+18}{(q^{4}+3q^{3}+4q^{2}+3q+1)(q^{5}+1)}$
$=\frac{2(9q^{4}-q^{3}-4q^{2}-q+9)}{[2]_{q}^{2}[3]_{q}(q^{5}+1)}$
$=\frac{2(9q^{4}-q^{3}-4q^{2}-q+9)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{10}(q)}$
$=18-56q+88q^{2}-96q^{3}+104q^{4}-154q^{5}+248q^{6}-336q^{7}+376q^{8}-392q^{9}+450q^{10}+\cdots$
Example A.4.
Consider the graphic arrangement $\mathcal{A}$ of the graph $K_{5}\setminus e$.
$\mathrm{Mag}(\mathcal{A})$ $=\frac{12(8q^{8}-9q^{7}+4q^{6}+3q^{5}+3q^{4}+4q^{2}-9q+8)}{[2]_{q}[4]_{q}[9]_{q}(q^{5}+1)}$
$=\frac{12(8q^{8}-9q^{7}+4q^{6}+3q^{5}+3q^{4}+4q^{2}-9q+8)}{\Phi_{2}(q)^{3}\Phi_{3}(q)\Phi_{4}(q)\Phi_{9}(q)\Phi_{10}(q)}$
$=96-396q+756q^{2}-924q^{3}+996q^{4}-1320q^{5}+1956q^{6}-2652q^{7}+3252q^{8}-3756q^{9}+4212q^{10}+\cdots$
where the numerator $P(q)=96q^{8}-108q^{7}+48q^{6}+36q^{5}+36q^{3}+48q^{2}-108q+96$ and the denominator $Q(q)=q^{17}+3q^{16}+5q^{15}+7q^{14}+8q^{13}+9q^{12}+11q^{11}+13q^{10}+15q^{9}+15q^{8}+13q^{7}+11q^{6}+9q^{5}+8q^{4}+7q^{3}+5q^{2}+3q+1$.
Example A.5.
Consider the bracelet arrangement $\mathcal{A}$ defined by $x_{1}x_{2}x_{3}(x_{1}+x_{4})(x_{2}+x_{4})(x_{3}+x_{4})(x_{1}+x_{2}+x_{4})(x_{1}+x_{3}+x_{4})(x_{2}+x_{3}+x_{4})$ (see Example 1.4 of [7]),
$\mathrm{Mag}(\mathcal{A})$ $=\frac{6(17q^{8}-4q^{7}-8q^{6}-9q^{5}+44q^{4}-9q^{3}-8q^{2}-4q+17)}{\Phi_{2}(q)^{3}\Phi_{3}(q)^{2}\Phi_{9}(q)\Phi_{10}(q)}$
$=102-432q+864q^{2}-1176q^{3}+1584q^{4}-2628q^{5}+4344q^{6}-6048q^{7}+7200q^{8}-8130q^{9}+9684q^{10}+\cdots$
where the numerator $P(q)=102q^{8}-24q^{7}-48q^{6}-54q^{5}+264q^{4}-54q^{3}-48q^{2}-24q+102$ and the denominator $Q(q)=q^{17}+4q^{16}+8q^{15}+11q^{14}+12q^{13}+13q^{12}+16q^{11}+20q^{10}+23q^{9}+23q^{8}+20q^{7}+16q^{6}+13q^{5}+12q^{4}+11q^{3}+8q^{2}+4q+1$.
Example A.6.
We collect here some results of magnitude Betti numbers using SageMath codes of [16].
[…]
0 4 0 0 0 0 0 0 0 0
1 0 8 0 0 0 0 0 0 0
2 0 0 12 0 0 0 0 0 0
3 0 0 0 16 0 0 0 0 0
4 0 0 0 0 20 0 0 0 0
5 0 0 0 0 0 24 0 0 0
6 0 0 0 0 0 0 28 0 0
7 0 0 0 0 0 0 0 32 0
8 0 0 0 0 0 0 0 0 36
Table 1. Magnitude Betti numbers $\beta_{k,l}$ for the Boolean arrangement $Bl_{2}=U_{2,2}$ defined by $xy=0$.
[…]
0 6 0 0 0 0 0 0 0 0
1 0 12 0 0 0 0 0 0 0
2 0 0 12 6 0 0 0 0 0
3 0 0 0 12 12 0 0 0 0
4 0 0 0 0 12 12 6 0 0
5 0 0 0 0 0 12 12 12 0
6 0 0 0 0 0 0 12 12 12
7 0 0 0 0 0 0 0 12 12
8 0 0 0 0 0 0 0 0 12
Table 2. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{2}=U_{2,3}$.
[…]
0 8 0 0 0 0 0 0 0 0
1 0 16 0 0 0 0 0 0 0
2 0 0 16 0 8 0 0 0 0
3 0 0 0 16 0 16 0 0 0
4 0 0 0 0 16 0 16 0 8
5 0 0 0 0 0 16 0 16 0
6 0 0 0 0 0 0 16 0 16
7 0 0 0 0 0 0 0 16 0
8 0 0 0 0 0 0 0 0 16
Table 3. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $B_{2}=U_{2,4}$.
[…]
0 24 0 0 0 0 0 0 0 0
1 0 72 0 0 0 0 0 0 0
2 0 0 96 48 0 0 0 0 0
3 0 0 0 120 96 0 24 0 0
4 0 0 0 0 144 96 48 72 0
5 0 0 0 0 0 168 96 96 96
6 0 0 0 0 0 0 192 96 96
7 0 0 0 0 0 0 0 216 96
8 0 0 0 0 0 0 0 0 240
Table 4. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{3}$.
[…]
0 120 0 0 0 0 0 0
1 0 480 0 0 0 0 0
2 0 0 840 360 0 0 0
3 0 0 0 1200 960 0 240
4 0 0 0 0 1560 1440 360
5 0 0 0 0 0 1920 1920
6 0 0 0 0 0 0 2280
Table 5. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{4}$.
[…]
0 14 0 0 0 0 0 0 0 0
1 0 48 0 0 0 0 0 0 0
2 0 0 96 0 0 0 0 0 0
3 0 0 0 144 14 0 0 0 0
4 0 0 0 0 192 48 0 0 0
5 0 0 0 0 0 240 96 0 0
6 0 0 0 0 0 0 288 144 14
7 0 0 0 0 0 0 0 336 192
8 0 0 0 0 0 0 0 0 384
Table 6. Magnitude Betti numbers $\beta_{k,l}$ for the generic arrangement $U_{3,4}$ defined by $xyz(x+y+z)$ Example A.2.
[…]
0 30 0 0 0 0 0 0
1 0 140 0 0 0 0 0
2 0 0 380 0 0 0 0
3 0 0 0 780 0 0 0
4 0 0 0 0 1340 30 0
5 0 0 0 0 0 2060 140
6 0 0 0 0 0 0 2940
Table 7. Magnitude Betti numbers $\beta_{k,l}$ for the generic arrangement $U_{4,5}$.
[…]
0 18 0 0 0 0 0 0 0 0
1 0 56 0 0 0 0 0 0 0
2 0 0 88 24 0 0 0 0 0
3 0 0 0 120 48 18 0 0 0
4 0 0 0 0 152 48 80 0 0
5 0 0 0 0 0 184 48 136 24
6 0 0 0 0 0 0 216 48 168
7 0 0 0 0 0 0 0 248 48
8 0 0 0 0 0 0 0 0 280
Table 8. Magnitude Betti numbers $\beta_{k,l}$ for the graphic arrangement of $K_{4}\setminus e$ Example A.3
[…]
0 32 0 0 0 0 0 0 0 0
1 0 96 0 0 0 0 0 0 0
2 0 0 120 72 0 0 0 0 0
3 0 0 0 144 144 0 0 32 0
4 0 0 0 0 168 144 72 0 96
5 0 0 0 0 0 192 144 144 0
6 0 0 0 0 0 0 216 144 144
7 0 0 0 0 0 0 0 240 144
8 0 0 0 0 0 0 0 0 264
Table 9. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(7,1)$
[…]
0 40 0 0 0 0 0 0 0 0
1 0 120 0 0 0 0 0 0 0
2 0 0 152 72 16 0 0 0 0
3 0 0 0 184 144 32 0 0 40
4 0 0 0 0 216 144 104 0 16
5 0 0 0 0 0 248 144 176 0
6 0 0 0 0 0 0 280 144 176
7 0 0 0 0 0 0 0 312 144
8 0 0 0 0 0 0 0 0 344
Table 10. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(8,1)$
[…]
0 48 0 0 0 0 0 0 0 0
1 0 144 0 0 0 0 0 0 0
2 0 0 192 48 48 0 0 0 0
3 0 0 0 240 96 96 0 0 0
4 0 0 0 0 288 96 144 0 48
5 0 0 0 0 0 336 96 192 0
6 0 0 0 0 0 0 384 96 192
7 0 0 0 0 0 0 0 432 96
8 0 0 0 0 0 0 0 0 480
Table 11. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(9,1)$
Reference
  • [1] T. Abe, T. Horiguchi, M. Masuda, S. Murai, and T. Sato (2020) Hessenberg varieties and hyperplane arrangements. J. Reine Angew. Math. 764, pp. 241–286. External Links: ISSN 0075-4102,1435-5345, Document, Link, MathReview (Marko Radovanović) Cited by: §2.7.
  • [2] M. Aguiar and S. Mahajan (2017) Topics in hyperplane arrangements. Mathematical Surveys and Monographs, Vol. 226, American Mathematical Society, Providence, RI. External Links: ISBN 978-1-4704-3711-4, Document, Link, MathReview (Tom C. Braden) Cited by: §2.6, Proposition 2.2, Remark 2.25, Theorem 2.33.
  • [3] A. Björner, P. H. Edelman, and G. M. Ziegler (1990) Hyperplane arrangements with a lattice of regions. Discrete Comput. Geom. 5 (3), pp. 263–288. External Links: ISSN 0179-5376,1432-0444, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.7, Proposition 2.36.
  • [4] A. Björner, M. Las Vergnas, B. Sturmfels, N. White, and G. M. Ziegler (1999) Oriented matroids. Second edition, Encyclopedia of Mathematics and its Applications, Vol. 46, Cambridge University Press, Cambridge. External Links: ISBN 0-521-77750-X, Document, Link, MathReview Entry Cited by: Theorem 1.1, Proposition 2.30, Proposition 2.34, Corollary 2.35.
  • [5] S. Brenner, J. Cardinal, T. McConville, A. Merino, and T. Mütze (2026) Combinatorial generation via permutation languages. VII. Supersolvable hyperplane arrangements. European J. Combin. 135, pp. Paper No. 104367. External Links: ISSN 0195-6698,1095-9971, Document, Link, MathReview Entry Cited by: §2.5.
  • [6] M. Cuntz, S. Elia, and J. Labbé (2022) Congruence normality of simplicial hyperplane arrangements via oriented matroids. Ann. Comb. 26 (1), pp. 1–85. External Links: ISSN 0218-0006,0219-3094, Document, Link, MathReview (Hery Randriamaro) Cited by: Example A.1.
  • [7] G. Denham, G. Dorpalen-Barry, and N. Proudfoot (2026) Cleanliness and the varchenko-gelfand algebra. arXiv preprint arXiv:2512.10077v2. Cited by: Example A.5.
  • [8] G. Denham and P. Hanlon (1999) Some algebraic properties of the Schechtman-Varchenko bilinear forms. In New perspectives in algebraic combinatorics (Berkeley, CA, 1996–97), Math. Sci. Res. Inst. Publ., Vol. 38, pp. 149–176. External Links: ISBN 0-521-77087-4, MathReview (P. Orlik) Cited by: §2.6.
  • [9] R. Diestel ([2025] ©2025) Graph theory. Sixth edition, Graduate Texts in Mathematics, Vol. 173, Springer, Berlin. External Links: ISBN 978-3-662-70106-5; 978-3-662-70107-2, MathReview Entry Cited by: §2.5.
  • [10] P. H. Edelman and J. W. Walker (1985) The homotopy type of hyperplane posets. Proc. Amer. Math. Soc. 94 (2), pp. 221–225. External Links: ISSN 0002-9939,1088-6826, Document, Link, MathReview (Udo Pachner) Cited by: Proposition 2.34.
  • [11] P. H. Edelman (1984) A partial order on the regions of ${\bf R}^{n}$ dissected by hyperplanes. Trans. Amer. Math. Soc. 283 (2), pp. 617–631. External Links: ISSN 0002-9947,1088-6850, Document, Link, MathReview (Robert A. Proctor) Cited by: Corollary 2.35.
  • [12] L. Giordani, T. Möller, P. Mücksch, and G. Roehrle (2025) On connected subgraph arrangements. arXiv preprint arXiv:2502.18144v1. Cited by: §2.7.
  • [13] K. Gomi (2025) Magnitude homology of geodesic space. arXiv preprint arXiv:1902.07044v3. Cited by: Proposition 2.21.
  • [14] B. Grünbaum (2009) A catalogue of simplicial arrangements in the real projective plane. Ars Math. Contemp. 2 (1), pp. 1–25. External Links: ISSN 1855-3966,1855-3974, Document, Link, MathReview Entry Cited by: Example A.1, §1, §3.3.
  • [15] Y. Gu (2018) Graph magnitude homology via algebraic morse theory. arXiv preprint arXiv:1809.07240v1. Cited by: Example 2.16.
  • [16] R. Hepworth and S. Willerton (2017) Categorifying the magnitude of a graph. Homology Homotopy Appl. 19 (2), pp. 31–60. External Links: ISSN 1532-0073,1532-0081, Document, Link, MathReview (Haimiao Chen) Cited by: Example A.6, §1, §2.3, Proposition 2.10, Theorem 2.12, Theorem 2.12, Proposition 2.8, Proposition 2.9, §4.5.
  • [17] W. Hochstättler and V. Welker (2019) The Varchenko determinant for oriented matroids. Math. Z. 293 (3-4), pp. 1415–1430. External Links: ISSN 0025-5874,1432-1823, Document, Link, MathReview (Daryl Funk) Cited by: §2.6.
  • [18] M. Jambu and L. Paris (1995) Combinatorics of inductively factored arrangements. European J. Combin. 16 (3), pp. 267–292. External Links: ISSN 0195-6698,1095-9971, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.7.
  • [19] R. Kaneta and M. Yoshinaga (2021) Magnitude homology of metric spaces and order complexes. Bull. Lond. Math. Soc. 53 (3), pp. 893–905. External Links: ISSN 0024-6093,1469-2120, Document, Link, MathReview (David Matthew Freeman) Cited by: §2.3, §2.3, Theorem 2.20, Theorem 2.23.
  • [20] V. Körber, T. Schnieders, J. Stricker, and J. Walizadeh (2025) Hamiltonian cycles in simplicial and supersolvable hyperplane arrangements. arXiv preprint arXiv:2508.14538v2. Cited by: §2.5.
  • [21] T. Leinster and M. W. Meckes (2017) The magnitude of a metric space: from category theory to geometric measure theory. In Measure theory in non-smooth spaces, Partial Differ. Equ. Meas. Theory, pp. 156–193. External Links: ISBN 978-3-11-055082-5; 978-3-11-055083-2, MathReview (Tobias W. Fritz) Cited by: §1.
  • [22] T. Leinster and M. Meckes (2023) Spaces of extremal magnitude. Proc. Amer. Math. Soc. 151 (9), pp. 3967–3973. External Links: ISSN 0002-9939,1088-6826, Document, Link, MathReview (Shlomo Reisner) Cited by: §3.1.
  • [23] T. Leinster and M. Shulman (2021) Magnitude homology of enriched categories and metric spaces. Algebr. Geom. Topol. 21 (5), pp. 2175–2221. External Links: ISSN 1472-2747,1472-2739, Document, Link, MathReview (Ahmet A. Husainov) Cited by: §1.
  • [24] T. Leinster (2013) The magnitude of metric spaces. Doc. Math. 18, pp. 857–905. External Links: ISSN 1431-0635,1431-0643, MathReview (J. Böhm) Cited by: §1.
  • [25] T. Leinster (2019) The magnitude of a graph. Math. Proc. Cambridge Philos. Soc. 166 (2), pp. 247–264. External Links: ISSN 0305-0041,1469-8064, Document, Link, MathReview (Lorenzo Traldi) Cited by: §1, §2.2, Proposition 2.3, Proposition 2.5, Theorem 2.6, Proposition 2.7.
  • [26] S. Oh, A. Postnikov, and H. Yoo (2008) Bruhat order, smooth Schubert varieties, and hyperplane arrangements. J. Combin. Theory Ser. A 115 (7), pp. 1156–1166. External Links: ISSN 0097-3165,1096-0899, Document, Link, MathReview (David E. Anderson) Cited by: §2.7.
  • [27] P. Orlik and H. Terao (1992) Arrangements of hyperplanes. Grundlehren der mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences], Vol. 300, Springer-Verlag, Berlin. External Links: ISBN 3-540-55259-6, Document, Link, MathReview (Michel Yves Jambu) Cited by: §2.4.
  • [28] The Sage Developers (2025) Sagemath, the Sage Mathematics Software System (Version 10.8). Note: https://www.sagemath.org Cited by: Appendix A.
  • [29] L. Solomon (1966) The orders of the finite Chevalley groups. J. Algebra 3, pp. 376–393. External Links: ISSN 0021-8693, Document, Link, MathReview (T. Ono) Cited by: Proposition 2.37.
  • [30] R. P. Stanley (2007) An introduction to hyperplane arrangements. In Geometric combinatorics, IAS/Park City Math. Ser., Vol. 13, pp. 389–496. External Links: ISBN 978-0-8218-3736-8; 0-8218-3736-2, Document, Link, MathReview Entry Cited by: §2.4, Remark 2.25.
  • [31] A. Varchenko (1993) Bilinear form of real configuration of hyperplanes. Adv. Math. 97 (1), pp. 110–144. External Links: ISSN 0001-8708,1090-2082, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.6, Theorem 2.33.
  • [32] Y. Yagi and M. Yoshinaga (2026) Reconstrucion of oriented matroids from varchenko-gelfand algebras. arXiv preprint arXiv:2509.19905v3. Cited by: §1.
  • [33] T. Zaslavsky (1975) Facing up to arrangements: face-count formulas for partitions of space by hyperplanes. Mem. Amer. Math. Soc. 1, pp. vii+102. External Links: ISSN 0065-9266,1947-6221, Document, Link, MathReview (E. Jucovič) Cited by: §2.4, Remark 2.25.

### 01 May 2026 v3

@@
 Abstract.
-We initiate the study of magnitude theory of real hyperplane arrangements.
-Magnitude is a cardinality-like invariant of metric spaces or enriched categories measuring the effective size.
-Its categorification, the magnitude homology, is a more powerful invariant.
-For a real hyperplane arrangement, or more generally, an oriented matroid, the tope graph encapsulates considerable amount of information.
-Since tope graphs are equipped with the shortest path metric, we feed them to the magnitude and magnitude homology machinery to derive new invariants of real hyperplane arrangements.
-We prove some structural results of the magnitude of arrangements, including reciprocity, palindromic numerator and denominator.
-For magnitude homology of arrangements, we give combinatorial descriptions in small length and prove that tope graphs are diagonal if and only if the arrangement is Boolean.
-We present a face decomposition of magnitude homology, using which we obtain a combinatorial formula of diagonal magnitude Betti numbers.
-Many open problems are posted for future study.
-In particular, we conjecture that magnitude and magnitude homology of arrangements are determined by the intersection lattice.
+We define and study the magnitude and magnitude homology of a real hyperplane arrangement by regarding its tope graph as a metric space.
+We prove several structural results for the magnitude of arrangements, including a symmetry formula, palindromicity of the numerator and denominator, a face decomposition formula, and results on the sign pattern of the magnitude power series.
+For the magnitude homology of arrangements, we obtain combinatorial formulas for small lengths and show that it detects Boolean arrangements.
+We also lift the face decomposition formula to a homological decomposition and derive explicit formulas for the diagonal magnitude Betti numbers.
+Another notable feature is that the magnitude Euler characteristic satisfies a reciprocity theorem analogous to Ehrhart–Macdonald reciprocity.
+We conclude by presenting several conjectures.
+In particular, we conjecture that the magnitude homology of an arrangement is torsion-free and is determined by the intersection lattice.
 Key words and phrases:
@@
 1. Introduction
-Let $\mathcal{A}$ be an arrangement of hyperplanes in $\mathbb{R}^{\ell}$ all passing through the origin.
-The hyperplanes in $\mathcal{A}$ cut $\mathbb{R}^{\ell}$ into a disjoint union of chambers.
-Let $\mathrm{Ch}(\mathcal{A})$ denote the set of chambers.
-For two chambers $C,C^{\prime}\in\mathrm{Ch}(\mathcal{A})$, we write $S(C,C^{\prime})$ as the set of hyperplanes in $\mathcal{A}$ that separate $C$ and $C^{\prime}$.
-The tope graph $\mathcal{T}(\mathcal{A})$ of $\mathcal{A}$ is the graph with vertex set $\mathrm{Ch}(\mathcal{A})$, and two vertices $C$ and $C^{\prime}$ are connected by an edge if $S(C,C^{\prime})$ consists of a single hyperplane.
-The graph metric $d$ on $\mathcal{T}(\mathcal{A})$ satisfies $d(C,C^{\prime})=\#S(C,C^{\prime})$.
-Note that $\mathcal{T}(\mathcal{A})$ is a partial cube, that is, isometric to a subgraph of the hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}$ equipped with the Hamming distance.
-For an oriented matroid $\mathcal{M}$, the tope graph $\mathcal{T}(\mathcal{M})$ can be defined similarly.
-Although seemingly elementary, the tope graphs turn out to be quite informative.
+Let $\mathcal{A}$ be an arrangement of hyperplanes passing through the origin in $\mathbb{R}^{d}$.
+The hyperplanes of $\mathcal{A}$ divide $\mathbb{R}^{d}$ into regions called chambers.
+We denote the set of chambers by $\mathrm{Ch}(\mathcal{A})$.
+The tope graph $\mathcal{T}(\mathcal{A})$ of $\mathcal{A}$ is the graph with vertex set $\mathrm{Ch}(\mathcal{A})$ and with edges joining two chambers separated by a single hyperplane.
+More generally, the tope graph is defined for an oriented matroid.
+Although its definition is elementary, the tope graph contains essentially all combinatorial information about a hyperplane arrangement:
 Theorem 1.1 (Theorem 4.2.14 of [4]).
@@
 We refer to Figure 1 of [34] for a detailed diagram of implications.
-In particular, if two arrangements have isomorphic tope graphs, then they have homoemorphic complexified complements, isomorphic intersection lattices and isomorphic filtered Varchenko–Gelfand algebras.
+In particular, if two arrangements have isomorphic tope graphs, then they have homeomorphic complexified complements, isomorphic intersection lattices, and isomorphic filtered Varchenko–Gelfand algebras.
 Hence we may study $\mathcal{A}$ via $\mathcal{T}(\mathcal{A})$.
-In this paper, we feed tope graphs into the emerging machinery of magnitude theory.
-Originally introduced by Leinster for enriched categories and metric spaces [24, 21], magnitude acts as a cardinality-like invariant measuring the “effective size” of a space.
-For finite graphs [25], it is a rational function defined via the inverse of the $q$-distance matrix, which is intimately related to the classical Varchenko matrix in the context of arrangements.
-The categorification of this concept, magnitude homology, developed by Hepworth and Willerton for graphs [16], and by Leinster and Shulman for enriched categories and metric spaces [23], provides a powerful bigraded homology theory that detects profound metric features.
-We initiate the study of magnitude and magnitude homology specifically tailored to the metric geometry of real hyperplane arrangements.
-Our motivation is two-fold.
-The first is that we introduce a new rational function invariant and a new bigraded homology invariant of real arrangements.
-As a special class of graphs, tope graphs of arrangements (or oriented matroids) possess many unique features, such as antipodal symmetry and organized face decomposition, that enable us to contribute to the theory of magnitude with new beautiful examples.
-At the decategorified level, we demonstrate that the magnitude of an arrangement, $\mathrm{Mag}(\mathcal{A},q)$, exhibits remarkable structural results.
-We prove a reciprocity formula relating $\mathrm{Mag}(\mathcal{A},q)$ and $\mathrm{Mag}(\mathcal{A},q^{-1})$, and establish that its reduced rational form consists of palindromic numerator and denominator whose poles are roots of unity.
-Supported by extensive computational evidence generated via SageMath on the Grünbaum catalogue [14], we further propose a conjectural Face Decomposition formula, which serves as a discrete $q$-analogue to the classical decomposition of the Euler characteristic of a zonotope.
-At the homological level, by exploiting the metric structure of gated subgraphs within partial cubes, which naturally arise from the faces of the arrangement, we systematically dissect the magnitude homology of $\mathcal{T}(\mathcal{A})$.
-After providing explicit combinatorial descriptions of the homology groups for small lengths, we establish a strict topological dichotomy for arrangements.
-We prove that the tope graph is diagonal, meaning its off-diagonal magnitude homology strictly vanishes, if and only if the arrangement is Boolean.
-We also prove a face decomposition of magnitude homology of arrangements, and as an application, we derive a formula of diagonal magnitude Betti numbers in terms of the intersection lattice.
+In this paper, by applying the emerging theory of magnitude to tope graphs, we define the magnitude and magnitude homology of arrangements and initiate the study of these invariants.
+Magnitude is a cardinality-like invariant of enriched categories and metric spaces, defined by Leinster [24, 21]; for metric spaces, it can be interpreted as measuring their “effective size.” The magnitude of a graph is defined by regarding the graph as a metric space via the shortest-path metric [25].
+The magnitude of a graph is a rational function defined using the inverse of the $q$-distance matrix; in the case of hyperplane arrangements, this matrix coincides with the classical Varchenko matrix.
+The notion of magnitude homology, a categorification of magnitude, was introduced for graphs by Hepworth and Willerton [17], and for enriched categories and metric spaces by Leinster and Shulman [23].
+It is a bigraded homology theory that reflects deep metric properties.
+Our motivation is twofold.
+First, magnitude and magnitude homology give new invariants of real hyperplane arrangements.
+What information about an arrangement is reflected by these new invariants remains mysterious and is worth exploring.
+Second, the tope graphs of arrangements form a class of graphs with many distinctive features, such as antipodal symmetry and decompositions into faces, and may therefore serve as a source of interesting examples in the magnitude theory of graphs.
+At the decategorified level, we show that the magnitude of an arrangement, $\mathrm{Mag}(\mathcal{A},q)$, satisfies remarkable structural properties.
+We prove a symmetry relating $\mathrm{Mag}(\mathcal{A},q)$ and $\mathrm{Mag}(\mathcal{A},q^{-1})$, which reflects the antipodal symmetry of the tope graph.
+The reduced rational form of $\mathrm{Mag}(\mathcal{A},q)$ has palindromic numerator and denominator, and all its poles are roots of unity.
+Moreover, if we define the interior magnitude by $\mathrm{Mag}^{\circ}(\mathcal{A},q)=(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}\mathrm{Mag}(\mathcal{A},q)$, then the following face decomposition formula holds:
+$\mathrm{Mag}(\mathcal{A},q)=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F},q).$
+Here $\mathcal{F}(\mathcal{A})$ denotes the set of faces (covectors) of $\mathcal{A}$, and $\mathcal{A}_{F}=\{H\in\mathcal{A}\mid F\subset H\}$.
+This serves as a discrete $q$-analogue of the classical decomposition of the Euler characteristic of a zonotope.
+We then investigate the sign pattern of the magnitude power series.
+In particular, we show that for arrangements of rank at most $3$, the signs of the magnitude coefficients eventually alternate.
+We also present an example with non-alternating magnitude.
+The magnitude homology $MH_{k,\ell}(\mathcal{A})$ of an arrangement $\mathcal{A}$ is a more refined invariant.
+By analyzing the structure of gated subgraphs arising from the faces of an arrangement, we establish the basic properties of this invariant.
+First, we describe all magnitude homology groups of length $\ell\leq 2$ in terms of the combinatorics of the arrangement.
+Next, we show that magnitude homology detects Boolean arrangements: an arrangement is Boolean if and only if its magnitude homology is concentrated on the diagonal.
+Furthermore, we define the interior magnitude homology $MH^{\circ}_{k,\ell}(\mathcal{A})$, which is a categorification of the interior magnitude $\mathrm{Mag}^{\circ}(\mathcal{A})$, and lift the face decomposition formula for magnitude to the homological level.
+In particular, we obtain the Euler characteristic reciprocity
+$\chi^{\circ}_{\ell+\#\mathcal{A}}(\mathcal{A})=(-1)^{\mathrm{rank}\mathcal{A}}\chi_{\ell}(\mathcal{A}),$
+where $\chi_{\ell}(\mathcal{A})$ (resp.
+$\chi^{\circ}_{\ell}(\mathcal{A})$) is the Euler characteristic of $MH_{*,\ell}(\mathcal{A})$ (resp.
+$MH^{\circ}_{*,\ell}(\mathcal{A})$).
+This can be regarded as a new combinatorial reciprocity theorem analogous to the Ehrhart–Macdonald reciprocity theorem for lattice polytopes [11, 26] and Stanley’s reciprocity theorem for posets [31].
+As an application of the homological face decomposition, we give a complete description of the diagonal magnitude Betti numbers in terms of the intersection lattice.
 The paper is organized as follows.
-Section 2 collects necessary recollections from metric graph theory and the theory of hyperplane arrangements, establishing the foundational link between gated subgraphs and covector composition.
-Section 3 investigates the basic properties and structural rigidity of the magnitude rational function.
-Section 4 explores magnitude homology, presenting the Boolean diagonality theorem and face decomposition theorem.
-Finally, Section 5 posts several open problems and conjectures for future study.
-Results of many computer experiments with SageMath form the appendix.
-Convention: When we talk about a topological property/construction of a poset, we mean that of its order complex.
+Section 2 recalls the necessary material from metric graph theory and the theory of hyperplane arrangements, establishing the foundational link between gated subgraphs and covector composition.
+Section 3 investigates the basic properties and structure of the magnitude rational function.
+Section 4 focuses on the sign pattern of magnitude power series.
+Section 5 explores magnitude homology, presenting the Boolean diagonality theorem and the face decomposition theorem.
+Finally, Section 6 poses several open problems and conjectures for future study.
+The appendix contains results from numerous SageMath computer experiments.
+Use of AI
+In writing this paper, we used AI in the following ways.
+  • A substantial part of the proof strategy for Theorem 3.7 was generated by ChatGPT-5.4 Pro.
+  • Part of the proof strategy for Theorem 4.5 was generated by ChatGPT-5.5 Pro.
+  • We used ChatGPT-5.5 Pro to proofread the manuscript for minor errors.
+We emphasize, however, that we have not used any AI-generated text without careful verification. The text of this paper has been written under the full responsibility of the authors.
+Convention
+In this paper, all graphs considered are finite, simple, and undirected unless otherwise specified.
+When we discuss a topological property or construction of a poset, we mean that of its order complex.
+We shall use the term “arrangement” to mean a real central hyperplane arrangement.
 2. Recollections
 2.1. Metric graph theory
-In this paper, all graphs we mention are supposed to be finite simple undirected graphs unless otherwise specified.
-For a graph $G=(V,E)$, a walk (or gallery) of length $k$ from vertex $x\in V$ to vertex $y\in V$ is a sequence of vertices $(x_{0},x_{1},\ldots,x_{k})$ such that $x_{0}=x$, $x_{k}=y$ and $\{x_{i},x_{i+1}\}\in E$ for $i=0,1,\ldots,k-1$.
-We define a distance function $d:V\times V\to\mathbb{Z}\cup\{\infty\}$ by declaring $d(x,y)$ is the minimal length of walks from $x$ to $y$.
+For a graph $G=(V,E)$, a walk (or gallery) of length $k$ from a vertex $x\in V$ to a vertex $y\in V$ is a sequence of vertices $(x_{0},x_{1},\ldots,x_{k})$ such that $x_{0}=x$, $x_{k}=y$, and $\{x_{i},x_{i+1}\}\in E$ for $i=0,1,\ldots,k-1$.
+We define a distance function $d:V\times V\to\mathbb{Z}\cup\{\infty\}$ by declaring $d(x,y)$ to be the minimum length among walks from $x$ to $y$.
 If there is no walk from $x$ to $y$, we set $d(x,y)=\infty$.
 Then $(G,d)$ is a generalized metric space.
-In particular, when $G$ is connected, that is there is a walk between any pair of vertices, $(G,d)$ is a metric space.
-We call the maximal value of $d$ the diameter of $G$, denoted by $\mathrm{diam}(G)$.
+In particular, when $G$ is connected, that is, when there is a walk between any pair of vertices, $(G,d)$ is a metric space.
+We call the maximum value of $d$ the diameter of $G$, denoted by $\operatorname{diam}(G)$.
 A subgraph $H$ of $G$ can be equipped with a metric in two ways.
-One is to restrict $d_{G}$ to $H$ and the other is $d_{H}$ on $H$ as a graph itself.
-We say $H$ is an isometric subgraph of $G$ if the two metrics coincide $d_{H}(x,y)=d_{G}(x,y)$ for $x,y\in V(H)$.
+One is the restriction of $d_{G}$ to $H$, and the other is the graph metric $d_{H}$ on $H$ itself.
+We say that $H$ is an isometric subgraph of $G$ if the two metrics coincide, that is, if $d_{H}(x,y)=d_{G}(x,y)$ for $x,y\in V(H)$.
 In a graph $G$, a walk from $x$ to $y$ of minimal length $d(x,y)$ is called a minimal walk.
-Given $x,y\in V$, let us define the closed interval $x,y]_{G}$ as the set of all vertices traversed by a minimal walk from $x$ to $y$.
-It is a poset equipped with partial order $u\prec_{G}v$ if there is a minimal walk from $x$ to $y$ traversing $u$ before traversing $v$.
+For $x,y\in V$, define the closed interval $x,y]_{G}$ as the set of all vertices traversed by a minimal walk from $x$ to $y$.
+We equip it with the partial order $u\prec_{G}v$ if there is a minimal walk from $x$ to $y$ that traverses $u$ before $v$.
 Note that $x,y]_{G}$ has a unique minimal element $x$ and a unique maximal element $y$.
 A vertex $u$ belongs to $x,y]_{G}$ if and only if $d(x,y)=d(x,u)+d(u,y)$.
-More generally, vertices $u_{1},\ldots,u_{k}$ form a strict linearly ordered chain in $x,y]_{G}$, that is $x\prec_{G}u_{1}\prec_{G}\cdots\prec_{G}u_{k}\prec_{G}y$ if and only if
-$d(x,y)=d(x,u_{1})+d(u_{1},u_{2})+\cdots+d(u_{k-1},u_{k})+d(u_{k},y).$
-We define the open interval $(x,y)_{G}$ as the subposet of $x,y]_{G}$ excluding $x$ and $y$.
-We say a subgraph $H$ of $G$ is convex in $G$ if for any $x,y\in V(H)$, any minimal walk from $x$ to $y$ traverses only vertices in $H$, equivalently $x,y]_{G}\subseteq V(H)$ for $x,y\in V(H)$.
-Apparently, any convex subgraph is an isometric subgraph, but the converse is false.
+More generally, vertices $u_{1},\ldots,u_{k}$ form a strictly linearly ordered chain in $x,y]_{G}$, that is, $x\prec_{G}u_{1}\prec_{G}\cdots\prec_{G}u_{k}\prec_{G}y$, if and only if
+(2.1) $d(x,y)=d(x,u_{1})+d(u_{1},u_{2})+\cdots+d(u_{k-1},u_{k})+d(u_{k},y).$
+We define the open interval $(x,y)_{G}$ as the subposet of $x,y]_{G}$ obtained by excluding $x$ and $y$.
+We say that a subgraph $H$ of $G$ is convex in $G$ if, for any $x,y\in V(H)$, every minimal walk from $x$ to $y$ traverses only vertices in $H$; equivalently, $x,y]_{G}\subseteq V(H)$ for $x,y\in V(H)$.
+Every convex subgraph is an isometric subgraph, but the converse is false.
 Definition 2.1.
-Let $G$ be a connected graph, $H$ an isometric subgraph of $G$ and $x\in V(G)$. We say $H$ is gated with respect to $x$ if there is $u\in V(H)$ such that for any $v\in V(H)$,
+Let $G$ be a connected graph, let $H$ be an isometric subgraph of $G$, and let $x\in V(G)$. We say that $H$ is gated with respect to $x$ if there is $u\in V(H)$ such that, for any $v\in V(H)$,
 $d(x,v)=d(x,u)+d(u,v).$
-Such $u$ must be unique since the above equation says $u$ is the closest vertex to $x$ in $H$.
-We may call $u$ the gate of $H$ with respect to $x$.
-Intuitively, any minimal walk from $x$ to a vertex of $H$ enters $H$ at the gate.
-We say $H$ is a gated subgraph of $G$ is it is gated with respect to all vertices of $G$.
+Such a vertex $u$ must be unique, since the above equation says that $u$ is the closest vertex to $x$ in $H$.
+We call $u$ the gate of $H$ with respect to $x$.
+Intuitively, every minimal walk from $x$ to a vertex of $H$ enters $H$ at the gate.
+We say that $H$ is a gated subgraph of $G$ if it is gated with respect to every vertex of $G$.
 In this case, we call the map $p:G\to H$ taking any vertex of $G$ to its gate in $H$ the gate projection.
 Note that the restriction of $p$ to $H$ is the identity map $\text{id}_{H}$.
-Proposition 2.2 (Proposition A.1 of [2]).
+Proposition 2.2 ([2, Proposition A.1]).
 A gated subgraph is necessarily convex.
 Proof.
-Suppose $H$ is gated in $G$. Choose any minimal walk in $G$ from $x$ to $y$ for $x,y\in V(H)$ and let $z$ be an arbitrary vertex in this walk, we need to show $z\in V(H)$. Since $H$ is gated, we have
+Suppose $H$ is gated in $G$.
+Choose a minimal walk in $G$ from $x$ to $y$ for $x,y\in V(H)$, and let $z$ be an arbitrary vertex in this walk.
+We need to show that $z\in V(H)$.
+Since $H$ is gated, we have
 d(z,x)=d(z,p(z))+d(p(z),x),d(z,y)=d(z,p(z))+d(p(z),y).d(z,x)=d(z,p(z))+d(p(z),x),~d(z,y)=d(z,p(z))+d(p(z),y).
@@
 This forces $d(z,p(z))=0$, i.e. $z=p(z)\in V(H)$. ∎
-In the scope of this paper, the homomorphisms between graphs we shall use are distance non-increasing maps (1-Lipschitz maps) $f:G\to H$, which are maps $f:V(G)\to V(H)$ such that $d_{H}(f(x),f(y))\leq d_{G}(x,y)$ for $x,y\in V(G)$.
+In this paper, the graph homomorphisms we use are distance non-increasing maps (1-Lipschitz maps) $f:G\to H$, that is, maps $f:V(G)\to V(H)$ such that $d_{H}(f(x),f(y))\leq d_{G}(x,y)$ for $x,y\in V(G)$.
 Note that gate projections are distance non-increasing.
@@
 Note that $Z_{G}(0)$ is the identity matrix and hence $\det Z_{G}(q)\in\mathbb{Z}[q]$ is a polynomial with constant term $1$.
-The rational function $1/\det Z_{G}(q)\in\mathbb{Q}(q)$ then can be expressed as a power series in $\mathbb{Z}[\![q]\!]$.
+The rational function $1/\det Z_{G}(q)\in\mathbb{Q}(q)$ can then be expressed as a power series in $\mathbb{Z}[\![q]\!]$.
 In particular, $Z_{G}(q)$ is invertible over $\mathbb{Q}(q)$, with inverse $Z_{G}(q)^{-1}=\frac{1}{\det Z_{G}(q)}\mathrm{adj}~Z_{G}(q)$, whose entries can be expressed as power series in $\mathbb{Z}[\![q]\!]$.
@@
 If $B$ is invertible, then the sum of all entries of $B^{-1}$ can be computed from $\mathbf{1}^{T}B^{-1}\mathbf{1}=\mathbf{1}^{T}\mathbf{x}$, where $\mathbf{x}$ is the unique solution of $B\mathbf{x}=\mathbf{1}$.
-For graph $G$, define the weight vector $w_{G}:=Z_{G}^{-1}\mathbf{1}$ as the unique solution to $Z_{G}\mathbf{x}=\mathbf{1}$, in other words, as a function $w_{G}:V\to\mathbb{Q}(q)$, $w_{G}(x)=\sum_{y\in V}Z_{G}(q)^{-1}_{x,y}$ is the $x$-row sum of $Z_{G}^{-1}$.
-Then the magnitude
+For a graph $G$, define the weight vector $w_{G}:=Z_{G}^{-1}\mathbf{1}$ as the unique solution to $Z_{G}\mathbf{x}=\mathbf{1}$.
+In other words, as a function $w_{G}:V\to\mathbb{Q}(q)$, $w_{G}(x)=\sum_{y\in V}Z_{G}(q)^{-1}_{x,y}$ is the $x$-row sum of $Z_{G}^{-1}$.
+Then the magnitude is given by
 $\mathrm{Mag}(G)=\mathbf{1}^{T}w_{G}=\sum_{x\in V}w_{G}(x).$
-There is a simple case that we can compute magnitude easily.
-Proposition 2.3 (Lemma 3.2 of [25]).
+The magnitude of a vertex-transitive graph can be computed explicitly.
+Proposition 2.3 ([25, Lemma 3.2]).
 If $G$ is vertex-transitive (i.e. $\mathrm{Aut}(G)$ acts transitively on $V$), then
-$\mathrm{Mag}(G)=\frac{\#V}{\sum_{x\in V}q^{d(g,x)}},$
+$\mathrm{Mag}(G)=\frac{\#V}{\sum_{x\in V}q^{d(g,x)}}$
 for any $g\in V$.
 Proof.
-The denominator $s(g)=\sum_{x\in V}q^{d(g,x)}$ is independent of $g\in V$ by the assumption that $G$ is vertex-transitive, and thus denoted by $s=s(g)$.
-Then it is easily seen that $w_{G}=[1/s\cdots 1/s]^{T}$ is the solution of $Z_{G}\mathbf{x}=\mathbf{1}$.
+The denominator $s(g)=\sum_{x\in V}q^{d(g,x)}$ is independent of $g\in V$ by the assumption that $G$ is vertex-transitive, and we denote it by $s=s(g)$.
+It is then easy to see that $w_{G}=[1/s\cdots 1/s]^{T}$ is the solution of $Z_{G}\mathbf{x}=\mathbf{1}$.
 We conclude $\mathrm{Mag}(G)=\#V/s$.
@@
 Example 2.4.
-Let $K_{n}$ be the complete graph on $n$ vertices and $C_{n}$ be the cycle graph on $n$ vertices. They are vertex-transitive. Then
+Let $K_{n}$ be the complete graph on $n$ vertices and let $C_{n}$ be the cycle graph on $n$ vertices. They are vertex-transitive. Their magnitudes are given by
 $\mathrm{Mag}(K_{n})$ $=\frac{n}{1+(n-1)q},$
@@
 $\mathrm{Mag}(C_{2n-1})$ $=\frac{(2n-1)(q-1)}{2q^{n}-q-1}=\frac{2n-1}{1+2q+2q^{2}+\cdots+2q^{n-1}}.$
-Magnitude behaves as cardinality in the following sense.
-Proposition 2.5 (Lemmas 3.5 and 3.6 of [25]).
-Let $G,H$ be graphs. Write $G\sqcup H$ and $G\square H$ the disjoint union and cartesian product of $G$ and $H$ respectively. Then
-$\mathrm{Mag}(G\sqcup H)=\mathrm{Mag}(G)+\mathrm{Mag}(H),~\mathrm{Mag}(G\square H)=\mathrm{Mag}(G)\mathrm{Mag}(H).$
-Under some conditions, magnitude satisfies inclusion-exclusion.
-Theorem 2.6 (Theorem 4.9 of [25]).
+Magnitude behaves like cardinality in the following sense.
+Proposition 2.5 ([25, Lemmas 3.5 and 3.6]).
+Let $G,H$ be graphs. Write $G\sqcup H$ and $G\square H$ for the disjoint union and Cartesian product of $G$ and $H$, respectively. Then
+$\mathrm{Mag}(G\sqcup H)=\mathrm{Mag}(G)+\mathrm{Mag}(H),\quad\mathrm{Mag}(G\square H)=\mathrm{Mag}(G)\mathrm{Mag}(H).$
+Under certain conditions, magnitude satisfies inclusion-exclusion.
+Theorem 2.6 ([25, Theorem 4.9]).
 If a graph $X$ has a decomposition into subgraphs $X=G\cup H$ such that $G\cap H$ is convex in $X$ and gated in $H$, then
@@
 For a graph $G$, we call a $(k+1)$-tuple $\vec{x}=(x_{0},x_{1},\ldots,x_{k})\in V^{k+1}$ of vertices of $G$ a $k$-chain of $G$ and define its length by $\ell(\vec{x})=d(x_{0},x_{1})+d(x_{1},x_{2})+\cdots+d(x_{k-1},x_{k})$.
-We say a $k$-chain $\vec{x}$ is proper if adjacent nodes are different $x_{0}\neq x_{1}\neq\cdots\neq x_{k}$.
-Let $P_{k}(G)$ be the set of proper $k$-chains of $G$, and $P_{k,\ell}(G)$ the set of proper $k$-chains of length $\ell$.
+We say that a $k$-chain $\vec{x}$ is proper if adjacent nodes are distinct, that is, if $x_{i}\neq x_{i+1}$ for all $i$.
+Let $P_{k}(G)$ be the set of proper $k$-chains of $G$, and let $P_{k,\ell}(G)$ be the set of proper $k$-chains of length $\ell$.
 The following result of Leinster is the starting point of the categorification of graph magnitude introduced by Hepworth–Willerton [17].
-Proposition 2.7 (Proposition 3.9 of [25]).
-For a graph $G$,
+Proposition 2.7 ([25, Proposition 3.9]).
+For a graph $G$, we have
 $\mathrm{Mag}(G)=\sum_{k=0}^{\infty}(-1)^{k}\sum_{\vec{x}\in P_{k}(G)}q^{\ell(\vec{x})}\in\mathbb{Z}[\![q]\!].$
@@
 $c_{\ell}=\sum_{k=0}^{\ell}(-1)^{k}\#P_{k,\ell}(G).$
-In particular, $V=c_{0}$ and $E=\frac{-c_{1}}{2}$.
+In particular, we have $c_{0}=\#V$ and $c_{1}=-2\#E$.
 Let $MC_{k,\ell}(G)$ be the free abelian group generated by $P_{k,\ell}(G)$.
-We define the boundary operator $\partial:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ as $\partial=\sum_{i=1}^{k-1}(-1)^{i-1}\partial_{i}$, where $\partial_{i}:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ is defined by
+We define the boundary operator $\partial:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ by $\partial=\sum_{i=1}^{k-1}(-1)^{i-1}\partial_{i}$, where $\partial_{i}:MC_{k,\ell}(G)\to MC_{k-1,\ell}(G)$ is defined by
 $\partial_{i}(x_{0},\ldots,x_{k})=\begin{cases}(x_{0},\ldots,\widehat{x_{i}},\ldots,x_{k}),&\text{ if }d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1});\\ 0,&\text{ otherwise},\end{cases}$
-where $\widehat{x_{i}}$ means deletion of $x_{i}$.
-One can verify $\partial^{2}=0$.
-Therefore we call the homology of the chain complex $(MC_{*,\ell}(G),\partial)$ the length $\ell$ magnitude homology of $G$, denoted by $MH_{*,\ell}(G)$.
+where $\widehat{x_{i}}$ denotes deletion of $x_{i}$.
+One can verify that $\partial^{2}=0$.
+We call the homology of the chain complex $(MC_{*,\ell}(G),\partial)$ the length $\ell$ magnitude homology of $G$, denoted by $MH_{*,\ell}(G)$.
 Since the boundary map preserves endpoints, we have a decomposition
 $MC_{k,\ell}(G)=\bigoplus_{u,v\in V(G)}MC_{k,\ell}(G;u,v),$
-where $MC_{k,\ell}(G;u,v)$ is the subgroup generated by chains $P_{k,\ell}(G;u,v)=\{\vec{x}\in P_{k,\ell}(G)\mid x_{0}=u,x_{k}=v\}$, forming subcomplexes and inducing decomposition of homology groups
+where $MC_{k,\ell}(G;u,v)$ is the subgroup generated by chains $P_{k,\ell}(G;u,v)=\{\vec{x}\in P_{k,\ell}(G)\mid x_{0}=u,\ x_{k}=v\}$.
+These subgroups form subcomplexes and induce a decomposition of homology groups
 (2.2) $MH_{k,\ell}(G)=\bigoplus_{u,v\in V(G)}MH_{k,\ell}(G;u,v).$
-Let us denote the magnitude Betti number $\beta_{k,\ell}(G)=\mathrm{rank}MH_{k,\ell}(G)$ and the length $\ell$ magnitude Euler characteristic $\chi_{\ell}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}(G)$.
-Proposition 2.8 (Propositions 2.9, 2.10 of [16]).
+We define the magnitude Betti number by $\beta_{k,\ell}(G)=\operatorname{rank}MH_{k,\ell}(G)$ and the length $\ell$ magnitude Euler characteristic by $\chi_{\ell}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}(G)$.
+Proposition 2.8 ([17, Propositions 2.9 and 2.10]).
 Let $G=(V,E)$ be a graph.
@@
   3. (3)
-    If $MH_{k,\ell}(G)\neq 0$, then $\ell/\mathrm{diam}(G)\leq k\leq\ell$ and the inequality $\ell/\mathrm{diam}(G)<k$ is strict if $\mathrm{diam}(G)>1$ and $\ell>0$.
-The following proposition is straightforward from definition, which says magnitude can be recovered from magnitude homology.
-Proposition 2.9 (Theorem 2.8 of [16]).
+    If $MH_{k,\ell}(G)\neq 0$, then $\ell/\operatorname{diam}(G)\leq k\leq\ell$, and the inequality $\ell/\operatorname{diam}(G)\leq k$ is strict if $\operatorname{diam}(G)>1$ and $\ell>0$.
+The following proposition is a straightforward consequence of Proposition 2.7 and says that magnitude can be recovered from magnitude homology.
+Proposition 2.9 ([17, Theorem 2.8]).
 Let $G$ be a graph, and $\mathrm{Mag}(G)=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell}\in\mathbb{Z}[\![q]\!]$. Then
-$c_{\ell}=\chi_{\ell}(G)=\sum_{k=0}^{\infty}(-1)^{k}\beta_{k,\ell}(G)=\sum_{k=\ell/\mathrm{diam}(G)}^{\ell}(-1)^{k}\beta_{k,\ell}(G).$
+$c_{\ell}=\chi_{\ell}(G)=\sum_{k=0}^{\infty}(-1)^{k}\beta_{k,\ell}(G)=\sum_{k=\ell/\operatorname{diam}(G)}^{\ell}(-1)^{k}\beta_{k,\ell}(G).$
 Hepworth–Willerton proved that magnitude homology is a functor.
-Proposition 2.10 (Proposition 3.3 of [16]).
+Proposition 2.10 ([17, Proposition 3.3]).
 Magnitude homology is a functor from the category of finite graphs with distance non-increasing maps to the category of bigraded abelian groups.
-A direct application of the functorality is as follow.
-Let $G$ be a connected graph and $H$ be a gated subgraph of $G$ with gate projection $p:G\to H$.
+A direct application of functoriality is as follows.
+Let $G$ be a connected graph and let $H$ be a gated subgraph of $G$ with gate projection $p:G\to H$.
 Let $\iota:H\to G$ be the inclusion.
@@
 There is a Künneth formula for magnitude homology of graphs.
-Theorem 2.12 (Theorem 5.3 of [16]).
-There is a natural short exact sequence which non-naturally splits
+Theorem 2.12 ([17, Theorem 5.3]).
+There is a natural short exact sequence that splits non-naturally
 $0\to MH_{*,*}(G)\otimes MH_{*,*}(H)\xrightarrow{\square}MH_{*,*}(G\square H)\to\mathrm{Tor}(MH_{*-1,*}(G),MH_{*,*}(H))\to 0,$
-where $\square$ is the exterior product defined in Definition 5.2 of [16].
+where $\square$ is the exterior product defined in [17, Definition 5.2].
 Example 2.13.
-For the complete graph $K_{n}$,
-$MH_{k,\ell}(K_{n})=\begin{cases}\bigoplus_{\vec{x}\in P_{\ell}(K_{n})}\mathbb{Z}\vec{x},&\text{ if }k=l;\\ 0,&\text{ otherwise}.\end{cases}$
-The Betti number $\beta_{\ell,\ell}(K_{n})=n(n-1)^{\ell}$. One can verify $\mathrm{Mag}(K_{n},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(K_{n})q^{\ell}=\frac{n}{1+(n-1)q}$.
+For the complete graph $K_{n}$, we have
+$MH_{k,\ell}(K_{n})=\begin{cases}\bigoplus_{\vec{x}\in P_{\ell}(K_{n})}\mathbb{Z}\vec{x},&\text{ if }k=\ell;\\ 0,&\text{ otherwise}.\end{cases}$
+The magnitude Betti number is given by $\beta_{\ell,\ell}(K_{n})=n(n-1)^{\ell}$. One can verify that $\mathrm{Mag}(K_{n},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(K_{n})q^{\ell}=\frac{n}{1+(n-1)q}$.
 Example 2.14.
-For the hypercube graph $Q_{d}=K_{2}^{\square d}$, we use the Künneth formula (Theorem 2.12) to obtain
+For the hypercube graph $Q_{d}=K_{2}^{\square d}$, the Künneth formula (Theorem 2.12) gives
 $MH_{k,\ell}(Q_{d})=\bigoplus_{\sum k_{i}=k,\sum\ell_{i}=\ell}MH_{k_{1},\ell_{1}}(K_{2})\otimes\cdots\otimes MH_{k_{d},\ell_{d}}(K_{2}),$
-and use the result of $K_{2}$, we conclude
-$MH_{k,\ell}(Q_{d})=\begin{cases}\mathbb{Z}^{\beta_{\ell,\ell}(Q_{d})},&\text{ if }k=l;\\ 0,&\text{ otherwise},\end{cases}$
-where $\beta_{\ell,\ell}(Q_{d})=2^{d}\binom{l+d-1}{d-1}$. One can verify $\mathrm{Mag}(Q_{d},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(Q_{d})q^{\ell}=\frac{2^{d}}{(1+q)^{d}}$.
+Using the result for $K_{2}$, we obtain
+$MH_{k,\ell}(Q_{d})=\begin{cases}\mathbb{Z}^{\beta_{\ell,\ell}(Q_{d})},&\text{ if }k=\ell;\\ 0,&\text{ otherwise},\end{cases}$
+where $\beta_{\ell,\ell}(Q_{d})=2^{d}\binom{\ell+d-1}{d-1}$. One can verify that $\mathrm{Mag}(Q_{d},q)=\sum_{\ell}(-1)^{\ell}\beta_{\ell,\ell}(Q_{d})q^{\ell}=\frac{2^{d}}{(1+q)^{d}}$.
 Definition 2.15.
-We say $G$ is diagonal if $MH_{k,\ell}(G)=0$ for $k\neq\ell$.
+We say that $G$ is diagonal if $MH_{k,\ell}(G)=0$ for $k\neq\ell$.
 Example 2.16.
-We have seen that $K_{n}$ and $Q_{d}$ are diagonal.
-In particular the cycle graph $C_{4}=Q_{2}$ is diagonal.
-However Gu [15] computed $MH_{k,\ell}(C_{n})$ for all $n\geq 5$ and showed that $C_{n}(n\geq 5)$ are not diagonal.
-Let us recall results from [19] in the context of graphs.
+We have seen in Examples 2.13 and 2.14 that $K_{n}$ and $Q_{d}$ are diagonal.
+In particular, the cycle graph $C_{4}=Q_{2}$ is diagonal.
+However, Gu [16] computed $MH_{k,\ell}(C_{n})$ for all $n\geq 5$ and showed that the graphs $C_{n}$ ($n\geq 5$) are not diagonal.
+Kaneta–Yoshinaga [20] constructed a computationally tractable approximation of magnitude homology.
 Let $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$.
 We say $\vec{x}$ is geodesic if $\ell(\vec{x})=d(x_{0},x_{k})$.
-Any sub-chain of a geodesic proper chain is again geodesic proper.
+Every subchain of a geodesic proper chain is again geodesic and proper.
 Any proper $1$-chain is geodesic.
 A proper $2$-chain $(x_{0},x_{1},x_{2})$ is geodesic if and only if $d(x_{0},x_{2})=d(x_{0},x_{1})+d(x_{1},x_{2})$.
-Consider $u,v\in V(G)$, such that $\ell=d(u,v)$, then any chain in $P_{k,\ell}(G;u,v)$ must be geodesic.
+If $u,v\in V(G)$ are vertices and $\ell=d(u,v)$, then any chain in $P_{k,\ell}(G;u,v)$ must be geodesic.
 Definition 2.17.
-Let us define the length $\ell$ geodesic magnitude complex
-$MC_{*,\ell}^{geod}(G):=\bigoplus_{u,v\in V(G):d(u,v)=\ell}MC_{*,\ell}(G;u,v).$
-The homology is called the length $\ell$ geodesic magnitude homology, which admits the decomposition
-$MH_{k,\ell}^{geod}(G)=\bigoplus_{u,v\in V(G):d(u,v)=\ell}MH_{k,\ell}(G;u,v).$
-Note that $MC_{*,\ell}^{geod}(G)$ and $MH_{k,\ell}^{geod}(G)$ are trivial if $\ell>\mathrm{diam}(G)$.
-Let us denote the geodesic magnitude Betti number $\beta_{k,\ell}^{geo}(G)=\mathrm{rank}MH_{k,\ell}^{geod}(G)$ and the length $\ell$ geodesic magnitude Euler characteristic $\chi_{\ell}^{geod}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}^{geod}(G)$.
-Notice that $\beta_{k,\ell}(G)\geq\beta_{k,\ell}^{geod}(G)$.
-Kaneta–Yoshinaga [19] generalized the above geodesic magnitude chains.
-Definition 2.18.
-A proper $3$-chain $(x_{0},x_{1},x_{2},x_{3})\in P_{3}(G)$ is called a $4$-cut of $G$ if $(x_{0},x_{1},x_{2})$ and $(x_{1},x_{2},x_{3})$ are geodesic but $(x_{0},x_{1},x_{2},x_{3})$ is not geodesic.
-Write $m_{G}$ as the minimum length of $4$-cuts of $G$.
-We say $m_{G}=\infty$ if $G$ has no $4$-cuts.
+We define the length $\ell$ geodesic magnitude complex by
+$MC_{*,\ell}^{\mathrm{geod}}(G):=\bigoplus_{\begin{subarray}{c}u,v\in V(G)\\ d(u,v)=\ell\end{subarray}}MC_{*,\ell}(G;u,v).$
+Its homology is called the length $\ell$ geodesic magnitude homology.
+The geodesic magnitude homology $MH^{\mathrm{geod}}_{k,\ell}(G)$ is a direct summand of $MH_{k,\ell}(G)$.
+It admits a decomposition
+$MH_{k,\ell}^{\mathrm{geod}}(G)=\bigoplus_{\begin{subarray}{c}u,v\in V(G)\\ d(u,v)=\ell\end{subarray}}MH_{k,\ell}(G;u,v).$
+Note that $MC_{*,\ell}^{\mathrm{geod}}(G)$ and $MH_{k,\ell}^{\mathrm{geod}}(G)$ are trivial if $\ell>\operatorname{diam}(G)$.
+We define the geodesic magnitude Betti number by $\beta_{k,\ell}^{\mathrm{geod}}(G)=\operatorname{rank}MH_{k,\ell}^{\mathrm{geod}}(G)$ and the length $\ell$ geodesic magnitude Euler characteristic by $\chi_{\ell}^{\mathrm{geod}}(G)=\sum_{k}(-1)^{k}\beta_{k,\ell}^{\mathrm{geod}}(G)$.
+Observe that $\beta_{k,\ell}(G)\geq\beta_{k,\ell}^{\mathrm{geod}}(G)$.
+Geodesic magnitude homology is a coarse approximation of magnitude homology.
+As a more refined approximation, Kaneta–Yoshinaga [20] introduced geodesically simple magnitude homology.
 For a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$, an internal node $x_{i}$ ($i=1,2,\ldots,k-1$) is called a smooth node of $\vec{x}$ if $d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$.
@@
 Note that endpoints are singular.
-We write $\varphi(\vec{x})$ as the sub-chain of $\vec{x}$ consisting of all singular nodes, called the frame of $\vec{x}$.
+We write $\varphi(\vec{x})$ for the subchain of $\vec{x}$ consisting of all singular nodes and call it the frame of $\vec{x}$.
 Note that $\varphi(\vec{x})$ is not necessarily proper.
@@
 A proper chain $\vec{x}$ is geodesically simple if $\ell(\varphi(\vec{x}))=\ell(\vec{x})$.
-Note that proper $1$-chains and proper $2$-chains are all geodesically simple, and proper chains with length smaller than $m_{G}$ are geodesically simple.
-Kaneta–Yoshinaga proved that the subgroups $MC_{*,\ell}^{gs}(G)$ of $MC_{*,\ell}(G)$ generated by geodesically simple chains is a subcomplex, whose homology is called the geodesically simple magnitude homology of $G$, denoted by $MH_{*,\ell}^{gs}(G)$, and the homology group $MH_{k,\ell}^{gs}(G)$ factors as a direct sum of framed magnitude homology groups as explained below.
+Definition 2.19.
+We define $MC_{k,\ell}^{\mathrm{gs}}(G)$ to be the subgroup of $MC_{k,\ell}(G)$ generated by geodesically simple chains.
+Then $MC_{*,\ell}^{\mathrm{gs}}(G)$ forms a subcomplex of $MC_{k,\ell}(G)$ [20, Propositions 3.6 and 3.7].
+Its homology is called the geodesically simple magnitude homology of $G$, denoted by $MH_{*,\ell}^{\mathrm{gs}}(G)$.
+How well geodesically simple magnitude homology approximates magnitude homology is controlled by the notion of a $4$-cut.
+Definition 2.20.
+A proper $3$-chain $(x_{0},x_{1},x_{2},x_{3})\in P_{3}(G)$ is called a $4$-cut of $G$ if $(x_{0},x_{1},x_{2})$ and $(x_{1},x_{2},x_{3})$ are geodesic but $(x_{0},x_{1},x_{2},x_{3})$ is not geodesic.
+Let $m_{G}$ be the minimum length of $4$-cuts of $G$.
+We set $m_{G}=\infty$ if $G$ has no $4$-cuts.
+Note that $m_{G}\geq 3$.
+Any proper chain with length smaller than $m_{G}$ is geodesically simple.
+Lemma 2.21.
+If $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k}(G)$ satisfies $\varphi(\vec{x})=(x_{0},x_{k})$ and $\vec{x}$ is not geodesically simple, i.e.
+$(x_{i-1},x_{i},x_{i+1})$ is geodesic for $i=1,\ldots,k-1$, but $\vec{x}$ is not geodesic, then $\vec{x}$ has a subchain that is a $4$-cut.
+Proof.
+Let $\vec{y}=(y_{0},\ldots,y_{p})$ be a contiguous subchain of $\vec{x}$, that is, $y_{0}=x_{i},y_{1}=x_{i+1},\ldots,y_{p}=x_{i+p}$ for some $i$, such that $\vec{y}$ is not geodesic and $p$ is as small as possible.
+This means $(y_{0},\ldots,y_{p-1})$ and $(y_{1},\ldots,y_{p})$ are geodesic.
+Then $p\geq 3$.
+We claim $\vec{z}:=(y_{0},y_{1},y_{p-1},y_{p})$ is a $4$-cut.
+First note that $y_{p-1}\neq y_{1}$ as they are connected by a geodesic $(y_{1},y_{2},\ldots,y_{p-1})$.
+That $(y_{0},y_{1},y_{p-1})$ and $(y_{1},y_{p-1},y_{p})$ are geodesic follows from the assumption that $(y_{0},\ldots,y_{p-1})$ and $(y_{1},\ldots,y_{p})$ are geodesic.
+Moreover, $\ell(\vec{z})=d(y_{0},y_{1})+d(y_{1},y_{p-1})+d(y_{p-1},y_{p})=\sum_{i=1}^{p}d(y_{i-1},y_{i})>d(y_{0},y_{p})$, since $\vec{y}$ is not geodesic.
+This shows $\vec{z}$ is not geodesic and hence is a $4$-cut.
+∎
 Let $\vec{a}\in P_{m,\ell}(G)$ and $k\geq m$.
 We define $P_{k,\vec{a}}(G):=\{\vec{x}\in P_{k,\ell}(G)\mid\varphi(\vec{x})=\vec{a}\}$.
-Note that $P_{k,\vec{a}}(G)$ consists of geodesically simple chains and would be empty if $\varphi(\vec{a})\neq\vec{a}$.
-We further define $MC_{k,\vec{a}}(G)$ as the free abelian group generated by $P_{k,\vec{a}}(G)$, which is a subgroup of $MC_{k,\ell}^{gs}(G)$.
-They form a subcomplex $MC_{*,\vec{a}}(G)$ whose homology $MH_{*,\vec{a}}(G)$ is called the $\vec{a}$-framed magnitude homology of $G$.
-Kaneta–Yoshinaga proved the chain complex decomposes
-(2.2) $MC_{k,\ell}^{gs}(G)=\bigoplus_{m\leq k,\vec{a}\in P_{m,\ell}(G)}MC_{k,\vec{a}}(G),$
-and so does the homology group.
-Theorem 2.20 (Theorem 3.12 of [19]).
-Let $G$ be a graph and $k,\ell>0$. We have
-$MH_{k,\ell}^{gs}(G)\cong\bigoplus_{m\leq k,\vec{a}\in P_{m,\ell}(G)}MH_{k,\vec{a}}(G).$
-Furthermore, if $k=1$ or $\ell<m_{G}$, then $MH_{k,\ell}^{gs}(G)\cong MH_{k,\ell}(G)$.
-Kaneta–Yoshinaga also related framed magnitude homology with open intervals determined by the frame.
+Note that $P_{k,\vec{a}}(G)$ consists of geodesically simple chains and is empty if $\varphi(\vec{a})\neq\vec{a}$.
+We further define $MC^{\mathrm{gs}}_{k,\vec{a}}(G)$ as the free abelian group generated by $P_{k,\vec{a}}(G)$, which is a subgroup of $MC_{k,\ell}^{\mathrm{gs}}(G)$.
+These groups form a subcomplex $MC^{\mathrm{gs}}_{*,\vec{a}}(G)$ whose homology $MH^{\mathrm{gs}}_{*,\vec{a}}(G)$ is called the $\vec{a}$-framed magnitude homology of $G$.
+Kaneta–Yoshinaga proved that the chain complex $MC_{*,\ell}^{\mathrm{gs}}(G)$ decomposes as
+(2.3) $MC_{k,\ell}^{\mathrm{gs}}(G)=\bigoplus_{\begin{subarray}{c}m\leq k\\ \vec{a}\in P_{m,\ell}(G)\end{subarray}}MC^{\mathrm{gs}}_{k,\vec{a}}(G),$
+and the same is true for the homology groups.
+Theorem 2.22 ([20, Theorem 3.12]).
+Let $G$ be a graph and $k,\ell>0$. We have a decomposition
+$MH_{k,\ell}^{\mathrm{gs}}(G)\cong\bigoplus_{\begin{subarray}{c}m\leq k\\ \vec{a}\in P_{m,\ell}(G)\end{subarray}}MH^{\mathrm{gs}}_{k,\vec{a}}(G).$
+Furthermore, if $k=1$ or $\ell<m_{G}$, then $MH_{k,\ell}^{\mathrm{gs}}(G)\cong MH_{k,\ell}(G)$.
+Kaneta–Yoshinaga also related framed magnitude homology to open intervals determined by the frame.
 Given $a,b\in V(G)$, consider the open interval $(a,b)_{G}$ introduced in Section 2.1.
-Also write $C_{*}(a,b)_{G}$ the reduced chain complex of the order complex of $(a,b)_{G}$.
-Proposition 2.21 (Proposition 2.3 of [13]).
-For $a,b\in G$ with $d(a,b)=\ell$, there is an isomorphism of chain complexes
+We also write $C_{*}(a,b)_{G}$ for the reduced chain complex of the order complex of $(a,b)_{G}$.
+Proposition 2.23 ([13, Proposition 2.3]).
+For $a,b\in V(G)$ with $d(a,b)=\ell$, there is an isomorphism of chain complexes
 $MC_{*,\ell}(G;a,b)\to C_{*-2}(a,b)_{G},$
-inducing an isomorphism of homology groups
+which induces an isomorphism of homology groups
 $MH_{k,\ell}(G;a,b)\xrightarrow{\cong}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
 Proof.
-For a proper chain $\vec{x}=(a,x_{1},\ldots,x_{k-1},b)\in P_{k,\ell}(G;a,b)$, it must be geodesic since $\ell=d(a,b)$.
-Then mapping $\vec{x}$ to the linearly order chain $x_{1}\prec_{G}\cdots\prec_{G}x_{k-1}$ in $C_{k-2}(a,b)_{G}$ defines the desired isomorphism.
+Any proper chain $\vec{x}=(a,x_{1},\ldots,x_{k-1},b)\in P_{k,\ell}(G;a,b)$ must be geodesic since $\ell=d(a,b)$.
+Mapping $\vec{x}$ to the linearly ordered chain $x_{1}\prec_{G}\cdots\prec_{G}x_{k-1}$ in $C_{k-2}(a,b)_{G}$ defines the desired isomorphism.
 ∎
 Corollary 2.24.
-The geodesic magnitude homology has the decomposition
-$MH_{k,\ell}^{geod}(G)\cong\bigoplus_{a,b\in V(G):d(a,b)=\ell}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
+The geodesic magnitude homology admits a decomposition
+$MH_{k,\ell}^{\mathrm{geod}}(G)\cong\bigoplus_{\begin{subarray}{c}a,b\in V(G)\\ d(a,b)=\ell\end{subarray}}\widetilde{H}_{k-2}(C_{*}(a,b)_{G}).$
 For a frame $\vec{a}=(a_{0},a_{1},\ldots,a_{m})\in P_{m,\ell}(G)$, we denote by $C_{*}(\vec{a})_{G}$ the total complex of $C_{*}(a_{0},a_{1})_{G}\otimes\cdots\otimes C_{*}(a_{m-1},a_{m})_{G}$.
-A more generalized version of Proposition 2.21 is as follow.
-Theorem 2.23 (Theorem 4.4 and Corollary 4.5 of [19]).
-Suppose a frame $\vec{a}$ has length $\ell(\vec{a})<m_{G}$, then there is an isomorphism of chain complexes
-$MC_{*,\vec{a}}(G)\to C_{*-2m}(\vec{a})_{G}$
-grouping smooth nodes in order between $a_{i}$ and $a_{i+1}$, which induces isomorphisms of homology groups
-$MH_{k,\vec{a}}(G)\cong\widetilde{H}_{k-2m}(C_{*}(\vec{a})_{G}).$
+A more general version of Proposition 2.23 is as follows.
+Theorem 2.25 ([20, Theorem 4.4 and Corollary 4.5]).
+Suppose a frame $\vec{a}$ has length $\ell(\vec{a})<m_{G}$. Then there is an isomorphism of chain complexes
+$MC^{\mathrm{gs}}_{*,\vec{a}}(G)\to C_{*-2m}(\vec{a})_{G},$
+which induces isomorphisms of homology groups
+$MH^{\mathrm{gs}}_{k,\vec{a}}(G)\cong\widetilde{H}_{k-2m}(C_{*}(\vec{a})_{G}).$
+Remark 2.26.
+The original statement of Theorem 4.4 in [20] does not contain the assumption $\ell(\vec{a})<m_{G}$.
+The result fails if a frame node $a_{i}$ of $\vec{a}$ can be smoothed out by nodes from neighboring intervals.
+For example, let $G=C_{4}$ be the cycle graph of $4$ vertices, labeled by $0,1,2,3$ in cyclic order.
+Choose $\vec{a}=(0,1,3)\in P_{2,3}(G)$.
+Then $(0,1)_{G}$ is empty and $(1,3)_{G}$ consists of two incomparable points $\{0,2\}$.
+Compare $MC^{\mathrm{gs}}_{3,\vec{a}}(G)$ with $C_{-1}(\vec{a})_{G}$.
+The element $\varnothing\otimes 0$ in $C_{-1}(\vec{a})_{G}$ does have a preimage $(0,1,0,3)$ in $MC^{\mathrm{gs}}_{3,\vec{a}}(G)$.
+However, the element $\varnothing\otimes 2$ in $C_{-1}(\vec{a})_{G}$ does not have a preimage because $(0,1,2,3)$ is a $4$-cut.
+Proof of Theorem 2.25.
+Let $\vec{a}=(a_{0},a_{1},\ldots,a_{m})\in P_{m,\ell}(G)$ such that $\varphi(\vec{a})=\vec{a}$.
+Then $\ell=\ell(\vec{a})<m_{G}$.
+If $\vec{x}\in P_{k,\vec{a}}(G)$ is a basis element of $MC^{\mathrm{gs}}_{k,\vec{a}}(G)$, then $\vec{x}$ has the form
+$\vec{x}=(a_{0},x^{1}_{1},\ldots,x^{1}_{k_{1}},a_{1},x^{2}_{1},\ldots,x^{2}_{k_{2}},a_{2},\ldots,a_{m}).$
+Note that $k=m+\sum_{i=1}^{m}k_{i}$.
+The smooth nodes between $a_{i-1}$ and $a_{i}$ form a linearly ordered chain $\langle x^{i}\rangle:=(x^{i}_{1}\prec_{G}\cdots\prec_{G}x^{i}_{k_{i}})$ in $C_{k_{i}-1}(a_{i-1},a_{i})_{G}$.
+Define $\Phi_{k}:MC^{\mathrm{gs}}_{k,\vec{a}}(G)\to C_{k-2m}(\vec{a})_{G}$ by
+$\Phi_{k}(\vec{x})=\langle x^{1}\rangle\otimes\langle x^{2}\rangle\otimes\cdots\otimes\langle x^{m}\rangle.$
+The map $\Phi_{k}$ is injective on basis elements.
+To prove that $\Phi_{k}$ is also surjective, consider a basis element $\langle y^{1}\rangle\otimes\cdots\otimes\langle y^{m}\rangle$ of $C_{k-2m}(\vec{a})_{G}$, where $\langle y^{i}\rangle=(y^{i}_{1}\prec_{G}\cdots\prec_{G}y^{i}_{k_{i}})\in C_{k_{i}-1}(a_{i-1},a_{i})_{G}$ for $i=1,\ldots,m$.
+Let
+$\vec{y}=(a_{0},y^{1}_{1},\ldots,y^{1}_{k_{1}},a_{1},y^{2}_{1},\ldots,y^{2}_{k_{2}},a_{2},\ldots,a_{m}).$
+We have $\vec{y}\in P_{k,\ell}(G)$.
+It remains to prove $\varphi(\vec{y})=\vec{a}$.
+First note that $(a_{i-1},y^{i}_{1},\ldots,y^{i}_{k_{i}},a_{i})$ is geodesic (2.1), and hence $y^{i}_{j}$ is smooth in $\vec{y}$.
+Next suppose, for contradiction, that $a_{i}$ is smooth in $\vec{y}$ for some $i=1,\ldots,m-1$, that is,
+$d(y^{i}_{k_{i}},y^{i+1}_{1})=d(y^{i}_{k_{i}},a_{i})+d(a_{i},y^{i+1}_{1}).$
+Here $y^{i}_{k_{i}}$ is understood as $a_{i-1}$ if $k_{i}=0$.
+Similarly $y^{i+1}_{1}$ is understood as $a_{i+1}$ if $k_{i+1}=0$.
+Note that $k_{i}$ and $k_{i+1}$ cannot both be zero, since otherwise $\varphi(\vec{a})\neq\vec{a}$.
+Consider the subchain
+$\vec{z}=(a_{i-1},y^{i}_{1},\ldots,y^{i}_{k_{i}},a_{i},y^{i+1}_{1},\ldots,y^{i+1}_{k_{i+1}},a_{i+1}).$
+It satisfies $\varphi(\vec{z})=(a_{i-1},a_{i+1})$ and $\ell(\vec{z})=d(a_{i-1},a_{i})+d(a_{i},a_{i+1})>d(a_{i-1},a_{i+1})$ by the assumption that $a_{i}$ is singular in $\vec{a}$.
+By Lemma 2.21, $\vec{z}$ contains a subchain $\vec{w}$ that is a $4$-cut.
+However, $\ell(\vec{w})\leq\ell(\vec{z})\leq\ell(\vec{y})=\ell$, which contradicts the assumption $\ell<m_{G}$.
+The verification that $\Phi_{k}$ defines a chain map is omitted.
+∎
 2.4. Hyperplane arrangements
-In this paper, all arrangements are real central hyperplane arrangements unless otherwise specified.
-Standard references are [27, 30].
-Let $\mathcal{A}$ be an arrangement in $\mathbb{R}^{\ell}$.
-We denote by $L(\mathcal{A})=\{\cap_{H\in\mathcal{B}}H\mid\mathcal{B}\subseteq\mathcal{A}\}$ the intersection lattice of $\mathcal{A}$, ordered by reverse inclusions.
+The standard references are [28, 32].
+Let $\mathcal{A}$ be an arrangement in $\mathbb{R}^{d}$.
+For two chambers $C,C^{\prime}\in\mathrm{Ch}(\mathcal{A})$, we write $S(C,C^{\prime})$ for the set of hyperplanes in $\mathcal{A}$ that separate $C$ and $C^{\prime}$.
+The tope graph $\mathcal{T}(\mathcal{A})$ of $\mathcal{A}$ is the graph with vertex set $\mathrm{Ch}(\mathcal{A})$ in which two vertices $C$ and $C^{\prime}$ are connected by an edge if $S(C,C^{\prime})$ consists of a single hyperplane.
+The graph metric $d$ on $\mathcal{T}(\mathcal{A})$ satisfies $d(C,C^{\prime})=\#S(C,C^{\prime})$.
+Note that $\mathcal{T}(\mathcal{A})$ is a partial cube, that is, it is isomorphic to an isometric subgraph of the hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}$.
+We denote by $L(\mathcal{A})=\{\cap_{H\in\mathcal{B}}H\mid\mathcal{B}\subseteq\mathcal{A}\}$ the intersection lattice of $\mathcal{A}$, ordered by reverse inclusion.
 An element of $L(\mathcal{A})$ is called a flat.
-Note that the maximal flat is $\hat{1}=\cap_{H\in\mathcal{A}}H=:\cap\mathcal{A}$, called the center of $\mathcal{A}$, and the minimal flat $\hat{0}$ is the ambient space $\mathbb{R}^{\ell}$.
+Note that the maximal flat is $\hat{1}=\cap_{H\in\mathcal{A}}H=:\cap\mathcal{A}$, called the center of $\mathcal{A}$, and the minimal flat $\hat{0}$ is the ambient space $\mathbb{R}^{d}$.
 The arrangement is called essential if the center is the origin.
-The rank of a flat $X\in L(\mathcal{A})$ is defined as the codimension $\mathrm{rank}X=\mathrm{codim}X$.
-The rank of $\mathcal{A}$ is defined as $\mathrm{rank}\mathcal{A}=\mathrm{rank}(\cap\mathcal{A})$.
-We write $L_{k}(\mathcal{A})$ as the set of flats of rank $k$ and $L_{\leq k}(\mathcal{A})$ the set of flats of rank at most $k$.
+The rank of a flat $X\in L(\mathcal{A})$ is defined as the codimension: $\operatorname{rank}X=\operatorname{codim}X$.
+The rank of $\mathcal{A}$ is defined as $\operatorname{rank}\mathcal{A}=\operatorname{rank}(\cap\mathcal{A})$.
+We write $L_{k}(\mathcal{A})$ for the set of flats of rank $k$ and $L_{\leq k}(\mathcal{A})$ for the set of flats of rank at most $k$.
 Using the Möbius function $\mu:L(\mathcal{A})\times L(\mathcal{A})\to\mathbb{Z}$, we define the characteristic polynomial of $\mathcal{A}$ by $\chi(\mathcal{A},t)=\sum_{X\in L(\mathcal{A})}\mu(\hat{0},X)t^{\dim X}$.
-It is classically known $\mathrm{Ch}(\mathcal{A})=|\chi(\mathcal{A},1)|$ [33].
-The arrangement $\mathcal{A}$ determines a stratification of $\mathbb{R}^{\ell}$, the strata are called faces (covectors).
-Let $\mathcal{F}(\mathcal{A})$ be the poset of faces, ordered by closure inclusions.
-The unique minimal face is the center $\cap\mathcal{A}=\cap_{H\in\mathcal{A}}H$, and maximal faces are the chambers $\mathrm{Ch}(\mathcal{A})$.
+It is classically known that $\mathrm{Ch}(\mathcal{A})=|\chi(\mathcal{A},-1)|$ [35].
+The arrangement $\mathcal{A}$ determines a stratification of $\mathbb{R}^{d}$; the strata are called faces (covectors).
+Let $\mathcal{F}(\mathcal{A})$ be the poset of faces, ordered by closure inclusion.
+The unique minimal face is the center $\cap\mathcal{A}=\cap_{H\in\mathcal{A}}H$, and the maximal faces are the chambers $\mathrm{Ch}(\mathcal{A})$.
 For faces $F,G\in\mathcal{F}(\mathcal{A})$, their meet $F\wedge G$ always exists, while their join $F\vee G$ may not exist.
-There is an order-reversing surjective map $s:\mathcal{F}(\mathcal{A})\to L(\mathcal{A})$ taking a face $F$ to its support $s(F)$, that is the minimal subspace of $\mathbb{R}^{\ell}$ containing $F$.
+There is an order-reversing surjective map $s:\mathcal{F}(\mathcal{A})\to L(\mathcal{A})$ taking a face $F$ to its support $s(F)$, that is, the minimal subspace of $\mathbb{R}^{d}$ containing $F$.
 We define the dimension of a face as the dimension of its support.
-For $F\in\mathcal{F}(\mathcal{A})$, let $\mathcal{A}_{F}=\{H\in\mathcal{A}\mid F\subseteq H\}$.
+For a face $F\in\mathcal{F}(\mathcal{A})$, let $\mathcal{A}_{F}=\{H\in\mathcal{A}\mid F\subseteq H\}$.
 Also, for a flat $X\in L(\mathcal{A})$, let $\mathcal{A}_{X}=\{H\in\mathcal{A}\mid X\subseteq H\}$ and $\mathcal{A}^{X}=\{H\cap X\mid H\in\mathcal{A}\setminus\mathcal{A}_{X}\}$.
@@
 Definition 2.27.
-For a flat $X\in L(A)$, we denote by
+For a flat $X\in L(\mathcal{A})$, we denote by
 $c^{X}:=\#s^{-1}(X)$
-the number of faces whose support is $X$, which is also equal to $\mathrm{Ch}(\mathcal{A}^{X})$.
-If $X$ is not the ambient space $\hat{0}$$\mathbb{R}^{\ell}$, we also define the beta invariant of $X$ by
-$\beta_{X}:=\beta(\mathcal{A}_{X}).$
-Recall that the beta invariant of an arrangement is defined as
-$\beta(\mathcal{A})=(-1)^{\mathrm{rank}\mathcal{A}-1}\chi^{\prime}(\mathcal{A},1)=|\chi^{\prime}(\mathcal{A},1)|.$
+the number of faces whose support is $X$, which is also equal to $\mathrm{Ch}(\mathcal{A}^{X})$. If $X$ is not the ambient space $\hat{0}=\mathbb{R}^{d}$, we also define the beta invariant of $X$ by
+$\beta_{X}:=\beta(\mathcal{A}_{X}),$
+where the beta invariant of an arrangement is defined as
+$\beta(\mathcal{A})=(-1)^{\operatorname{rank}\mathcal{A}-1}\chi^{\prime}(\mathcal{A},1)=|\chi^{\prime}(\mathcal{A},1)|.$
 Remark 2.28.
 The beta invariant $\beta_{X}$ counts the bounded chambers in the deconing of $\mathcal{A}_{X}$.
-In other words, let $F$ be a face whose support is $X$ and pick $H\in\mathcal{A}$ such that $F\subseteq H$, then $\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid\overline{C}\cap H=\overline{F}\}$ (See [33, 2, 30]).
-Hence the product $c^{X}\beta_{X}$ is also a counting function
+In other words, let $F$ be a face whose support is $X$ and pick $H\in\mathcal{A}$ such that $F\subseteq H$.
+Then $\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid\overline{C}\cap H=\overline{F}\}$ (see [35, 2, 32]).
+Hence the product $c^{X}\beta_{X}$ is also a counting function:
 $c^{X}\beta_{X}=\frac{1}{2}\#\{C\in\mathrm{Ch}(\mathcal{A})\mid s(\overline{C}\cap H)=X\}.$
-For arrangement $\mathcal{A}$ in $V=\mathbb{R}^{\ell}$, we fix a defining form $\alpha_{H}\in V^{*}$ for each $H\in\mathcal{A}$ and consider the line segment $L_{H}=[-\alpha_{H},\alpha_{H}]$ in $V^{*}$.
-We write $Z(\mathcal{A})=\sum_{H\in\mathcal{A}}L_{H}$ as the Minkowski sum of the line segments and call it the zonotope of $\mathcal{A}$.
+For an arrangement $\mathcal{A}$ in $V=\mathbb{R}^{d}$, we fix a defining form $\alpha_{H}\in V^{*}$ for each $H\in\mathcal{A}$ and consider the line segment $L_{H}=[-\alpha_{H},\alpha_{H}]$ in $V^{*}$.
+We write $Z(\mathcal{A})=\sum_{H\in\mathcal{A}}L_{H}$ for the Minkowski sum of the line segments and call it the zonotope of $\mathcal{A}$.
 The poset of nonempty faces of this polytope is anti-isomorphic to $\mathcal{F}(\mathcal{A})$.
-For a face $F\in\mathcal{F}(A)$, we say the corresponding face $z_{F}$ of $Z(\mathcal{A})$ dual to $F$.
-In particular, 0-faces (vertices) of $Z(\mathcal{A})$ are dual to chambers of $\mathcal{A}$.
-With these $\alpha_{H}$ chosen, each face $F\in\mathcal{F}(\mathcal{A})$ can be represented by a sign vector $F=(F_{H})_{H\in\mathcal{A}}\subseteq\{+,-,0\}^{\mathcal{A}}$, where $F_{H}=\mathrm{sgn}(\alpha_{H}(F))$.
-Declare $0<\pm$ and this determines a partial order on sign vectors coordinate-wisely, which coincides with the face order of $\mathcal{F}(\mathcal{A})$.
-We define the Tits product on $\mathcal{F}(\mathcal{A})$ using sign vectors as follow.
-For $F,G\in\mathcal{F}(\mathcal{A})$, let $FG\in\mathcal{F}(A)$ be the sign vector
+For a face $F\in\mathcal{F}(\mathcal{A})$, we say that the corresponding face $z_{F}$ of $Z(\mathcal{A})$ is dual to $F$.
+In particular, the 0-faces (vertices) of $Z(\mathcal{A})$ are dual to the chambers of $\mathcal{A}$.
+With these $\alpha_{H}$ chosen, each face $F\in\mathcal{F}(\mathcal{A})$ can be represented by a sign vector $F=(F_{H})_{H\in\mathcal{A}}\in\{+,-,0\}^{\mathcal{A}}$, where $F_{H}=\mathrm{sgn}(\alpha_{H}(F))$.
+Declaring $0<\pm$ determines a partial order on the set of sign vectors, which coincides with the face order of $\mathcal{F}(\mathcal{A})$.
+We define the Tits product on $\mathcal{F}(\mathcal{A})$ using sign vectors as follows.
+For $F,G\in\mathcal{F}(\mathcal{A})$, let $FG\in\mathcal{F}(\mathcal{A})$ be the sign vector
 $(FG)_{H}=\begin{cases}F_{H},\text{ if }F_{H}\neq 0;\\ G_{H},\text{ if }F_{H}=0.\end{cases}$
-This product does not depend on the choices of defining forms, since geometrically, $FG$ is the face we land in when we start from a relative interior point of $F$, move out a small enough distance towards a relative interior point of $G$ (see Figure 1).
+This product does not depend on the choice of defining forms: geometrically, $FG$ is the face reached by starting from a relative interior point of $F$ and moving a sufficiently small distance toward a relative interior point of $G$ (see Figure 1).
 […]
@@
 Let $F$ be a ray (blue) and $G$ be a chamber (red).
-(a) The product FGFG represents taking a small step from $F$ towards $G$, which immediately lands in the adjacent chamber.
-(b) The product GFGF represents taking a small step from $G$ towards $F$.
+(a) The product FGFG represents taking a small step from $F$ toward $G$, which immediately lands in the adjacent chamber.
+(b) The product GFGF represents taking a small step from $G$ toward $F$.
 Because $G$ is an open chamber, a sufficiently small step remains entirely within $G$, so $GF=G$.
 2.5. Tope graphs and graph-theoretic properties
-For an arrangement $\mathcal{A}$ in $\mathbb{R}^{\ell}$, notice that the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the $1$-skeleton of $Z(\mathcal{A})$.
-For a face $F\in\mathcal{F}(A)$, we let $\mathcal{T}(\mathcal{A})_{F}$ be the induced subgraph of $\mathcal{T}(\mathcal{A})$ with vertices $\mathrm{Ch}(\mathcal{A})_{F}=\{C\in\mathrm{Ch}(\mathcal{A})\mid F\leq C\}$, or equivalently, $\mathcal{T}(\mathcal{A})_{F}$ is the $1$-skeleton of the face $z_{F}$ of $Z(\mathcal{A})$.
+For an arrangement $\mathcal{A}$ in $\mathbb{R}^{d}$, note that the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the $1$-skeleton of $Z(\mathcal{A})$.
+For a face $F\in\mathcal{F}(\mathcal{A})$, we let $\mathcal{T}(\mathcal{A})_{F}$ be the induced subgraph of $\mathcal{T}(\mathcal{A})$ with vertices $\mathrm{Ch}(\mathcal{A})_{F}=\{C\in\mathrm{Ch}(\mathcal{A})\mid F\leq C\}$; equivalently, $\mathcal{T}(\mathcal{A})_{F}$ is the $1$-skeleton of the face $z_{F}$ of $Z(\mathcal{A})$.
 We call $\mathcal{T}(\mathcal{A})_{F}$ the $F$-subgraph of $\mathcal{T}(\mathcal{A})$.
@@
 Lemma 2.29.
-For $F\in\mathcal{F}(\mathcal{A})$, $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph of $\mathcal{T}(\mathcal{A})$.
+For $F\in\mathcal{F}(\mathcal{A})$, the subgraph $\mathcal{T}(\mathcal{A})_{F}$ of $\mathcal{T}(\mathcal{A})$ is gated with gate projection $D\mapsto FD$.
 Proof.
-We prove that the map $p_{F}:\mathcal{T}(\mathcal{A})\to\mathcal{T}(\mathcal{A})_{F}$ defined by $p_{F}(D)=FD$ is the gate projection, that is, we prove the equation
+We must show that the equality
 (2.4) $d(D,C)=d(D,FD)+d(FD,C)$
 holds for $D\in\mathrm{Ch}(\mathcal{A})$ and $C\in\mathrm{Ch}(\mathcal{A})_{F}$.
-In terms of sign vectors, the distance $d(X,Y)=\sum_{H\in\mathcal{A}}d(X,Y)_{H}$ is the Hamming distance (number of non-equal signs), where
+In terms of sign vectors, the distance $d(X,Y)=\sum_{H\in\mathcal{A}}d(X,Y)_{H}$ is the Hamming distance (the number of unequal signs), where
 $d(X,Y)_{H}=\begin{cases}1,\text{ if }X_{H}\neq Y_{H};\\ 0,\text{ if }X_{H}=Y_{H}.\end{cases}$
-Let us prove both sides of equation (2.3) have the same evaluation on each $H\in\mathcal{A}$.
-If $H\notin\mathcal{A}_{F}$, then $F_{H}\neq 0$ and $(FD)_{H}=F_{H}$, also since $C\in\mathrm{Ch}(\mathcal{A})_{F}$, we have $C_{H}=F_{H}$.
-Therefore (2.3) evaluated at $H$ is $d(D,C)_{H}=d(D,FD)_{H}+0$, which is trivially true since both sides equal $1$ if $D_{H}\neq F_{H}$ and 0 if $D_{H}=F_{H}$.
-Next suppose $H\in\mathcal{A}_{F}$, we have $F_{H}=0$.
-Then $d(D,FD)_{H}=0$.
-The equation (2.3) evaluated at $H$ is $d(D,C)_{H}=0+d(FD,C)_{H}$, which is triviall true since $D_{H}=(FD)_{H}$.
+We prove that both sides of equation (2.4) have the same contribution on each $H\in\mathcal{A}$.
+If $H\notin\mathcal{A}_{F}$, then $F_{H}\neq 0$ and $(FD)_{H}=F_{H}$; also, since $C\in\mathrm{Ch}(\mathcal{A})_{F}$, we have $C_{H}=F_{H}$.
+Therefore (2.4) evaluated at $H$ is $d(D,C)_{H}=d(D,FD)_{H}+0$, which is true since both sides equal $1$ if $D_{H}\neq F_{H}$ and 0 if $D_{H}=F_{H}$.
+Next suppose $H\in\mathcal{A}_{F}$.
+Then $F_{H}=0$ and $d(D,FD)_{H}=0$.
+The equation (2.4) evaluated at $H$ is $d(D,C)_{H}=0+d(FD,C)_{H}$, which is true since $D_{H}=(FD)_{H}$.
 ∎
-Here we record some important examples of tope graphs.
+We record some important examples of tope graphs.
 Example 2.30.
-Let $\mathcal{A}=Bl_{d}$ be the $d$-th Boolean arrangement, that is, the arrangement of $d$ coordinate hyperplanes of $\mathbb{R}^{d}$.
-It is clear that the tope graph $\mathcal{T}(\mathcal{A})$ is the whole hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{d}$.
+Let $\mathcal{A}=Bl_{d}$ be the $d$-th Boolean arrangement, that is, the arrangement of $d$ coordinate hyperplanes in $\mathbb{R}^{d}$.
+Then the tope graph $\mathcal{T}(\mathcal{A})$ is the whole hypercube graph $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{d}$.
 Example 2.31.
@@
 Generic arrangements are realizations of uniform matroids $U_{r,n}$.
-By abuse of notations, we use $U_{r,n}$ to denote a generic arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$.
+By abuse of notation, we use $U_{r,n}$ to denote a generic arrangement of $n$ hyperplanes in $\mathbb{R}^{r}$.
 For example, consider a generic arrangement $\mathcal{A}=U_{r,r+1}$.
-The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is isomorphic to the induced subgraph of $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{r+1}$ with all vertices (sign vectors) except the all ++ vector and the all - vector.
+The tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the induced subgraph of $Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}\cong Q_{r+1}$ obtained by removing the all-++ vector and the all-- vector.
 Example 2.32.
 Consider a finite Coxeter system $(W,S)$ and the associated Coxeter arrangement $\mathcal{A}_{W}$.
-The zonotope $Z(\mathcal{A}_{W})$ is the $W$-permutohedron, and the tope graph $\mathcal{T}(\mathcal{A})$ is isomorphic to the Cayley graph of $(W,S)$.
-Tope graphs have the following graph-theoretic properties.
-Recall that a chamber $C$ is simplicial if its intersection with the unit sphere of the ambient space is combinatorially a simplex.
-An arrangement is simplicial if all chambers all simplicial.
-Proposition 2.30 (Propositions 4.2.15 and 4.4.8 of [4]).
-Let $\mathcal{A}$ be an arrangement of $\mathrm{rank}\mathcal{A}=r$. The tope graph $\mathcal{T}=\mathcal{T}(\mathcal{A})$ has the following properties as an unlabeled graph.
+The zonotope $Z(\mathcal{A}_{W})$ is the $W$-permutahedron, and the tope graph $\mathcal{T}(\mathcal{A}_{W})$ is isomorphic to the Cayley graph of $(W,S)$.
+Tope graphs have the following graph-theoretic properties. Recall that a chamber $C$ is simplicial if it is a simplicial cone. An arrangement is simplicial if all its chambers are simplicial.
+Proposition 2.33 ([4, Propositions 4.2.15 and 4.4.8]).
+Let $\mathcal{A}$ be an arrangement of $\operatorname{rank}\mathcal{A}=r$. The tope graph $\mathcal{T}=\mathcal{T}(\mathcal{A})$ has the following properties as an unlabeled graph.
   1. (1)
-    For each vertex $C$, there is a unique vertex C-C such that $d(C,-C)=\mathrm{diam}(\mathcal{T})$.
+    For each vertex $C$, there is a unique vertex C-C such that $d(C,-C)=\operatorname{diam}(\mathcal{T})$.
   2. (2)
-    The mapping $C\mapsto-C$ is a fixed point free automorphism of $\mathcal{T}$.
+    The mapping $C\mapsto-C$ is a fixed-point-free automorphism of $\mathcal{T}$.
   3. (3)
@@
   4. (4)
-    For each vertex $C$, $\deg(C)\geq r$, where the equality holds iff $C$ is simplicial.
+    For each vertex $C$, we have $\deg(C)\geq r$, where equality holds if and only if $C$ is simplicial.
   5. (5)
     $\mathcal{T}$ is $r$-connected.
-Recall that a graph is called $k$-regular if every vertex has degree $k$.
-Corollary 2.31.
-$\mathcal{A}$ is simplicial of rank $r$ if and only if $\mathcal{T}(\mathcal{A})$ is $r$-regular.
-A matching of a graph is a collection of edges where no two edges share a common incident vertex.
-We say a matching is perfect if every vertex of the graph is incident to exactly one edge in the matching.
-The marriage theorem (see Theorem 2.1.2 of [9]) has a corollary (see Corollary 2.1.3 of [9]) asserting that every $k$-regular ($k\geq 1$) bipartite graph has a perfect matching.
-Corollary 2.32.
-If $\mathcal{A}$ is simplicial, then $\mathcal{T}(\mathcal{A})$ has a perfect matching.
-Very recently, Hamiltonian cycles in tope graphs of simplicial and supersolvable arrangements are studied in [5, 20].
 2.6. Varchenko matrix
-Varchenko introduced a bilinear form and computed the determinant [31]. Generalization to oriented matroids can be found in [17]. See [8] and Sections 8.4-8.5 of [2] for details.
-Here we only consider the equal weight case. For a real hyperplane arrangement $\mathcal{A}$, we define the $q$-Varchenko matrix
+Varchenko introduced a bilinear form associated with an arrangement and computed its determinant [33].
+Generalizations to oriented matroids can be found in [18].
+See [8] and [2, Sections 8.4–8.5] for details.
+Here we only consider the equal-weight case. For a real hyperplane arrangement $\mathcal{A}$, we define the $q$-Varchenko matrix by
 $\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]_{C,D\in\mathrm{Ch}(\mathcal{A})}.$
-In other words, $\mathcal{V}_{q}(\mathcal{A})$ is the Zeta matrix of the tope graph $\mathcal{T}(\mathcal{A})$ (see Section 2.2). It can also be defined for oriented matroids.
-Theorem 2.33 ([31], see also Theorem 8.23 of [2]).
+In other words, $\mathcal{V}_{q}(\mathcal{A})$ is the Zeta matrix of the tope graph $\mathcal{T}(\mathcal{A})$ (see Section 2.2). This matrix can also be defined for oriented matroids.
+Theorem 2.34 ([33], see also [2, Theorem 8.23]).
 The determinant of the $q$-Varchenko matrix is
-$\det\mathcal{V}_{q}(\mathcal{A})=\prod_{X\in L(\mathcal{A})\setminus\{\hat{0}\}}(1-q^{2|\mathcal{A}_{X}|})^{c^{X}\beta_{X}}$
+$\det\mathcal{V}_{q}(\mathcal{A})=\prod_{X\in L(\mathcal{A})\setminus\{\hat{0}\}}(1-q^{2|\mathcal{A}_{X}|})^{c^{X}\beta_{X}}.$
 2.7. Tope poset and generating function
-Fix a base chamber $B\in\mathrm{Ch}(\mathcal{A})$ and define a partial order on $\mathrm{Ch}(\mathcal{A})$ by $C\preceq C^{\prime}$ if $S(B,C)\subseteq S(B,C^{\prime})$.
+Fix a chamber $B\in\mathrm{Ch}(\mathcal{A})$ and define a partial order on $\mathrm{Ch}(\mathcal{A})$ by $C\preceq C^{\prime}$ if $S(B,C)\subseteq S(B,C^{\prime})$.
 This poset is called the tope poset of $\mathcal{A}$ based at $B$, denoted by $\mathcal{T}(\mathcal{A},B)$.
-The poset $\mathcal{T}(\mathcal{A},B)$ has bottom element $B$ and top element B-B, graded by rank function $\rho_{B}(C)=d(B,C)$.
-In terms of notations in Section 2.1, $\mathcal{T}(\mathcal{A},B)=[B,-B]_{\mathcal{T}(\mathcal{A})}$.
-Proposition 2.34 (Theorem 2.2 of [10], see also Theorem 4.4.2 of [4]).
-An open interval $(C_{1},C_{2})$ inside $\mathcal{T}(\mathcal{A},B)$ is homotopy equivalent to a sphere $S^{\mathrm{rank}\mathcal{A}-\dim F-2}$, if the closed interval $C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$; contractible otherwise.
-Corollary 2.35 (Theorem 1.11 of [11], see also Corollary 4.4.3 of [4]).
-The Möbius function on $\mathcal{T}(\mathcal{A},B)$ is
-$\mu(C_{1},C_{2})=\begin{cases}(-1)^{\mathrm{rank}\mathcal{A}-\dim F},&\text{ if }[C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F},\exists F\in\mathcal{F}(\mathcal{A});\\ 0,&\text{ otherwise}.\end{cases}$
-Proposition 2.36 (Theorem 3.4 of [3]).
+The poset $\mathcal{T}(\mathcal{A},B)$ has bottom element $B$ and top element B-B, graded by the rank function $\rho_{B}(C)=d(B,C)$.
+In the notation of Section 2.1, $\mathcal{T}(\mathcal{A},B)=[B,-B]_{\mathcal{T}(\mathcal{A})}$.
+Proposition 2.35 ([9, Theorem 2.2], see also [4, Theorem 4.4.2]).
+An open interval $(C_{1},C_{2})$ in $\mathcal{T}(\mathcal{A},B)$ is homotopy equivalent to a sphere $S^{\operatorname{rank}\mathcal{A}-\dim F-2}$ if $C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$, and is contractible otherwise.
+Corollary 2.36 ([10, Theorem 1.11], see also [4, Corollary 4.4.3]).
+The Möbius function on $\mathcal{T}(\mathcal{A},B)$ is given by
+$\mu(C_{1},C_{2})=\begin{cases}(-1)^{\operatorname{rank}\mathcal{A}-\dim F},&\text{ if }[C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F}\text{ for some }F\in\mathcal{F}(\mathcal{A});\\ 0,&\text{ otherwise}.\end{cases}$
+Proposition 2.37 ([3, Theorem 3.4]).
 If $\mathcal{A}$ is simplicial, then $\mathcal{T}(\mathcal{A},B)$ is a lattice for any $B\in\mathrm{Ch}(\mathcal{A})$.
-Let $D_{\mathcal{A},B}(q)=\sum_{C\in\mathrm{Ch(\mathcal{A})}}q^{d(B,C)}$ be the rank generation function of $\mathcal{T}(\mathcal{A},B)$.
-Note that it is the sum of entries of the $B$-row/column of the $q$-Varchenko matrix $\mathcal{V}_{q}(\mathcal{A})$.
+Let $D_{\mathcal{A},B}(q)=\sum_{C\in\mathrm{Ch}(\mathcal{A})}q^{d(B,C)}$ be the rank-generating function of $\mathcal{T}(\mathcal{A},B)$.
+Note that it is the sum of the entries in the $B$-row/column of the $q$-Varchenko matrix $\mathcal{V}_{q}(\mathcal{A})$.
 Proposition 2.38 ([30]).
-If $\mathcal{A}$ is a Coxeter arrangement with exponents $e_{1},\ldots,e_{\ell}$, then
+If $\mathcal{A}$ is a Coxeter arrangement with exponents $e_{1},\ldots,e_{\ell}$, then we have
 $D_{\mathcal{A},B}(q)=\prod_{i=1}^{\ell}[1+e_{i}]_{q}$
-for every $B\in\mathrm{Ch}(\mathcal{A})$, where $n]_{q}=1+q+q^{2}+\cdots+q^{n-1}$ is the $q$-analog.
-The generating function decomposition formula is generalized to supersolvable arrangements [3], inductively factored arrangements [18]. See also [1, 26, 12].
-3. Magnitude of tope graphs
-For a real hyperplane arrangement $\mathcal{A}$, we define the magnitude $\mathrm{Mag}(\mathcal{A})=\mathrm{Mag}(\mathcal{A},q)$ of $\mathcal{A}$ as the magnitude of $\mathcal{T}(\mathcal{A})$.
+for every $B\in\mathrm{Ch}(\mathcal{A})$, where $n]_{q}=1+q+q^{2}+\cdots+q^{n-1}$ is the $q$-number.
+The generating-function decomposition formula has been generalized to supersolvable arrangements [3] and inductively factored arrangements [19]. See also [1, 27, 12].
+3. Magnitude of arrangements
+For a real central hyperplane arrangement $\mathcal{A}$, we define the magnitude $\mathrm{Mag}(\mathcal{A})=\mathrm{Mag}(\mathcal{A},q)$ of $\mathcal{A}$ as the magnitude of its tope graph $\mathcal{T}(\mathcal{A})$.
 3.1. Basic properties
-We first prove a reciprocity result.
+We first prove the symmetry of the magnitude with respect to $q\leftrightarrow q^{-1}$.
 Proposition 3.1.
-Let $\mathcal{A}$ be a central hyperplane arrangement with $N=\#\mathcal{A}$ hyperplanes. Then
+Let $\mathcal{A}$ be an arrangement with $N=\#\mathcal{A}$ hyperplanes. Then
 $\mathrm{Mag}(\mathcal{A},q^{-1})=q^{N}\mathrm{Mag}(\mathcal{A},q).$
@@
 For any two chambers $C,D\in\mathrm{Ch}(\mathcal{A})$, we have d(C,D)+d(C,D)=Nd(C,D)+d(C,-D)=N.
-Then the $q$-Varchenko matrix $\mathcal{V}_{q}=\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]$ satisfies
+Thus the $q$-Varchenko matrix $\mathcal{V}_{q}=\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]$ satisfies
 $q^{N}\mathcal{V}_{q^{-1}}=[q^{N-d(C,D)}]=[q^{d(C,-D)}]=\mathcal{V}_{q}P_{\tau},$
-where $P_{\tau}$ is the permutation matrix induced by $\tau$. By definition of magnitude,
+where $P_{\tau}$ is the permutation matrix induced by $\tau$. By the definition of magnitude, we have
 $\mathrm{Mag}(\mathcal{A},q^{-1})=\mathbf{1}^{T}(\mathcal{V}_{q^{-1}})^{-1}\mathbf{1}=\mathbf{1}^{T}(q^{-N}\mathcal{V}_{q}P_{\tau})^{-1}\mathbf{1}=q^{N}\mathbf{1}^{T}P_{\tau}^{-1}\mathcal{V}_{q}^{-1}\mathbf{1}=q^{N}\mathbf{1}^{T}\mathcal{V}_{q}^{-1}\mathbf{1}=q^{N}\mathrm{Mag}(\mathcal{A},q),$
 where we used $\mathbf{1}^{T}P_{\tau}^{-1}=\mathbf{1}^{T}$. ∎
-Since tope graphs are partial cubes, the magnitude $\mathrm{Mag}(\mathcal{A})$ has the one-point property by [22].
+Since tope graphs are partial cubes, the magnitude $\mathrm{Mag}(\mathcal{A})$ satisfies the one-point property established in [22].
 Proposition 3.2.
-Let $\mathcal{A}$ be a central hyperplane arrangement, then $\mathrm{Mag}(\mathcal{A},1)=1$.
-Next we prove some main structural results of $\mathrm{Mag}(\mathcal{A})$.
+If $\mathcal{A}$ is an arrangement, then $\mathrm{Mag}(\mathcal{A},1)=1$.
+Next we prove some structural results for $\mathrm{Mag}(\mathcal{A})$.
 Theorem 3.3.
-For central arrangement $\mathcal{A}$, let $\mathrm{Mag}(\mathcal{A},q)=P(q)/Q(q)$ be the reduced form, with the leading coefficient of $Q(q)$ positive. Then
+For an arrangement $\mathcal{A}$, let $\mathrm{Mag}(\mathcal{A},q)=P(q)/Q(q)$ be the reduced expression of the magnitude, with the leading coefficient of $Q(q)$ positive. Then
   1. (1)
@@
   3. (3)
-    Roots of $Q$ are all roots of unity, excluding $1$.
+    All roots of $Q$ are roots of unity other than $1$.
 Proof.
@@
 Since $\mathrm{Mag}(\mathcal{A},0)=a_{0}/b_{0}=\#\mathrm{Ch}(\mathcal{A})$ (Proposition 2.7), we have $a_{0}\neq 0$ and $b_{0}\neq 0$.
-By the reciprocity formula (Proposition 3.1),
+By the symmetry formula (Proposition 3.1), we have
 $q^{N}=\frac{\mathrm{Mag}(\mathcal{A},q^{-1})}{\mathrm{Mag}(\mathcal{A},q)}=\frac{P(q^{-1})Q(q)}{P(q)Q(q^{-1})}.$
-Multiplying both sides by $q^{m-n}$, we have
+Multiplying both sides by $q^{m-n}$, we obtain
 $q^{N+m-n}=\frac{q^{m}P(q^{-1})Q(q)}{P(q)q^{n}Q(q^{-1})}=\frac{(a_{m}+\cdots+a_{0}q^{m})(b_{n}q^{n}+\cdots+b_{0})}{(a_{m}q^{m}+\cdots+a_{0})(b_{n}+\cdots+b_{0}q^{n})}.$
-Comparing degrees, we conclude N+mn=0N+m-n=0 and (1) is proved.
-Let us write $P^{*}(q)=q^{m}P(q^{-1})=a_{m}+\cdots+a_{0}q^{m}$ and $Q^{*}(q)=q^{n}Q(q^{-1})=b_{n}+\cdots+b_{0}q^{n}$ as the reverse polynomials. We have proved
+Comparing degrees, we obtain N+mn=0N+m-n=0, and (1) is proved.
+Write $P^{*}(q)=q^{m}P(q^{-1})=a_{m}+\cdots+a_{0}q^{m}$ and $Q^{*}(q)=q^{n}Q(q^{-1})=b_{n}+\cdots+b_{0}q^{n}$ for the reverse polynomials. We have proved
 $\mathrm{Mag}(\mathcal{A},q)=\frac{P(q)}{Q(q)}=\frac{P^{*}(q)}{Q^{*}(q)}.$
-Note that the fraction $P^{*}(q)/Q^{*}(q)$ is reduced since if $P^{*}$ and $Q^{*}$ share a common root $\alpha\neq 0$, $P$ and $Q$ would share a common root $1/\alpha$, contradicting to the assumption $P/Q$ is reduced.
-Now we must have $P^{*}(q)=rP(q)$ and $Q^{*}(q)=rQ(q)$ for some constant $r\neq 0$.
+Note that the fraction $P^{*}(q)/Q^{*}(q)$ is reduced, since if $P^{*}$ and $Q^{*}$ share a common root $\alpha\neq 0$, then $P$ and $Q$ would share the common root $1/\alpha$, contradicting the assumption that $P/Q$ is reduced.
+Hence we must have $P^{*}(q)=rP(q)$ and $Q^{*}(q)=rQ(q)$ for some constant $r\neq 0$.
 Substituting $q=1$ into the equations $Q^{*}(q)=q^{n}Q(1/q)$ and $Q^{*}(q)=rQ(q)$, we obtain
 $Q^{*}(1)=Q(1)=rQ(1).$
-Since $\mathrm{Mag}(\mathcal{A},1)=P(1)/Q(1)=1$ (Proposition 3.2) and $P/Q$ is reduced, $Q(1)\neq 0$, we conclude $r=1$. Then $P^{*}=P$ and $Q^{*}=Q$, so (2) is proved.
-Since $\mathcal{V}_{q}^{-1}=\frac{1}{\det\mathcal{V}_{q}}\mathrm{adj~}\mathcal{V}_{q}$, roots of $Q(q)$ are all roots of $\det\mathcal{V}_{q}$, which are roots of unity by Theorem 2.33.
-We have already seen $Q(1)\neq 0$.
-So (3) is proved.
+Since $\mathrm{Mag}(\mathcal{A},1)=P(1)/Q(1)=1$ (Proposition 3.2) and $P/Q$ is reduced, we have $Q(1)\neq 0$, so $r=1$. Then $P^{*}=P$ and $Q^{*}=Q$, and (2) is proved.
+Since $\mathcal{V}_{q}^{-1}=\frac{1}{\det\mathcal{V}_{q}}\mathrm{adj~}\mathcal{V}_{q}$, the roots of $Q(q)$ are roots of $\det\mathcal{V}_{q}$, which are roots of unity by Theorem 2.34.
+We have already seen that $Q(1)\neq 0$.
+Thus (3) is proved.
 ∎
-If an arrangement $\mathcal{A}$ decomposes as a direct sum $\mathcal{A}_{1}\oplus\mathcal{A}_{2}$, then the tope graph decomposes as cartesian product $\mathcal{T}(\mathcal{A})=\mathcal{T}(\mathcal{A}_{1})\square\mathcal{T}(\mathcal{A}_{2})$.
-So by Proposition 2.5, we have a decomposition of magnitude.
+If an arrangement $\mathcal{A}$ decomposes as a direct sum $\mathcal{A}_{1}\oplus\mathcal{A}_{2}$, then the tope graph decomposes as the Cartesian product $\mathcal{T}(\mathcal{A})=\mathcal{T}(\mathcal{A}_{1})\square\mathcal{T}(\mathcal{A}_{2})$.
+Therefore, by Proposition 2.5, we have a decomposition of magnitude.
 Proposition 3.4.
 $\mathrm{Mag}(\mathcal{A}_{1}\oplus\mathcal{A}_{2})=\mathrm{Mag}(\mathcal{A}_{1})\mathrm{Mag}(\mathcal{A}_{2})$.
-Coxeter arrangements have the natural action of the Coxeter groups on the chambers, which makes the magnitude easy to compute.
+Coxeter arrangements carry a natural action of the corresponding Coxeter groups on the chambers, which makes their magnitudes easy to compute.
 Proposition 3.5.
-For Coxeter arrangement $\mathcal{A}_{W}$ with exponents $\{e_{1},\ldots,e_{\ell}\}$, we have
+For a Coxeter arrangement $\mathcal{A}_{W}$ with exponents $\{e_{1},\ldots,e_{\ell}\}$, we have
 $\mathrm{Mag}(\mathcal{A}_{W})=\frac{\#W}{\prod_{i}[1+e_{i}]_{q}}.$
 Proof.
-The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is the $1$-skeleton of the $W$-permutohedron, or equivalently the Cayley graph of $W$ with respect to the standard generating set (Example 2.29), hence is vertex-transitive.
+The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is the $1$-skeleton of the $W$-permutahedron, or equivalently the Cayley graph of $W$ with respect to the standard generating set (Example 2.32), and hence is vertex-transitive.
 The result follows from Proposition 2.3 and Proposition 2.38.
 ∎
-3.2. Face decomposition conjecture
-For the purpose of reciprocity, let us define the following.
+3.2. Face decomposition formula
+For reciprocity purposes, let us define the following.
 Definition 3.6.
 We define the interior magnitude of $\mathcal{A}$ by
-$\mathrm{Mag}^{\circ}(\mathcal{A},q):=(-1)^{\mathrm{rank}\mathcal{A}}\mathrm{Mag}(\mathcal{A},q^{-1})=(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}\mathrm{Mag}(\mathcal{A},q),$
-where the latter equality is Proposition 3.1.
-Let us propose a main conjecture of the magnitude of arrangements.
-Conjecture 3.7.
-We have a face decomposition of the magnitude.
-$\mathrm{Mag}(\mathcal{A})$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\mathrm{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})$
-$=\sum_{X\in L(\mathcal{A})}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})=\sum_{X\in L(\mathcal{A})}(-1)^{\mathrm{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X}),$
-where $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$. Equivalently,
-$\mathrm{Mag}(\mathcal{A})$ $=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\mathrm{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}$
-$=\frac{\sum_{X\in L(\mathcal{A})\setminus\{\cap\mathcal{A}\}}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{X\in L(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\mathrm{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}.$
+$\mathrm{Mag}^{\circ}(\mathcal{A},q):=(-1)^{\operatorname{rank}\mathcal{A}}\mathrm{Mag}(\mathcal{A},q^{-1})=(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}\mathrm{Mag}(\mathcal{A},q),$
+where the latter equality follows from Proposition 3.1.
+We prove the following “face decomposition formula” for the magnitude of an arrangement.
+Theorem 3.7.
+For any arrangement $\mathcal{A}$, we have
+$\mathrm{Mag}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F}).$
 Remark 3.8.
-This formula is the discrete $q$-analogue of the decomposition of the Euler characteristic of a zonotope into its open faces.
-Conjecture 3.7 is verified by computer for all examples listed in Section 3.3 and the appendix, and by easy computations for Boolean arrangements and all rank $2$ arrangements.
+Theorem 3.7 can be reformulated as
+$\mathrm{Mag}(\mathcal{A})=\sum_{X\in L(\mathcal{A})}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})=\sum_{X\in L(\mathcal{A})}(-1)^{\operatorname{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X}),$
+where $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$. It can also be written as
+$\mathrm{Mag}(\mathcal{A})=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})}{1-(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})}{1-(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}}.$
+This can be regarded as a recursive formula for computing $\mathrm{Mag}(\mathcal{A})$.
+Remark 3.9.
+Taking the limit $q\to 1$ gives $\mathrm{Mag}(\mathcal{A},1)=1$ and $\mathrm{Mag}^{\circ}(\mathcal{A},1)=(-1)^{\operatorname{rank}\mathcal{A}}$ by Proposition 3.2.
+Therefore, Theorem 3.7 can be regarded as a $q$-analogue of the decomposition of the Euler characteristic of the zonotope into its open faces.
 Corollary 3.10.
-Conjecture 3.7 implies that $\mathrm{Mag}(\mathcal{A})$ is determined by $L(\mathcal{A})$.
+The magnitude $\mathrm{Mag}(\mathcal{A})$ of an arrangement $\mathcal{A}$ is determined by $L(\mathcal{A})$.
+Lemma 3.11.
+For any arrangement $\mathcal{A}$ and $D\in\mathrm{Ch}(\mathcal{A})$, we have
+$\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}+d(D,FD)}=1.$
+Proof.
+If $\operatorname{rank}\mathcal{A}=0$, then $\mathcal{A}$ is the empty arrangement and there is nothing to prove, so let us assume $r:=\operatorname{rank}\mathcal{A}\geq 1$.
+Passing to the essentialization does not change the chamber set, the tope graph, or the quantities $\mathcal{A}_{F}$ and $\operatorname{codim}F=\operatorname{rank}(\mathcal{A}_{F})$, so we may work in an essential arrangement of rank $r$.
+Choose defining forms $\alpha_{H}$ so that $D_{H}=+$ for every $H\in\mathcal{A}$. For a face $F$, define $M_{D}(F)\subset\mathcal{A}$ by
+$M_{D}(F):=\{H\in\mathcal{A}:F_{H}\neq+\}.$
+Since $F_{H}=0$ exactly when $H\in\mathcal{A}_{F}$, while $d(D,FD)$ counts the hyperplanes on which $F_{H}=$-, we have $M_{D}(F)=\#\mathcal{A}_{F}+d(D,FD)$. Therefore it is enough to prove
+$S_{D}(q):=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#M_{D}(F)}=1.$
+Expand each power as
+$q^{\#M_{D}(F)}=(1+(q-1))^{\#M_{D}(F)}=\sum_{T\subseteq M_{D}(F)}(q-1)^{\#T}.$
+After exchanging the order of summation we obtain
+$S_{D}(q)=\sum_{T\subseteq\mathcal{A}}(q-1)^{\#T}E_{T},\qquad E_{T}:=\sum_{F:\,T\subseteq M_{D}(F)}(-1)^{\operatorname{codim}F}.$
+Hence the desired identity will follow once we show $E_{T}=\mathbf{1}_{T=\varnothing}$, i.e.
+$E_{T}=0$ if $T\neq\varnothing$ and $E_{\varnothing}=1$.
+We interpret $E_{T}$ as a reduced Euler characteristic.
+Let $S^{r-1}$ be the unit sphere in the ambient space $V=\mathbb{R}^{r}$.
+For each hyperplane $H\in\mathcal{A}$, let
+$H_{D}^{-}:=\{x\in S^{r-1}:\alpha_{H}(x)\leq 0\}$
+be the closed hemisphere opposite to $D$, and for $T\subseteq\mathcal{A}$ define
+$\Sigma_{T}:=\bigcap_{H\in T}H_{D}^{-}\subseteq S^{r-1},\qquad\Sigma_{\varnothing}:=S^{r-1}.$
+The arrangement induces a regular CW decomposition of $S^{r-1}$ where the open cells are $F\cap S^{r-1}$ for non-minimal faces $F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}$.
+An open cell $F\cap S^{r-1}$ lies in $\Sigma_{T}$ if and only if $T\subseteq M_{D}(F)$.
+Therefore
+$\chi(\Sigma_{T})=\sum_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}\\ T\subseteq M_{D}(F)\end{subarray}}(-1)^{\dim F-1}.$
+Since $\operatorname{codim}F=r-\dim F$ and the minimal face $\cap\mathcal{A}$ contributes $(-1)^{r}$ to $E_{T}$, we obtain
+$E_{T}$ $=(-1)^{r}+\sum_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}\\ T\subseteq M_{D}(F)\end{subarray}}(-1)^{r-\dim F}=(-1)^{r-1}\left(\chi(\Sigma_{T})-1\right).$
+If $T=\varnothing$, then $\Sigma_{\varnothing}=S^{r-1}$, so $E_{\varnothing}=(-1)^{r-1}(\chi(S^{r-1})-1)=1$.
+If $T\neq\varnothing$, then $\Sigma_{T}$ is the intersection of a nonempty collection of closed hemispheres, so $\Sigma_{T}$ is homeomorphic to $D^{r-1}$.
+Therefore $E_{T}=(-1)^{r-1}(\chi(D^{r-1})-1)=0$.
+This shows $E_{T}=\mathbf{1}_{T=\varnothing}$.
+∎
+Proof of Theorem 3.7.
+Fix a face $F\in\mathcal{F}(\mathcal{A})$.
+Let $w_{F}\colon\mathrm{Ch}(\mathcal{A})_{F}\to\mathbb{Q}(q)$ be the weight vector of $\mathcal{T}(\mathcal{A})_{F}\cong\mathcal{T}(\mathcal{A}_{F})$, so that
+$\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}q^{d(E,C)}w_{F}(C)=1$
+for every $E\in\mathrm{Ch}(\mathcal{A})_{F}$.
+Extend $w_{F}$ by zero outside $\mathrm{Ch}(\mathcal{A})_{F}$.
+Now define a function $\widetilde{w}_{\mathcal{A}}\colon\mathrm{Ch}(\mathcal{A})\to\mathbb{Q}(q)$ by
+$\widetilde{w}_{\mathcal{A}}(C):=\sum_{F\leq C}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}w_{F}(C).$
+Taking the sum over all chambers gives
+$\sum_{C\in\mathrm{Ch}(\mathcal{A})}\widetilde{w}_{\mathcal{A}}(C)$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}w_{F}(C)$
+$=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F},q)$
+$=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F},q).$
+Therefore Theorem 3.7 will follow once we prove that $\widetilde{w}_{\mathcal{A}}$ is the actual weight vector of $\mathcal{A}$.
+Fix a chamber $D\in\mathrm{Ch}(\mathcal{A})$.
+We compute the $D$-th coordinate of $\mathcal{V}_{q}(\mathcal{A})\widetilde{w}_{\mathcal{A}}$:
+$\sum_{C\in\mathrm{Ch}(\mathcal{A})}q^{d(D,C)}\widetilde{w}_{\mathcal{A}}(C)$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}q^{d(D,C)}w_{F}(C).$
+Since $\mathcal{T}(\mathcal{A})_{F}$ is gated with gate projection $D\mapsto FD$ (Lemma 2.29), we have
+$d(D,C)=d(D,FD)+d(FD,C)$
+for every $C\in\mathrm{Ch}(\mathcal{A})_{F}$. Using this identity and the defining equation for the local weight vector $w_{F}$, we obtain
+$\sum_{C\in\mathrm{Ch}(\mathcal{A})}q^{d(D,C)}\widetilde{w}_{\mathcal{A}}(C)$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}+d(D,FD)}\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}q^{d(FD,C)}w_{F}(C)$
+$=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}+d(D,FD)}.$
+By Lemma 3.11, this equals $1$ for every $D\in\mathrm{Ch}(\mathcal{A})$. ∎
 3.3. Examples
-We use the notion $\mathcal{A}(n,k)$ from Grünbaum’s catalogue [14].
+We use the notation $\mathcal{A}(n,k)$ from Grünbaum’s catalogue [15].
 Example 3.12.
-The near-pencil $\mathcal{A}(n,0)$ has tope graph $C_{2(n-1)}\Box K_{2}$ and hence
+The near-pencil $\mathcal{A}(n,0)$ has the tope graph $C_{2(n-1)}\Box K_{2}$ and hence
 $\mathrm{Mag}(\mathcal{A}(n,0))=\frac{4(n-1)}{[2]_{q}^{2}[n-1]_{q}}.$
 Example 3.13.
-Note that $\mathcal{A}(6,1)$ is the Coxeter arrangement of type $A_{3}$ or $D_{3}$, and $\mathcal{A}(9,1)$ is the Coxeter arrangement of type $B_{3}$, which can be computed from Proposition 3.5.
+Note that $\mathcal{A}(6,1)$ is the Coxeter arrangement of type $A_{3}$ or $D_{3}$, and $\mathcal{A}(9,1)$ is the Coxeter arrangement of type $B_{3}$.
+Therefore, the magnitudes of both arrangements can be computed from Proposition 3.5.
 In what follows, $\Phi_{k}(q)\in\mathbb{Z}[q]$ is the $k$-th cyclotomic polynomial.
@@
 $=48-144q+192q^{2}-192q^{3}+240q^{4}-336q^{5}+432q^{6}-528q^{7}+624q^{8}-720q^{9}+816q^{10}+\cdots$
-More examples are in the appendix.
-4. Magnitude homology of tope graphs
-We continue denoting invariants of $\mathcal{T}(\mathcal{A})$ simply by invariants of $\mathcal{A}$, such as $MH_{k,\ell}(\mathcal{A}):=MH_{k,\ell}(\mathcal{T}(\mathcal{A}))$, $\beta_{k,\ell}(\mathcal{A}):=\mathrm{rank}MH_{k,\ell}(\mathcal{A})$ and $\chi_{\ell}(\mathcal{A})=\sum_{k}(-1)^{k}\beta_{k,\ell}(\mathcal{A})$.
+Additional examples are given in the appendix.
+4. Computations for rank 3 arrangements
+4.1. Magnitude of rank 3 arrangements
+Using the face decomposition formula (Theorem 3.7), we can explicitly compute the magnitude of rank $3$ arrangements.
+Let $\mathcal{A}$ be an arrangement of rank $3$ with $n=\#\mathcal{A}$.
+For $k\geq 2$, we write $n_{k}$ for the number of faces $F\in\mathcal{F}(\mathcal{A})$ with $\operatorname{codim}F=2$ and $\mathcal{A}_{F}=k$.
+Theorem 4.1.
+Let $\mathcal{A}$ be an arrangement of rank $3$ with $\mathcal{A}=n$. Then we have
+$\mathrm{Mag}(\mathcal{A})=\dfrac{1}{1+q^{n}}\biggl(\#\mathrm{Ch}(\mathcal{A})-\sum_{k\geq 2}\dfrac{2kn_{k}q(1-q^{k-1})}{(1+q)(1-q^{k})}\biggr).$
+Proof.
+Let $\mathcal{F}_{d}(\mathcal{A})$ denote the set of faces of codimension $d$ of $\mathcal{A}$. By Theorem 3.7, we have
+$\mathrm{Mag}(\mathcal{A})$ ${}=\sum_{F\in\mathcal{F}_{0}(\mathcal{A})}1-\sum_{F\in\mathcal{F}_{1}(\mathcal{A})}\dfrac{2q}{1+q}+\sum_{F\in\mathcal{F}_{2}(\mathcal{A})}\dfrac{2\#\mathcal{A}_{F}q^{\#\mathcal{A}_{F}}(1-q)}{(1+q)(1-q^{\#\mathcal{A}_{F}})}-q^{n}\mathrm{Mag}(\mathcal{A},q)$
+${}=\#\mathrm{Ch}(\mathcal{A})-\#\mathcal{F}_{1}(\mathcal{A})\cdot\dfrac{2q}{1+q}+\sum_{k\geq 2}\dfrac{2kn_{k}q^{k}(1-q)}{(1+q)(1-q^{k})}-q^{n}\mathrm{Mag}(\mathcal{A},q).$
+For each $F\in\mathcal{F}_{2}(\mathcal{A})$, the number of faces $G\in\mathcal{F}_{1}(\mathcal{A})$ of codimension $1$ adjacent to $F$ is exactly $2\#\mathcal{A}_{F}$.
+Conversely, for each $G\in\mathcal{F}_{1}(\mathcal{A})$, there are exactly two faces $F\in\mathcal{F}_{2}(\mathcal{A})$ of codimension $2$ adjacent to $G$.
+Therefore we have
+$\mathcal{F}_{1}(\mathcal{A})=\dfrac{1}{2}\sum_{F\in\mathcal{F}_{2}(\mathcal{A})}2\#\mathcal{A}_{F}=\sum_{k\geq 2}kn_{k}.$
+Using this identity, we obtain
+$(1+q^{n})\mathrm{Mag}(\mathcal{A})$ ${}=\#\mathrm{Ch}(\mathcal{A})-\sum_{k\geq 2}\dfrac{2kn_{k}q}{1+q}+\sum_{k\geq 2}\dfrac{2kn_{k}q^{k}(1-q)}{(1+q)(1-q^{k})}$
+${}=\#\mathrm{Ch}(\mathcal{A})-\sum_{k\geq 2}\dfrac{2kn_{k}q(1-q^{k-1})}{(1+q)(1-q^{k})}.$
+This completes the proof. ∎
+Example 4.2.
+The arrangement $\mathcal{A}(6,1)$ from Grünbaum’s catalogue [15] satisfies
+$\mathrm{Ch}(\mathcal{A}(6,1))=24\quad\text{and}\quad n_{k}=\begin{cases}6,&\text{ if }k=2;\\ 8,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4.\end{cases}$
+By Theorem 4.1, its magnitude is given by
+$\mathrm{Mag}(\mathcal{A}(6,1))$ ${}=\dfrac{1}{1+q^{6}}\biggl(24-\dfrac{24q}{(1+q)^{2}}-\dfrac{48q}{1+q+q^{2}}\biggr)$
+${}=\frac{24}{(1+q)^{2}(1+q+q^{2})(1+q^{2})}.$
+This coincides with the formula given in Example 3.13.
+4.2. Alternating sign property
+In all the examples shown in Appendix A, the coefficients in the series expansion of the magnitude of each arrangement seem to have alternating signs. That is, if we write
+$\mathrm{Mag}(\mathcal{A})=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell},$
+then $(-1)^{\ell}c_{\ell}\geq 0$ appears to hold.
+Here we prove that if $\operatorname{rank}\mathcal{A}\leq 3$, then $(-1)^{\ell}c_{\ell}>0$ holds for all sufficiently large $\ell$.
+On the other hand, in the next subsection, we construct an example for which the inequality $(-1)^{\ell}c_{\ell}\geq 0$ fails for small $\ell$.
+Definition 4.3.
+For each $k\geq 2$, we define a rational function $G_{k}(q)$ and a sequence $(g_{k,\ell})_{\ell\geq 0}$ by
+$G_{k}(q)=\dfrac{q(1-(-q)^{k-1})}{(1-q)(1-(-q)^{k})}=\sum_{\ell=0}^{\infty}g_{k,\ell}q^{\ell}.$
+Theorem 4.1 implies that, for an arrangement $\mathcal{A}$ of rank $3$ with $\mathcal{A}=n$, we have
+(4.1) $\mathrm{Mag}(\mathcal{A},-q)=\dfrac{1}{1+(-q)^{n}}\biggl(\#\mathrm{Ch}(\mathcal{A})+\sum_{k\geq 2}2kn_{k}G_{k}(q)\biggr).$
+Lemma 4.4.
+Let $\mathcal{A}$ be an arrangement of rank $3$. Let $n_{k}$ be the number of faces $F\in\mathcal{F}(\mathcal{A})$ with $\operatorname{codim}F=2$ and $\mathcal{A}_{F}=k$. Then we have $n_{2}>0$.
+Proof.
+We may assume that $\mathcal{A}$ is essential.
+Consider the zonotope $Z(\mathcal{A})$ of $\mathcal{A}$, which is a polytope in $\mathbb{R}^{3}$.
+The facets of $Z(\mathcal{A})$ are polygons with an even number of vertices.
+The number of facets with $2k$ vertices is exactly $n_{k}$.
+If $n_{2}=0$, then every facet of $Z(\mathcal{A})$ must have at least $6$ vertices.
+However, Euler’s polyhedral formula shows that this is impossible.
+Therefore we have $n_{2}>0$.
+∎
+Theorem 4.5.
+Let $\mathcal{A}$ be an arrangement with $\operatorname{rank}\mathcal{A}\leq 3$. Let $c_{\ell}$ denote the coefficient of $q^{\ell}$ in the series expansion of $\mathrm{Mag}(\mathcal{A})$:
+$\mathrm{Mag}(\mathcal{A})=\sum_{\ell=0}^{\infty}c_{\ell}q^{\ell}.$
+Then we have $(-1)^{\ell}c_{\ell}>0$ for all sufficiently large $\ell$.
+Proof.
+When $\operatorname{rank}\mathcal{A}\leq 2$, the claim can be checked directly.
+Therefore, we may assume that $\operatorname{rank}\mathcal{A}=3$.
+Let $\mathcal{A}=n$.
+Write $n_{k}$ for the number of faces $F\in\mathcal{F}(\mathcal{A})$ with $\operatorname{codim}F=2$ and $\mathcal{A}_{F}=k$.
+Define $(e_{\ell})_{\ell\geq 0}$ by
+$\mathrm{Mag}(\mathcal{A},-q)=\dfrac{1}{1+(-q)^{n}}\biggl(\#\mathrm{Ch}(\mathcal{A})+\sum_{k\geq 2}2kn_{k}G_{k}(q)\biggr)=\sum_{\ell=0}^{\infty}e_{\ell}q^{\ell}.$
+It suffices to show that $e_{\ell}>0$ for all sufficiently large $\ell$. Put
+$H(q):=\#\mathrm{Ch}(\mathcal{A})+\sum_{k=2}^{\infty}2kn_{k}G_{k}(q),\qquad H(q)=\sum_{\ell\geq 0}a_{\ell}q^{\ell}.$
+We first record two elementary facts about the coefficients $g_{k,\ell}$ of $G_{k}$. If $k$ is even, then
+$G_{k}(q)=\frac{q(1+q^{k-1})}{(1-q)(1-q^{k})}=\frac{q}{1-q}\sum_{j\geq 0}\left(q^{kj}+q^{kj+k-1}\right).$
+Hence $g_{k,\ell}$ is non-negative and non-decreasing in $\ell$. In particular, for $k=2$ we have $g_{2,\ell}=\ell$. If $k$ is odd, then
+$G_{k}(q)=\frac{q(1-q^{k-1})}{(1-q)(1+q^{k})}=\frac{q(1+q+\cdots+q^{k-2})}{1+q^{k}}.$
+Thus the coefficient sequence $(g_{k,\ell})_{\ell\geq 0}$ is periodic and bounded. Since $n_{2}>0$ by Lemma 4.4, there is a constant $C>0$ such that $a_{\ell}\geq 4n_{2}\ell-C$ for all $\ell\geq 0$.
+We now distinguish the parity of $n$. First suppose that $n$ is odd. Then $1+(-q)^{n}=1-q^{n}$, so
+$\mathrm{Mag}(\mathcal{A},-q)=\frac{H(q)}{1-q^{n}}=H(q)\sum_{j\geq 0}q^{nj}.$
+Therefore, with $N=\lfloor\ell/n\rfloor$,
+$e_{\ell}=\sum_{j=0}^{N}a_{\ell-nj}\geq\sum_{j=0}^{N}\bigl(4n_{2}(\ell-nj)-C\bigr).$
+The right-hand side is a polynomial of degree $1$ in $\ell$ with leading coefficient $4n_{2}(N+1)>0$, so it tends to $\infty$ as $\ell\to\infty$.
+Hence $e_{\ell}>0$ for all sufficiently large $\ell$ in this case.
+Next suppose that $n$ is even. Then $1+(-q)^{n}=1+q^{n}$ and
+$\frac{1}{1+q^{n}}=\sum_{j\geq 0}(-1)^{j}q^{nj}.$
+For even $k$, let $b_{k,\ell}$ be the coefficient of $q^{\ell}$ in the series expansion of $\frac{G_{k}(q)}{1+q^{n}}$. Then we have
+$b_{k,\ell}=\sum_{j=0}^{\lfloor\ell/n\rfloor}(-1)^{j}g_{k,\ell-nj}.$
+Since $(g_{k,\ell})_{\ell}$ is non-negative and non-decreasing, pairing adjacent terms gives $b_{k,\ell}\geq 0$ for every $\ell$.
+For $k=2$, using $g_{2,m}=m$, we also get $b_{2,\ell}\to+\infty$.
+We claim that the odd $k$ terms, after division by $1+q^{n}$, are bounded.
+Indeed, for odd $k$ and even $n$,
+$\frac{G_{k}(q)}{1+q^{n}}=\frac{q(1+q+\cdots+q^{k-2})}{(1+q^{k})(1+q^{n})}.$
+Moreover $\gcd(1+q^{k},1+q^{n})=1$, because $k$ is odd and $n$ is even.
+Hence the denominator has only simple roots of unity, and the coefficient sequence is bounded.
+The coefficient sequence of $\mathrm{Ch}(\mathcal{A})/(1+q^{n})$ is also bounded.
+Thus there is a constant $D>0$ such that the total contribution of $\mathrm{Ch}(\mathcal{A})$ and all odd $k$ terms to $e_{\ell}$ is at least D-D for every $\ell$.
+Combining these observations, and using Lemma 4.4, we obtain
+$e_{\ell}\geq 4n_{2}b_{2,\ell}-D.$
+Since $b_{2,\ell}\to+\infty$, the right-hand side is positive for all sufficiently large $\ell$. This proves the theorem. ∎
+4.3. Arrangements with non-alternating magnitude
+In this section, we use the construction of Burr–Grünbaum–Sloane [5] for the orchard problem to construct an example of an arrangement whose magnitude does not have the alternating-sign property.
+Recall the formula (4.1) for $\mathrm{Mag}(\mathcal{A},-q)$. The series expansion of $G_{3}(q)$ is given by
+$G_{3}(q)=\dfrac{q}{1-q+q^{2}}=q+q^{2}-q^{4}-q^{5}+q^{7}+q^{8}-\cdots.$
+This suggests that for an arrangement with large $n_{3}$, the series expansion of $\mathrm{Mag}(\mathcal{A},-q)$ might have negative coefficients. This motivates the construction given below.
+Definition 4.6.
+Let $S\subset\mathbb{R}^{2}$ be a finite set of points in $\mathbb{R}^{2}$.
+For each $k\geq 2$, define $m_{k}$ to be the number of lines that contain exactly $k$ points from $S$.
+Define an arrangement $\mathcal{A}(S)$ in $\mathbb{R}^{3}$ by
+$\mathcal{A}(S)=\{(x,y,1)^{\perp}\mid(x,y)\in S\}.$
+The arrangement $\mathcal{A}(S)$ is essential unless $S$ is contained in a single line. It satisfies $n_{k}=2m_{k}$ for $k\geq 2$.
+We use the following result to construct an $S$ with large $m_{3}$.
+Theorem 4.7 ([5, Theorem 1], see also [14, Proposition 2.6]).
+For any $N\geq 3$, there exists a finite set $S\subset\mathbb{R}^{2}$ with #S=N\#S=N such that
+$m_{k}=\begin{cases}N-1-2\cdot\mathbf{1}_{3\mid N},&\text{ if }k=2;\\ \biggl\lfloor\dfrac{N(N-3)}{6}+1\biggr\rfloor,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4,\end{cases}$
+where $\mathbf{1}_{3\mid N}$ is equal to $1$ when $3$ divides $N$ and 0 otherwise.
+Example 4.8.
+As an example of Theorem 4.7, we construct a set of points $S\subset\mathbb{R}^{2}$ with #S=18\#S=18 such that
+$m_{k}=\begin{cases}15,&\text{ if }k=2;\\ 46,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4.\end{cases}$
+Let $X$ be the set of non-singular points of the real projective algebraic curve
+$Y^{2}=X^{2}(X-Z).$
+Then the points of $X$ naturally form a group isomorphic to $\mathbb{R}/\mathbb{Z}$. An explicit parametrization is given by
+$\phi(t)=\bigl[\cot^{2}(\pi t)+1:\cot(\pi t)(\cot^{2}(\pi t)+1):1\bigr]\quad(t\in\mathbb{R}/\mathbb{Z}).$
+Let $P_{j}\in\mathbb{R}^{2}$ be the image of $\phi(j/18)$ under the projective transformation $X:Y:Z]\mapsto[X:Y:X+Y+Z]$.
+Then $P_{i},P_{j},P_{k}$ lie on a single line if and only if $i+j+k\equiv 0\pmod{18}$.
+Define $S\subset\mathbb{R}^{2}$ by
+$S=\{P_{j}\mid j=0,1,\dots,17\}.$
+There are exactly $46$ triples $\{i,j,k\}\subset\mathbb{Z}/18\mathbb{Z}$ such that $i+j+k\equiv 0\pmod{18}$, so we have $m_{3}=46$.
+Also, $m_{k}=0$ for $k\geq 4$.
+Since $m_{2}+3m_{3}=\binom{18}{2}$ by double-counting, we have $m_{2}=15$.
+Example 4.9.
+We give an explicit construction of an arrangement $\mathcal{A}$ such that the series expansion of $\mathrm{Mag}(\mathcal{A})$ does not have alternating signs.
+Let $S\subset\mathbb{R}^{2}$ be a finite set as in Example 4.8.
+Then the associated arrangement $\mathcal{A}(S)$ satisfies
+$\mathcal{A}(S)=18,\quad n_{k}=\begin{cases}30,&\text{ if }k=2;\\ 92,&\text{ if }k=3;\\ 0,&\text{ if }k\geq 4,\end{cases}\quad\text{and}\quad\#\mathrm{Ch}(\mathcal{A})=216.$
+Figure 2 shows the projectivization of $\mathcal{A}(S)$. By Theorem 4.1, the magnitude of $\mathcal{A}(S)$ is given by
+$\mathrm{Mag}(\mathcal{A})$ ${}=\dfrac{1}{1+q^{18}}\biggl(216-\dfrac{120q}{(1+q)^{2}}-\dfrac{552q}{1+q+q^{2}}\biggr)$
+${}=216-672q+792q^{2}-360q^{3}-72q^{4}-48q^{5}+720q^{6}-1392q^{7}+1512q^{8}-\cdots.$
+In particular, the coefficient of $q^{4}$ is 72<0-72<0.
+Figure 2. An arrangement whose magnitude does not have alternating signs
+5. Magnitude homology of arrangements
+We continue to denote invariants of $\mathcal{T}(\mathcal{A})$ simply by the corresponding invariants of $\mathcal{A}$, such as $MH_{k,\ell}(\mathcal{A}):=MH_{k,\ell}(\mathcal{T}(\mathcal{A}))$, $\beta_{k,\ell}(\mathcal{A}):=\operatorname{rank}MH_{k,\ell}(\mathcal{A})$, and $\chi_{\ell}(\mathcal{A})=\sum_{k}(-1)^{k}\beta_{k,\ell}(\mathcal{A})$.
 Also recall the notation $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$ for $X\in L(\mathcal{A})$.
-4.1. Small length
-Let us start our study of $MH_{k,\ell}(\mathcal{A})$ from investigating small length $\ell\leq 2$.
+5.1. Small lengths
+We begin our study of $MH_{k,\ell}(\mathcal{A})$ by investigating small lengths $\ell\leq 2$.
 Proposition 5.1.
-Let $\mathcal{A}$ be a real central hyperplane arrangement. Then the only nontrivial $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are as follows.
+Let $\mathcal{A}$ be an arrangement. Then the only nontrivial $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are as follows.
   1. (1)
@@
 Proof.
-(1) and (2) follow from (1) and (2) of Proposition 2.8 and noting that an edge in $\mathcal{T}(\mathcal{A})$ corresponds to a chamber of restriction $\mathcal{A}^{H}$ for some $H\in\mathcal{A}$.
-For (3), since $m_{\mathcal{T}(\mathcal{A})}\geq 3$, we may apply Theorem 2.20 for $\ell=2<m_{\mathcal{T}(\mathcal{A})}$.
-$MH_{2,2}(\mathcal{A})\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})\cup P_{2,2}(\mathcal{A})}MH_{2,\vec{a}}(\mathcal{A}).$
+(1) and (2) follow from (1) and (2) of Proposition 2.8, together with the observation that an edge in $\mathcal{T}(\mathcal{A})$ corresponds to a chamber of the restriction $\mathcal{A}^{H}$ for some $H\in\mathcal{A}$.
+For (3), since $m_{\mathcal{T}(\mathcal{A})}\geq 3$, we may apply Theorem 2.22 for $\ell=2<m_{\mathcal{T}(\mathcal{A})}$:
+$MH_{2,2}(\mathcal{A})\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})\cup P_{2,2}(\mathcal{A})}MH^{\mathrm{gs}}_{2,\vec{a}}(\mathcal{A}).$
 By Theorem 2.25, the framed magnitude homology of $\mathcal{A}$ can be computed from interval posets.
@@
 If $\mathcal{A}_{X}>2$, then $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ is a singleton.
-Since $MH_{2,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{0}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})$, taking reduced homology, only the first type $(a_{0},a_{1})$ with $\mathcal{A}_{X}=2$ contributes to the magnitude homology a copy of $\mathbb{Z}$.
-Then the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{1,2}(\mathcal{A})$ is $4\sum_{X\in L_{2}(\mathcal{A}),\#\mathcal{A}_{X}=2}c^{X}$ copies of $\mathbb{Z}$.
+Since $MH^{\mathrm{gs}}_{2,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{0}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})$, taking reduced homology shows that only the first type of $(a_{0},a_{1})$, with $\mathcal{A}_{X}=2$, contributes a copy of $\mathbb{Z}$ to the magnitude homology.
+Hence the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{1,2}(\mathcal{A})$ is $4\sum_{X\in L_{2}(\mathcal{A}),\#\mathcal{A}_{X}=2}c^{X}$ copies of $\mathbb{Z}$.
 Similarly, for $\vec{a}=(a_{0},a_{1},a_{2})\in P_{2,2}(\mathcal{A})$, if $a_{0}\neq a_{2}$, then $\vec{a}$ is geodesic and hence $\varphi(\vec{a})\neq\vec{a}$, making the $\vec{a}$-framed magnitude homology trivial.
-So we only need to consider $\vec{a}=(a_{0},a_{1},a_{0})\in P_{2,2}(\mathcal{A})$, where $a_{0}$ and $a_{1}$ are adjacent.
+Thus we only need to consider $\vec{a}=(a_{0},a_{1},a_{0})\in P_{2,2}(\mathcal{A})$, where $a_{0}$ and $a_{1}$ are adjacent.
 The interval posets $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ and $(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})}$ are empty.
-Then $MH_{2,(a_{0},a_{1},a_{0})}(\mathcal{A})\cong\widetilde{H}_{-2}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}\otimes C(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})})\cong\mathbb{Z}$.
+Then $MH^{\mathrm{gs}}_{2,(a_{0},a_{1},a_{0})}(\mathcal{A})\cong\widetilde{H}_{-2}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}\otimes C(a_{1},a_{0})_{\mathcal{T}(\mathcal{A})})\cong\mathbb{Z}$.
 Each such $(a_{0},a_{1},a_{0})$ corresponds to a directed edge of $\mathcal{T}(\mathcal{A})$.
-Then the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{2,2}(\mathcal{A})$ is $\beta_{1,1}(\mathcal{A})$ copies of $\mathbb{Z}$.
-Now (3) is proved.
-To finish the proof, we need to show other $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are trivial. By Proposition 2.8 (3), we only need to prove $MH_{1,2}(\mathcal{A})=0$. We use Theorem 2.20 again
-$MH_{1,2}(A)\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})}MH_{1,\vec{a}}(\mathcal{A}).$
-Consider $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$, the interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ cannot be empty.
-Hence $MH_{1,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{-1}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})=0$.
+Hence the total contribution to $MH_{2,2}(\mathcal{A})$ from $P_{2,2}(\mathcal{A})$ is $\beta_{1,1}(\mathcal{A})$ copies of $\mathbb{Z}$.
+This proves (3).
+To finish the proof, we need to show that all other $MH_{k,\ell}(\mathcal{A})$ for $\ell\leq 2$ are trivial.
+By Proposition 2.8 (3), we only need to prove $MH_{1,2}(\mathcal{A})=0$.
+We use Theorem 2.22 again:
+$MH_{1,2}(\mathcal{A})\cong\bigoplus_{\vec{a}\in P_{1,2}(\mathcal{A})}MH^{\mathrm{gs}}_{1,\vec{a}}(\mathcal{A}).$
+Consider $\vec{a}=(a_{0},a_{1})\in P_{1,2}(\mathcal{A})$.
+The interval poset $(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})}$ cannot be empty.
+Hence $MH^{\mathrm{gs}}_{1,(a_{0},a_{1})}(\mathcal{A})\cong\widetilde{H}_{-1}(C(a_{0},a_{1})_{\mathcal{T}(\mathcal{A})})=0$.
 ∎
 5.2. Geodesic magnitude homology and diagonal tope graphs
-Let us determine the geodesic magnitude homology of arrangements.
+We now determine the geodesic magnitude homology of arrangements.
+Recall that $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$.
+Let $c_{X}=\#\mathrm{Ch}(\mathcal{A}_{X})$.
+Both $c^{X}$ and $c_{X}$ are determined by the intersection lattice $L(\mathcal{A})$.
 Theorem 5.2.
-Let $\mathcal{A}$ be an arrangement. Then its geodesic magnitude homology is as follow
-$MH_{k,\ell}^{geod}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A}):\mathrm{rank}\mathcal{A}_{F}=k,\#\mathcal{A}_{F}=\ell}\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A}_{F})}$
+Let $\mathcal{A}$ be an arrangement. Then its geodesic magnitude homology is given as follows:
+$MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ \operatorname{rank}\mathcal{A}_{F}=k,\ \#\mathcal{A}_{F}=\ell\end{subarray}}\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A}_{F})}\cong\bigoplus_{\begin{subarray}{c}X\in L_{k}(\mathcal{A})\\ \#\mathcal{A}_{X}=\ell\end{subarray}}\mathbb{Z}^{c^{X}c_{X}}.$
+In particular, $MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})$ is determined by $L(\mathcal{A})$.
 Proof.
-By Corollary 2.22,
-$MH_{k,\ell}^{geod}(\mathcal{A})\cong\bigoplus_{a,b\in\mathrm{Ch}(\mathcal{A}):d(a,b)=\ell}\widetilde{H}_{k-2}(C_{*}(a,b)_{\mathcal{T}(\mathcal{A})}).$
-By Proposition 2.34, the interval $(a,b)_{\mathcal{T}(\mathcal{A})}$ is homotopy equivalent to $S^{\mathrm{rank}\mathcal{A}_{F}-2}$ if $a,b]_{\mathcal{T}(\mathcal{A})}=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$ and contractible otherwise.
-Therefore each $F\in\mathcal{F}(\mathcal{A})$ such that $k=\mathrm{rank}\mathcal{A}_{F}$ and $\ell=\#\mathcal{A}_{F}$ contribute to $MH_{k,\ell}^{geod}(\mathcal{A})$ a copy of $\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A})_{F}}$ and there is no other summand.
+By Corollary 2.24, we have
+$MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})\cong\bigoplus_{\begin{subarray}{c}a,b\in\mathrm{Ch}(\mathcal{A})\\ d(a,b)=\ell\end{subarray}}\widetilde{H}_{k-2}(C_{*}(a,b)_{\mathcal{T}(\mathcal{A})}).$
+By Proposition 2.35, the interval $(a,b)_{\mathcal{T}(\mathcal{A})}$ is homotopy equivalent to $S^{\operatorname{rank}\mathcal{A}_{F}-2}$ if $a,b]_{\mathcal{T}(\mathcal{A})}=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$, and is contractible otherwise.
+Therefore each $F\in\mathcal{F}(\mathcal{A})$ such that $k=\operatorname{rank}\mathcal{A}_{F}$ and $\ell=\#\mathcal{A}_{F}$ contributes a copy of $\mathbb{Z}^{\#\mathrm{Ch}(\mathcal{A})_{F}}$ to $MH_{k,\ell}^{\mathrm{geod}}(\mathcal{A})$, and there is no other summand.
 ∎
-Recall that $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$ and let $c_{X}=\#\mathrm{Ch}(\mathcal{A}_{X})$. We can give a combinatorial description of geodesic magnitude Betti number.
 Corollary 5.3.
-The geodesic magnitude Betti number
-$\beta_{k,\ell}^{geod}(\mathcal{A})=\sum_{X\in L_{k}(\mathcal{A}):\#\mathcal{A}_{X}=\ell}c^{X}c_{X}.$
-In particular, geodesic magnitude homology is determined by $L(\mathcal{A})$.
+Let $\mathcal{A}$ be an arrangement. Then its geodesic magnitude Betti numbers are given by
+$\beta_{k,\ell}^{\mathrm{geod}}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L_{k}(\mathcal{A})\\ \#\mathcal{A}_{X}=\ell\end{subarray}}c^{X}c_{X}.$
 As an application, we give a characterization of diagonal tope graphs.
@@
 We know that the Boolean arrangement $Bl_{d}$ has tope graph $\mathcal{T}(Bl_{d})=Q_{d}$, which is diagonal (see Example 2.14).
-Suppose $\mathcal{A}$ is not Boolean, then $n=\#\mathcal{A}>r=\mathrm{rank}\mathcal{A}$.
-Therefore
-$\beta_{r,n}(\mathcal{A})\geq\beta_{r,n}^{geod}(\mathcal{A})=\#\mathrm{Ch}(\mathcal{A}),$
-where the equation is from Corollary 4.3. This means $\mathcal{A}$ is not diagonal. ∎
+Suppose $\mathcal{A}$ is not Boolean.
+Then $n=\#\mathcal{A}>r=\operatorname{rank}\mathcal{A}$.
+We have
+$\beta_{r,n}(\mathcal{A})\geq\beta_{r,n}^{\mathrm{geod}}(\mathcal{A})=\#\mathrm{Ch}(\mathcal{A}),$
+where the equality follows from Corollary 5.3. This means that $\mathcal{A}$ is not diagonal. ∎
 5.3. Magnitude homology with support
-Consider a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A})$, let us define $S(\vec{x})=S(x_{0},x_{1})\cup\cdots\cup S(x_{k-1},x_{k})$ as the set of hyperplanes crossed by $\vec{x}$.
-Notice that $S(\vec{x})\leq\ell(\vec{x})$.
-It is not difficult to verify that $x_{i}$ is smooth in $\vec{x}$, i.e.
-$d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$, is equivalent to $S(x_{i-1},x_{i+1})=S(x_{i-1},x_{i})\sqcup S(x_{i},x_{i+1})$.
-So if a proper chain $\vec{y}$ appears in the boundary of $\vec{x}$, then $S(\vec{x})=S(\vec{y})$.
+Consider a proper chain $\vec{x}=(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A})$, and define $S(\vec{x})=S(x_{0},x_{1})\cup\cdots\cup S(x_{k-1},x_{k})$ as the set of hyperplanes crossed by $\vec{x}$.
+Observe that $S(\vec{x})\leq\ell(\vec{x})$.
+One verifies that $x_{i}$ is smooth in $\vec{x}$, i.e.
+$d(x_{i-1},x_{i+1})=d(x_{i-1},x_{i})+d(x_{i},x_{i+1})$, if and only if $S(x_{i-1},x_{i+1})=S(x_{i-1},x_{i})\sqcup S(x_{i},x_{i+1})$.
+Thus, if a proper chain $\vec{y}$ appears in the boundary of $\vec{x}$, then $S(\vec{x})=S(\vec{y})$.
 In other words, the boundary operator $\partial:MC_{k,\ell}(\mathcal{A})\to MC_{k-1,\ell}(\mathcal{A})$ preserves crossed hyperplanes.
-This motivated us to define the following.
+This motivates the following definition.
 Definition 5.5.
-For arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$, we define the $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A})$ generated by $P_{k,\ell}(\mathcal{A};\mathcal{B})=\{\vec{x}\in P_{k,\ell}(\mathcal{A})\mid S(\vec{x})=\mathcal{B}\}$.
+For an arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$, we define the $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A})$ generated by $P_{k,\ell}(\mathcal{A};\mathcal{B})=\{\vec{x}\in P_{k,\ell}(\mathcal{A})\mid S(\vec{x})=\mathcal{B}\}$.
 The above discussion shows that $MC_{*,\ell}(\mathcal{A};\mathcal{B})$ forms a subcomplex of $MC_{*,\ell}(\mathcal{A})$, called the $\mathcal{B}$-supported magnitude complex of $\mathcal{A}$.
 Its homology $MH_{*,\ell}(\mathcal{A};\mathcal{B})$ is called the $\mathcal{B}$-supported magnitude homology of $\mathcal{A}$.
-Definition 4.6.
-When $\mathcal{B}=\mathcal{A}$, we call $MC_{*,\ell}^{\circ}(\mathcal{A}):=MC_{*,\ell}(\mathcal{A};\mathcal{A})$ the length $\ell$ interior magnitude complex of $\mathcal{A}$ and $MH_{*,\ell}^{\circ}(\mathcal{A}):=MH_{*,\ell}(\mathcal{A};\mathcal{A})$ the length $\ell$ interior magnitude homology of $\mathcal{A}$.
-Let us also define the interior magnitude Betti number $\beta_{k,\ell}^{\circ}(\mathcal{A}):=\mathrm{rank}MH_{k,\ell}^{\circ}(\mathcal{A})$ and the length $\ell$ interior magnitude Euler characteristic $\chi_{\ell}^{\circ}(\mathcal{A}):=\sum_{k}(-1)^{k}\beta_{k,\ell}^{\circ}(\mathcal{A})$.
 The following result is a direct consequence of the above observation that the magnitude boundary operator preserves crossed hyperplanes.
 Proposition 5.6.
-For arrangement $\mathcal{A}$, we have the decomposition of magnitude chain groups
+For an arrangement $\mathcal{A}$, we have the following decomposition of magnitude chain groups:
 $MC_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MC_{k,\ell}(\mathcal{A};\mathcal{B}),$
-inducing a decomposition of homology groups
+which induces a decomposition of homology groups
 $MH_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MH_{k,\ell}(\mathcal{A};\mathcal{B}).$
-Furthermore $MC_{k,\ell}(\mathcal{A};\mathcal{B})=0$ and $MH_{k,\ell}(\mathcal{A};\mathcal{B})=0$ if $\ell<\#\mathcal{B}$.
-4.4. Face decomposition of magnitude homology
-We would like to further decompose the magnitdue complex/homology with support. To this end, we introduce another notation.
+Furthermore, $MC_{k,\ell}(\mathcal{A};\mathcal{B})=MH_{k,\ell}(\mathcal{A};\mathcal{B})=0$ if $\ell<\#\mathcal{B}$.
+Using the notion of magnitude homology with support, we define interior magnitude homology. Later, we show that this is a categorification of the interior magnitude defined in Definition 3.6.
 Definition 5.7.
-For arrangement $\mathcal{A}$ and a subarrangement $\mathcal{B}\subseteq\mathcal{A}$.
+Let $\mathcal{A}$ be an arrangement and let $X\in L(\mathcal{A})$ be a flat. We define the $X$-supported magnitude complex of $\mathcal{A}$ by
+$MC_{*,\ell}(\mathcal{A};X):=\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}\\ \cap\mathcal{B}=X\end{subarray}}MC_{*,\ell}(\mathcal{A};\mathcal{B}).$
+Its homology $MH_{*,\ell}(\mathcal{A};X)$ is called the $X$-supported magnitude homology of $\mathcal{A}$. In particular, we define
+$MC_{*,\ell}^{\circ}(\mathcal{A}):=MC_{*,\ell}(\mathcal{A};\cap\mathcal{A}),$
+and call it the length $\ell$ interior magnitude complex of $\mathcal{A}$.
+We call $MH_{*,\ell}^{\circ}(\mathcal{A}):=H_{*}(MC_{*,\ell}^{\circ}(\mathcal{A}))$ the length $\ell$ interior magnitude homology of $\mathcal{A}$.
+We also define the interior magnitude Betti number $\beta_{k,\ell}^{\circ}(\mathcal{A}):=\operatorname{rank}MH_{k,\ell}^{\circ}(\mathcal{A})$ and the length $\ell$ interior magnitude Euler characteristic $\chi_{\ell}^{\circ}(\mathcal{A}):=\sum_{k}(-1)^{k}\beta_{k,\ell}^{\circ}(\mathcal{A})$.
+5.4. Face decomposition for magnitude homology
+Here we prove a homological version of the face decomposition formula established in Theorem 3.7. To this end, we introduce another notation.
+Definition 5.8.
+Let $\mathcal{A}$ be an arrangement and let $\mathcal{B}\subseteq\mathcal{A}$ be a subarrangement.
 Consider an induced subgraph $G$ of $\mathcal{T}(\mathcal{A})$.
-We define the localized $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(G;\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ generated by chains entirely contained in $G$,
+We define the localized $\mathcal{B}$-supported magnitude chain group $MC_{k,\ell}(G;\mathcal{B})$ as the subgroup of $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ generated by chains entirely contained in $G$:
 $P_{k,\ell}(G;\mathcal{B})=\{(x_{0},\ldots,x_{k})\in P_{k,\ell}(\mathcal{A};\mathcal{B})\mid x_{0},\ldots,x_{k}\in V(G)\}.$
@@
 Furthermore, by Lemma 2.29, the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$ is gated, and therefore convex (Proposition 2.2).
-Because the boundary operator $\partial_{i}$ merely deletes the chamber $x_{i}$ from the sequence, all remaining chambers in the chain $\partial_{i}\vec{x}$ are still contained within the vertex set of $\mathcal{T}(\mathcal{A})_{F}$.
+Because the boundary operator $\partial_{i}$ only deletes the chamber $x_{i}$ from the sequence, all remaining chambers in the chain $\partial_{i}\vec{x}$ are still contained in the vertex set of $\mathcal{T}(\mathcal{A})_{F}$.
 Thus, $\partial(MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B}))\subseteq MC_{k-1,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})$.
 ∎
-Now we are ready to state the main structural result of the magnitude homology of an arrangement.
+Proposition 5.10.
+Let $\mathcal{A}$ be an arrangement and let $X\in L(\mathcal{A})$ be a flat. Then the $X$-supported magnitude complex can be decomposed as follows:
+$MC_{*,\ell}(\mathcal{A};X)\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}MC_{*,\ell}^{\circ}(\mathcal{A}_{F}).$
+Proof.
+By Definition 5.7, $MC_{*,\ell}(\mathcal{A};X)=\bigoplus_{\mathcal{B}\subseteq\mathcal{A},\cap\mathcal{B}=X}MC_{*,\ell}(\mathcal{A};\mathcal{B})$.
+For a subarrangement $\mathcal{B}$ with $\cap\mathcal{B}=X$, consider a chain $\vec{x}\in P_{k,\ell}(\mathcal{A};\mathcal{B})$.
+By Definition 5.5, $S(\vec{x})=\mathcal{B}$.
+The chain $\vec{x}$ never crosses any hyperplane in $\mathcal{A}\setminus\mathcal{A}_{X}$.
+Therefore, every chamber in $\vec{x}$ has a constant nonzero sign on each hyperplane $H\in\mathcal{A}\setminus\mathcal{A}_{X}$.
+These signs determine a chamber $F\in\mathrm{Ch}(\mathcal{A}^{X})$, or equivalently a face $F\in\mathcal{F}(\mathcal{A})$ with $s(F)=X$.
+Therefore the chain $\vec{x}$ lies in the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$.
+By Lemma 2.29, $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph of $\mathcal{T}(\mathcal{A})$ isometrically isomorphic to $\mathcal{T}(\mathcal{A}_{F})$.
+This induces a direct-sum decomposition of the $\mathcal{B}$-supported chains over the faces supported on $X$:
+$MC_{k,\ell}(\mathcal{A};\mathcal{B})\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}MC_{k,\ell}(\mathcal{A}_{F};\mathcal{B}).$
+Summing both sides over $\mathcal{B}\subseteq\mathcal{A}$ such that $\cap\mathcal{B}=X$, we obtain
+$MC_{*,\ell}(\mathcal{A};X)\cong\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}_{F}\\ \cap\mathcal{B}=X\end{subarray}}MC_{*,\ell}(\mathcal{A}_{F};\mathcal{B}).$
+Since we are summing over $F$ supported on $X$, we have $\mathcal{A}_{F}=\mathcal{A}_{X}$ and $\cap\mathcal{A}_{F}=X$, so the inner sum is exactly $MC_{*,\ell}^{\circ}(\mathcal{A}_{F})$.
+This proves the desired formula.
+∎
+We are now ready to state the main structural result for the magnitude homology of an arrangement.
 Theorem 5.11.
-For an arrangement $\mathcal{A}$, we have the face decomposition of its magnitude complex
+For an arrangement $\mathcal{A}$, we have the following face decomposition of its magnitude complex:
 $MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC_{k,\ell}^{\circ}(\mathcal{A}_{F}),$
-inducing the face decomposition of magnitude homology
+which induces the following face decomposition of magnitude homology:
 $MH_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MH_{k,\ell}^{\circ}(\mathcal{A}_{F}).$
 Proof.
-By Proposition 4.7, the global magnitude chain complex decomposes into a direct sum over all subarrangements $\mathcal{B}\subseteq\mathcal{A}$,
+By Proposition 5.6, the magnitude chain complex decomposes into a direct sum over all subarrangements $\mathcal{B}\subseteq\mathcal{A}$,
 $MC_{k,\ell}(\mathcal{A})=\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}MC_{k,\ell}(\mathcal{A};\mathcal{B}).$
-For a fixed subarrangement $\mathcal{B}$, consider a chain $\vec{x}\in P_{k,\ell}(\mathcal{A};\mathcal{B})$.
-By Definition 4.5, $S(\vec{x})=\mathcal{B}$, meaning the chain never crosses any hyperplane in the complement $\mathcal{A}\setminus\mathcal{B}$.
-Therefore, every chamber in $\vec{x}$ shares a strictly constant, non-zero sign on all hyperplanes in $\mathcal{A}\setminus\mathcal{B}$.
-This constant sign pattern uniquely defines a face $F\in\mathcal{F}(\mathcal{A})$ such that $F_{H}=0$ for $H\in\mathcal{B}$, and $F_{H}\in\{+,-\}$ for $H\notin\mathcal{B}$.
-By definition, the subarrangement associated with this face is exactly $\mathcal{A}_{F}=\mathcal{B}$.
-Because the sign pattern on $\mathcal{A}\setminus\mathcal{B}$ is uniquely determined for each chain, the subgroup $MC_{k,\ell}(\mathcal{A};\mathcal{B})$ further decomposes into a direct sum over the faces $F$ that satisfy $\mathcal{A}_{F}=\mathcal{B}$,
-$MC_{k,\ell}(\mathcal{A};\mathcal{B})=\bigoplus_{F\in\mathcal{F}(\mathcal{A}):\mathcal{A}_{F}=\mathcal{B}}MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B}).$
-By Lemma 2.26, the $F$-subgraph $\mathcal{T}(\mathcal{A})_{F}$ is a gated subgraph isometrically isomorphic to the tope graph of the localized subarrangement $\mathcal{A}_{F}$.
-Under this isomorphism, the $\mathcal{B}$-supported complex maps identically to the full interior complex of $\mathcal{A}_{F}$.
-Therefore, by Definition 4.6,
-$MC_{k,\ell}(\mathcal{T}(\mathcal{A})_{F};\mathcal{B})\cong MC_{k,\ell}(\mathcal{A}_{F};\mathcal{A}_{F})=MC^{\circ}_{k,\ell}(\mathcal{A}_{F}).$
-Substituting this isomorphism back into the original sum over subarrangements $\mathcal{B}$ yields the result
-$MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{\mathcal{B}\subseteq\mathcal{A}}\left(\bigoplus_{F:\mathcal{A}_{F}=\mathcal{B}}MC^{\circ}_{k,\ell}(\mathcal{A}_{F})\right)=\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC^{\circ}_{k,\ell}(\mathcal{A}_{F}).$
-Since this is a decomposition into subcomplexes, we have the face decomposition of magnitude homology
-$MH_{k,\ell}(\mathcal{A})\cong\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MH_{k,\ell}^{\circ}(\mathcal{A}_{F}).$
-For an arrangement $\mathcal{A}$, let us write
-$\chi^{\circ}(\mathcal{A},q):=\sum_{\ell=0}^{\infty}\chi_{\ell}^{\circ}(\mathcal{A})q^{\ell}\in\mathbb{Z}[\![q]\!].$
-Then we have a reformulation of Conjecture 3.7.
+Grouping the $\mathcal{B}$-summands according to $\cap\mathcal{B}$, we have
+$MC_{k,\ell}(\mathcal{A})=\bigoplus_{X\in L(\mathcal{A})}\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}\\ \cap\mathcal{B}=X\end{subarray}}MC_{k,\ell}(\mathcal{A};\mathcal{B})=\bigoplus_{X\in L(\mathcal{A})}MC_{k,\ell}(\mathcal{A};X).$
+By Proposition 5.10, the $X$-supported complex decomposes into interior complexes
+$MC_{k,\ell}(\mathcal{A})\cong\bigoplus_{X\in L(\mathcal{A})}\bigoplus_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\\ s(F)=X\end{subarray}}MC_{k,\ell}^{\circ}(\mathcal{A}_{F})=\bigoplus_{F\in\mathcal{F}(\mathcal{A})}MC_{k,\ell}^{\circ}(\mathcal{A}_{F}).$
+The decomposition of homology follows immediately. ∎
 Corollary 5.12.
-Let $r=\mathrm{rank}\mathcal{A}$ and $n=\#\mathcal{A}$. The following are equivalent
-  1. (1)
-    $\mathrm{Mag}(\mathcal{A},q)=\sum_{F}\mathrm{Mag}^{\circ}(\mathcal{A}_{F},q)$;
-  2. (2)
-    $\chi^{\circ}(\mathcal{A},q)=\mathrm{Mag}^{\circ}(\mathcal{A},q)$;
-  3. (3)
-    $\chi_{\ell+n}^{\circ}(\mathcal{A})=(-1)^{r}\chi_{\ell}(\mathcal{A})$ for $\ell\geq 0$.
+Let $\mathcal{A}$ be an arrangement with $\operatorname{rank}\mathcal{A}=r$ and $\mathcal{A}=n$. Then we have
+$\mathrm{Mag}^{\circ}(\mathcal{A})=\sum_{\ell=0}^{\infty}\chi_{\ell}^{\circ}(\mathcal{A})q^{\ell}.$
+Moreover, we have $\chi_{\ell+n}^{\circ}(\mathcal{A})=(-1)^{r}\chi_{\ell}(\mathcal{A})$ for $\ell\geq 0$, and $\chi_{\ell}^{\circ}(\mathcal{A})=0$ for $\ell<n$.
 Proof.
-Since Euler characteristic is additive over direct sums, the face decomposition of magnitude homology (Theorem 4.10) yields
-$\chi_{\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi_{\ell}^{\circ}(\mathcal{A}_{F}).$
-Multiplying by $q^{\ell}$ and summing over $\ell$, we obtain
-$\mathrm{Mag}(\mathcal{A},q)=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi^{\circ}(\mathcal{A}_{F},q),$
-By Möbius inversion on $\mathcal{F}(\mathcal{A})$, we see that (1) is equivalent to (2). The equivalence between (2) and (3) is straightforward by definition. ∎
+Define $\Delta(\mathcal{A})$ to be the difference of the two expressions:
+$\Delta(\mathcal{A})=\mathrm{Mag}^{\circ}(\mathcal{A})-\sum_{\ell=0}^{\infty}\chi_{\ell}^{\circ}(\mathcal{A})q^{\ell}.$
+We prove $\Delta(\mathcal{A})=0$ by induction on the rank of $\mathcal{A}$. By Theorem 3.7 and Theorem 5.11, we have
+$\mathrm{Mag}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F}),\quad\chi_{\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\chi^{\circ}_{\ell}(\mathcal{A}_{F}).$
+Since $\mathrm{Mag}(\mathcal{A})$ can be written as $\sum_{\ell}\chi_{\ell}(\mathcal{A})q^{\ell}$, we obtain
+$\sum_{F\in\mathcal{F}(\mathcal{A})}\Delta(\mathcal{A}_{F})=0.$
+The induction hypothesis shows that $\Delta(\mathcal{A}_{F})=0$ for every face $F$ other than $\bigcap\mathcal{A}$.
+Thus $\Delta(\mathcal{A})=0$.
+The final assertion follows from the identity $\mathrm{Mag}^{\circ}(\mathcal{A})=(-1)^{r}q^{n}\mathrm{Mag}(\mathcal{A})$.
+∎
 Remark 5.13.
-Note that if there is a (quasi-)isomorphism of chain complexes
-$MC_{*,\ell}(\mathcal{A})\to MC_{*+r,\ell+n}^{\circ}(\mathcal{A})=MC_{*+r,\ell+n}(\mathcal{A};\mathcal{A})$
-for all $\ell\geq 0$, then the conditions in Corollary 4.11 hold.
+The identity $\chi_{\ell+n}^{\circ}(\mathcal{A})=(-1)^{r}\chi_{\ell}(\mathcal{A})$ in Corollary 5.12 is analogous to the Ehrhart–Macdonald reciprocity theorem for lattice polytopes [11, 26] and Stanley’s reciprocity theorem for posets [31].
+Remark 5.14.
+In light of Corollary 5.12, it is natural to expect that there exists an isomorphism of homology groups
+$MH_{k,\ell}(\mathcal{A})\cong MH_{k+r,\ell+n}^{\circ}(\mathcal{A})$
+for all $k,\ell\geq 0$, and that the interior magnitude homology $MH_{k,\ell}^{\circ}(\mathcal{A})$ vanishes for $0\leq\ell<n$.
 5.5. Magnitude homology on the diagonal
-In this subsection, we give a combinatorial formula of the magnitude homology of arrangements on the diagonal $MH_{\ell,\ell}(\mathcal{A})$.
-Using Theorem 4.10, we can determine the interior magnitude homology of Boolean arrangements.
+In this subsection, we give a combinatorial formula for the diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ of an arrangement $\mathcal{A}$.
+First, we determine the interior magnitude homology of Boolean arrangements using Theorem 5.11.
 Corollary 5.15.
-The interior magnitude homology of Boolean arrangement $MH_{k,\ell}^{\circ}(Bl_{d})$ concentrates on the diagonal $k=\ell$ and $MH_{\ell,\ell}^{\circ}(Bl_{d})$ is free abelian of rank
+For the Boolean arrangement $Bl_{d}$, the interior magnitude homology $MH_{k,\ell}^{\circ}(Bl_{d})$ is concentrated on the diagonal $k=\ell$, and $MH_{\ell,\ell}^{\circ}(Bl_{d})$ is free abelian of rank
 $\beta_{\ell,\ell}^{\circ}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$.
@@
 We know from Corollary 5.4 and Example 2.14 that $\mathcal{T}(Bl_{d})$ is diagonal and $MH_{\ell,\ell}(Bl_{d})$ is free abelian of rank $\beta_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell+d-1}{d-1}$.
-By Theorem 4.10, the interior magnitude homology $MH_{k,\ell}^{\circ}(Bl_{d})$ is a direct summand of $MH_{k,\ell}(Bl_{d})$, and hence concentrates on the diagonal $k=\ell$.
+By Theorem 5.11, the interior magnitude homology $MH_{k,\ell}^{\circ}(Bl_{d})$ is a direct summand of $MH_{k,\ell}(Bl_{d})$, and hence is concentrated on the diagonal $k=\ell$.
 Theorem 5.11 also gives
 $\beta_{\ell,\ell}(Bl_{d})=\sum_{F\in\mathcal{F}(Bl_{d})}\beta_{\ell,\ell}^{\circ}((Bl_{d})_{F}).$
-Notice that if $F\in\mathcal{F}(Bl_{d})$ has $\mathrm{codim}F=j$, the localization $(Bl_{d})_{F}$ has essentialization $Bl_{j}$.
-Then applying Möbius inversion on $\mathcal{F}(\mathcal{A})$ yields the desired formula.
+If $F\in\mathcal{F}(Bl_{d})$ has $\operatorname{codim}F=j$, then the localization $(Bl_{d})_{F}$ has essentialization $Bl_{j}$.
+Applying Möbius inversion on $\mathcal{F}(Bl_{d})$ yields the desired formula.
 ∎
-Let us prove the following vanishing result of diagonal interior magnitude homology for non-Boolean arrangements.
+Next we prove the following vanishing result for diagonal interior magnitude homology of non-Boolean arrangements.
 Theorem 5.16.
-Let $\mathcal{A}$ be an arrangement of $\mathrm{rank}\mathcal{A}=r$ and $\mathcal{A}=n$.
-Suppose $\mathcal{A}$ is not Boolean $r<n$, then the diagonal interior magnitude homology of $\mathcal{A}$ vanishes for all length
-$MH_{\ell,\ell}^{\circ}(\mathcal{A})=0,\forall\ell\geq 0.$
+Let $\mathcal{A}$ be an arrangement of $\operatorname{rank}\mathcal{A}=r$ and $\mathcal{A}=n$.
+Suppose $\mathcal{A}$ is not Boolean, so that $r<n$.
+Then the diagonal interior magnitude homology of $\mathcal{A}$ vanishes in all lengths:
+$MH_{\ell,\ell}^{\circ}(\mathcal{A})=0,\quad\forall\ell\geq 0.$
 Proof.
-Choose an isometric embedding $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{n}=\{+,-\}^{n}$ of $\mathcal{T}(\mathcal{A})$ into the hypercube $Q_{n}$, which is not surjective since $r<n$.
-Since $\phi$ is distance preserving, it induces injective chain maps $\phi_{\#}:MC_{*,\ell}(\mathcal{A})\hookrightarrow MC_{*,\ell}(Q_{n})$.
-Let us take $c\in MC_{\ell,\ell}^{\circ}(\mathcal{A})\cap\ker\partial\subseteq MC_{\ell,\ell}(\mathcal{A})$.
-Since $\partial c=0$, $\phi_{\#}(c)\in MC_{\ell,\ell}(Q_{n})$ is a cycle.
-For fixed $\ell$, $MC_{\ell+1,\ell}(Q_{n})=0$, by abuse of notation, we say $\phi_{\#}(c)$ is a homology class in $MH_{\ell,\ell}(Q_{n})$.
-Recall the isomorphism given in Example 2.14
-$\nabla:\bigoplus_{\ell_{1}+\cdots+\ell_{n}=\ell}MH_{\ell_{1},\ell_{1}}(Q_{1})\otimes\cdots\otimes MH_{\ell_{n},\ell_{n}}(Q_{1})\xrightarrow{\cong}MH_{\ell,\ell}(Q_{n}),$
+By Definition 5.7, we have
+$MH_{\ell,\ell}^{\circ}(\mathcal{A})=\bigoplus_{\begin{subarray}{c}\mathcal{B}\subseteq\mathcal{A}\\ \cap\mathcal{B}=\cap\mathcal{A}\end{subarray}}MH_{\ell,\ell}(\mathcal{A};\mathcal{B}).$
+We shall prove that each summand vanishes.
+In what follows, we suppose that $\mathcal{A}$ is essential without loss of generality, because passing to the essentialization does not alter the tope graph or the intersection lattice.
+For $\mathcal{B}\subseteq\mathcal{A}$ such that $\cap\mathcal{B}=\cap\mathcal{A}=\{0\}$, let $m=\#\mathcal{B}$.
+Then $m\geq r$.
+Let $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{\mathcal{A}}=\{+,-\}^{\mathcal{A}}$ be the embedding of $\mathcal{T}(\mathcal{A})$ into the hypercube $Q_{\mathcal{A}}$.
+Since $\phi$ is distance preserving, it induces an injective chain map $\phi_{\#}:MC_{*,\ell}(\mathcal{A})\hookrightarrow MC_{*,\ell}(Q_{\mathcal{A}})$.
+Let $c\in MC_{\ell,\ell}(\mathcal{A};\mathcal{B})$ be a cycle.
+It maps to a cycle $\phi_{\#}(c)\in MC_{\ell,\ell}(Q_{\mathcal{A}})$.
+By the endpoint decomposition (2.2), we may suppose without loss of generality that $c$ is a linear combination of $\mathcal{B}$-supported walks with a common starting vertex and a common terminal vertex.
+This means that the sequence of chambers in each walk of $c$ flips exactly the $m$ coordinates corresponding to $\mathcal{B}$, while the signs on the remaining nmn-m coordinates corresponding to $\mathcal{A}\setminus\mathcal{B}$ remain equal to those of the starting vertex.
+Geometrically, this means that the entire cycle $\phi_{\#}(c)$ is confined to an $m$-dimensional face of $Q_{\mathcal{A}}$, denoted by $Q_{\mathcal{B}}$ by abuse of notation.
+We claim that this subcube $Q_{\mathcal{B}}$ contains a vertex $v_{0}$ that is not a vertex of $\mathcal{T}(\mathcal{A})$.
+Indeed, in the case $m>r$, the hyperplanes of $\mathcal{B}$ contain linear dependencies.
+Consequently, the total number of geometrically realizable sign patterns on $\mathcal{B}$ is less than $2^{m}$.
+However, the subcube $Q_{\mathcal{B}}$ contains all $2^{m}$ sign patterns.
+Therefore, at least one sign pattern in this subcube is not geometrically realizable, so at least one vertex $v_{0}$ inside the subcube $Q_{\mathcal{B}}$ does not correspond to a chamber in $\mathcal{T}(\mathcal{A})$.
+In the case $m=r$, $\mathcal{B}$ is a Boolean subarrangement.
+Its $2^{r}$ chambers form full-dimensional orthants.
+Because $\mathcal{A}$ is not Boolean, there exists a hyperplane $H\in\mathcal{A}\setminus\mathcal{B}$, and we must have $r\geq 2$.
+This hyperplane $H$ cannot intersect the interior of every orthant.
+Thus, there exists a pair of antipodal chambers $C_{v},C_{-v}\in\mathrm{Ch}(\mathcal{B})$ separated by $H$ (that is, $H$ does not cut their interiors), where $v\in\{+,-\}^{\mathcal{B}}$ is the sign vector of $C_{v}$.
+We may assume that the extension of $v$ to $\tilde{v}\in\{+,-\}^{\mathcal{A}}$ using the signs of the starting vertex of $c$, regarded as a vertex of $Q_{\mathcal{A}}$, is geometrically realizable, i.e.
+is a vertex of $\mathcal{T}(\mathcal{A})$.
+Then the extension of v-v to $\widetilde{-v}\in\{+,-\}^{\mathcal{A}}$ is not geometrically realizable, since by definition $\widetilde{-v}_{H}=\widetilde{v}_{H}$, contradicting the geometric condition that $C_{v}$ and $C_{-v}$ are separated by $H$.
+Thus, we have shown the existence of such a forbidden vertex $v_{0}$ in the subcube $Q_{\mathcal{B}}$.
+For a fixed $\ell$, $MC_{\ell+1,\ell}(Q_{\mathcal{A}})=0$; by abuse of notation, we regard $\phi_{\#}(c)$ as a homology class in $MH_{\ell,\ell}(Q_{\mathcal{A}})$.
+Recall that the isomorphism given in Example 2.14
+$\nabla:\bigoplus_{\ell_{1}+\cdots+\ell_{n}=\ell}MH_{\ell_{1},\ell_{1}}(Q_{1})\otimes\cdots\otimes MH_{\ell_{n},\ell_{n}}(Q_{1})\xrightarrow{\cong}MH_{\ell,\ell}(Q_{\mathcal{A}}),$
 is induced by the Eilenberg–Zilber chain map [17].
-We may assume there are basis elements $b^{\lambda}$ of the left hand side, one for each summand, indexed by a composition $\lambda=(\ell_{1},\ldots,\ell_{n})$ of $\ell$, such that $\phi_{\#}(c)=\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})$ for some scalars $d_{\lambda}$.
+We may assume that there are basis elements $b^{\lambda}$ of the left-hand side, one for each summand, indexed by a composition $\lambda=(\ell_{1},\ldots,\ell_{n})$ of $\ell$, such that $\phi_{\#}(c)=\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})$ for some scalars $d_{\lambda}$.
 We must show that $d_{\lambda}=0$.
-Since $c\in MC_{\ell,\ell}^{\circ}(\mathcal{A})$, it is a linear combination of walks that cross all hyperplanes in $\mathcal{A}$.
-Its image $\phi_{\#}(c)$ in $MC_{\ell,\ell}(Q_{n})$ is a linear combination of walks in $Q_{n}$ along which every coordinate gets flipped at least once.
-This means the basis elements $b^{\lambda}$ involved are indexed by compositions $\lambda=(\ell_{1},\ldots,\ell_{n})$ with all $\ell_{i}\geq 1$.
-We write such $b^{\lambda}=b_{1}^{\lambda}\otimes\cdots\otimes b_{n}^{\lambda}$, where $b_{i}^{\lambda}\in MH_{\ell_{i},\ell_{i}}(Q_{1})$ is an $\ell_{i}$-step walk in the $i$-th $Q_{1}$.
-So $b^{\lambda}$ consists of totally $\ell$ steps in these $n$ dimensions with distribution $\lambda$.
-By definition of the Eilenberg-Zilber map, $\nabla(b^{\lambda})$ is the signed sum of all possible shuffles of all these $\ell$ steps.
-It is a linear combination of walks in $Q_{n}$ from the vertex $(s(b_{1}^{\lambda}),\ldots,s(b_{n}^{\lambda}))$ to the vertex $(t(b_{1}^{\lambda}),\ldots,t(b_{n}^{\lambda}))$, where $s(b_{i}^{\lambda})$ and $t(b_{i}^{\lambda})$ represent the starting vertex and terminating vertex in $Q_{1}$ respectively.
-Since all $\ell_{i}\geq 1$, any vertex in $Q_{n}$ must be visited by at least one walk in $\nabla(b^{\lambda})$.
-The next key observation is that if $\lambda\neq\mu$, then $\nabla(b^{\lambda})$ and $\nabla(b^{\mu})$ share no common walk due to distribution difference.
-Now recall that $\phi:\mathcal{T}(\mathcal{A})\hookrightarrow Q_{n}$ is not surjective, there is a vertex $v_{0}\in Q_{n}$ that is not a vertex of $\mathcal{T}(\mathcal{A})$.
-We observed that any $\nabla(b^{\lambda})$ contains a walk passing through $v_{0}$ (see Figure 2), but $c\in MC_{\ell,\ell}(\mathcal{A})$ contains no walk passing through $v_{0}$.
-We must have $d_{\lambda}=0$ in the sum $\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})=\phi_{\#}(c)$.
-Therefore we conclude $c=0$, and $MH_{\ell,\ell}^{\circ}(\mathcal{A})=0$.
+Since $c$ crosses all hyperplanes in $\mathcal{B}$, its image $\phi_{\#}(c)$ in $MC_{\ell,\ell}(Q_{\mathcal{A}})$ is a linear combination of walks in $Q_{\mathcal{A}}$ along which every $\mathcal{B}$-coordinate is flipped at least once.
+This means that the basis elements $b^{\lambda}$ involved are indexed by compositions $\lambda=(\ell_{1},\ldots,\ell_{n})$ with all $\ell_{i}\geq 1$ for $i\in\mathcal{B}$.
+We write such an element as $b^{\lambda}=b_{1}^{\lambda}\otimes\cdots\otimes b_{n}^{\lambda}$, where $b_{i}^{\lambda}\in MH_{\ell_{i},\ell_{i}}(Q_{1})$ is an $\ell_{i}$-step walk in the $i$-th $Q_{1}$.
+Thus $b^{\lambda}$ consists of $\ell$ steps in total in these $n$ dimensions with distribution $\lambda$.
+By definition of the Eilenberg–Zilber map, $\nabla(b^{\lambda})$ is the signed sum of all possible shuffles of all these $\ell$ steps.
+It is a linear combination of walks in $Q_{\mathcal{A}}$ from the vertex $(s(b_{1}^{\lambda}),\ldots,s(b_{n}^{\lambda}))$ to the vertex $(t(b_{1}^{\lambda}),\ldots,t(b_{n}^{\lambda}))$, where $s(b_{i}^{\lambda})$ and $t(b_{i}^{\lambda})$ represent the starting vertex and terminal vertex in $Q_{1}$, respectively.
+Since all $\ell_{i}\geq 1$ for $i\in\mathcal{B}$, any vertex in the subcube $Q_{\mathcal{B}}$ must be visited by at least one walk in $\nabla(b^{\lambda})$.
+The next key observation is that if $\lambda\neq\mu$, then $\nabla(b^{\lambda})$ and $\nabla(b^{\mu})$ share no common walk because their step distributions differ.
 $G^{+}(+,+,+)$$G^{-}(-,-,-)$(+,,)(+,-,-)Start: (+,+,)(+,+,-)(,+,)(-,+,-)(,+,+)(-,+,+)(,,+)(-,-,+)End: (+,,+)(+,-,+)Step 1 ($Z_{1}$)Step 2 ($X_{1}$)Step 3 ($X_{2}$)Step 4 ($Y_{1}$)
@@
 Suppose $b^{\lambda}=(+\xrightarrow{X_{1}}-\xrightarrow{X_{2}}+)\otimes(+\xrightarrow{Y_{1}}-)\otimes(-\xrightarrow{Z_{1}}+)$, where $\lambda=(2,1,1)$.
-A 4-step shuffle walk $Z_{1}X_{1}X_{2}Y_{1}$ (orange/dotted) in $\nabla(b^{\lambda})$ is shown hitting the forbidden vertex $G^{+}$.
-Also notice that this walk cannot appear in other $\nabla(b^{\mu})$.
+A 4-step shuffle walk $Z_{1}X_{1}X_{2}Y_{1}$ (orange/dotted) in $\nabla(b^{\lambda})$ is shown passing through the forbidden vertex $G^{+}$.
+Note also that this walk cannot appear in any other $\nabla(b^{\mu})$.
+Now we apply the above observations to the forbidden vertex $v_{0}$.
+Each $\nabla(b^{\lambda})$ contains a walk passing through $v_{0}$ (see Figure 3), but $c\in MC_{\ell,\ell}(\mathcal{A})$ contains no walk passing through $v_{0}$.
+We must have $d_{\lambda}=0$ in the sum $\sum_{\lambda}d_{\lambda}\nabla(b^{\lambda})=\phi_{\#}(c)$.
+Therefore $c=0$ and $MH_{\ell,\ell}(\mathcal{A};\mathcal{B})=0$.
+∎
 Remark 5.17.
@@
 The diagonal magnitude homology $MH_{\ell,\ell}(\mathcal{A})$ is free abelian of rank $\beta_{\ell,\ell}(\mathcal{A})$, which is determined by the Boolean flats of the intersection lattice $L(\mathcal{A})$.
-Precisely, for $\ell\geq 1$,
-$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \mathrm{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\mathrm{rank}X}\binom{\ell-1}{\mathrm{rank}X-1}$.
-Consequently, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{max}-1$, where $d_{max}=\max\{\mathrm{rank}X\mid X\in L(\mathcal{A}),\mathrm{rank}X=\#\mathcal{A}_{X}\}$.
+More precisely, for $\ell\geq 1$,
+$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \operatorname{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\operatorname{rank}X}\binom{\ell-1}{\operatorname{rank}X-1}$.
+Consequently, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{\max}-1$, where
+$d_{\max}=\max\{\operatorname{rank}X\mid X\in L(\mathcal{A}),\ \operatorname{rank}X=\#\mathcal{A}_{X}\}.$
 Proof.
@@
 By Theorem 5.11, the global magnitude homology admits a face decomposition.
-By restricting this decomposition to the diagonal ($k=\ell$) and taking Betti numbers, we obtain
+Restricting this decomposition to the diagonal ($k=\ell$) and taking Betti numbers, we obtain
 $\beta_{\ell,\ell}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\beta^{\circ}_{\ell,\ell}(\mathcal{A}_{F}).$
 By Theorem 5.16, the interior Betti number $\beta^{\circ}_{\ell,\ell}(\mathcal{A}_{F})$ vanishes unless the localization $\mathcal{A}_{F}$ is Boolean, i.e.
-$\mathcal{A}_{F}=\mathrm{rank}\mathcal{A}_{F}$.
-Therefore, the sum collapses to the faces whose support flat $X=s(F)$ satisfies $\mathrm{rank}X=\#\mathcal{A}_{X}$.
-For these faces, $\mathcal{A}_{F}\cong Bl_{\mathrm{rank}X}$.
-We can group the non-zero terms of the summation by their support flats $X\in L(\mathcal{A})$.
+$\mathcal{A}_{F}=\operatorname{rank}\mathcal{A}_{F}$.
+Therefore, the sum collapses to the faces whose support flat $X=s(F)$ satisfies $\operatorname{rank}X=\#\mathcal{A}_{X}$.
+For these faces, $\mathcal{A}_{F}\cong Bl_{\operatorname{rank}X}$.
+We can group the nonzero terms of the summation by their support flats $X\in L(\mathcal{A})$.
 For each valid flat $X$, the number of faces in $\mathcal{A}$ with support $X$ is exactly $c^{X}$.
-Substituting the known formula for the interior diagonal Betti numbers of the Boolean arrangement (Corollary 4.13), $\beta^{\circ}_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$ where $d=\mathrm{rank}X$, we obtain
-$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \mathrm{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\mathrm{rank}X}\binom{\ell-1}{\operatorname{rank}X-1}$.
-For $\ell\geq\mathrm{rank}X$, the binomial $\binom{\ell-1}{\mathrm{rank}X-1}$ is a polynomial in $\ell$ of degree $\mathrm{rank}X-1$.
-Therefore for $\ell$ sufficiently large, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{max}-1$.
+Substituting the known formula for the interior diagonal Betti numbers of the Boolean arrangement (Corollary 5.15), namely $\beta^{\circ}_{\ell,\ell}(Bl_{d})=2^{d}\binom{\ell-1}{d-1}$ where $d=\operatorname{rank}X$, we obtain
+$\beta_{\ell,\ell}(\mathcal{A})=\sum_{\begin{subarray}{c}X\in L(\mathcal{A})\\ \operatorname{rank}X=\#\mathcal{A}_{X}\end{subarray}}c^{X}\cdot 2^{\operatorname{rank}X}\binom{\ell-1}{\operatorname{rank}X-1}$.
+For $\ell\geq\operatorname{rank}X$, the binomial $\binom{\ell-1}{\operatorname{rank}X-1}$ is a polynomial in $\ell$ of degree $\operatorname{rank}X-1$.
+Therefore, for $\ell$ sufficiently large, $\beta_{\ell,\ell}(\mathcal{A})$ grows as a polynomial in $\ell$ of degree $d_{\max}-1$.
 ∎
@@
 6. Open problems
-Besides the face decomposition conjecture (Conjecture 3.7, see also Corollary 4.11 and Remark 4.12), we propose the following conjectures based on observations of examples in the appendix.
+Based on observations from examples in the appendix, we propose the following conjectures.
 Conjecture 6.1.
@@
   3. (3)
-    The coefficients of the power series expression of $\mathrm{Mag}(\mathcal{A},q)$ have alternating signs (equivalently $(-1)^{\ell}\chi_{\ell}(\mathcal{A})\geq 0,\forall\ell\geq 0$).
+    $(-1)^{\ell}\chi_{\ell}(\mathcal{A})\geq 0$ for sufficiently large $\ell$.
   4. (4)
@@
   5. (5)
-    $MH_{k,\ell}(\mathcal{A})$ is torsion free.
+    $MH_{k,\ell}(\mathcal{A})$ is torsion-free.
   6. (6)
-    If $\mathcal{A}$ is not Boolean, then $\beta_{0,0}(\mathcal{A})=\beta_{r,n}(\mathcal{A})$.
-Acknowledgement
-The author thanks Weili Guo and Masahiko Yoshinaga for comments on the preliminary draft of this paper.
+    Suppose that $r\geq 2$ and that $\mathcal{A}$ is indecomposable, that is, $\mathcal{A}$ cannot be expressed as a direct sum $\mathcal{A}_{1}\oplus\mathcal{A}_{2}$ of nonzero arrangements.
+Then $\beta_{0,0}=\beta_{r,n}$.
+Acknowledgements
+The authors thank Weili Guo, Yoh Kitajima, Takuya Saito, and Masahiko Yoshinaga for their comments on a preliminary version of this paper.
 Appendix A Computer experiments
-The following examples are computed by using SageMath [28].
+The following examples are computed with SageMath [29].
 Example A.1.
-We list the results of magnitudes of simplicial arrangements $\mathcal{A}(n,k)$ in Grünbaum’s catalogue [14] for $n\leq 13$.
-Note that $\mathcal{A}(n,0),\mathcal{A}(6,1),\mathcal{A}(9,1)$ are recorded in Examples 3.10 and 3.11.
-We use the database of [6] to pull the list of normal vectors.
+We list the magnitudes of simplicial arrangements $\mathcal{A}(n,k)$ in Grünbaum’s catalogue [15] for $n\leq 13$.
+Note that $\mathcal{A}(n,0),\mathcal{A}(6,1),\mathcal{A}(9,1)$ are recorded in Examples 3.12 and 3.13.
+We use the database of [6] to obtain the list of normal vectors.
 $\mathrm{Mag}(\mathcal{A}(7,1))$ $=\frac{32q^{4}-40q^{2}+32}{(q^{4}+3q^{3}+4q^{2}+3q+1)(q^{7}+1)}$
@@
 Example A.3.
-Consider the non-simplicial arrangement $\mathcal{A}$ defined by $xyz(x+z)(y+z)$, which is also the graphic arrangement $\mathcal{A}$ of the graph $K_{4}\setminus e$.
-Note that $K_{4}\setminus e$ is chordal.
+Consider the non-simplicial arrangement $\mathcal{A}$ defined by $xyz(x+z)(y+z)$, which is also the graphic arrangement of the graph $K_{4}\setminus e$. Note that $K_{4}\setminus e$ is chordal.
 $\mathrm{Mag}(\mathcal{A})$ $=\frac{18q^{4}-2q^{3}-8q^{2}-2q+18}{(q^{4}+3q^{3}+4q^{2}+3q+1)(q^{5}+1)}$
@@
 $=96-396q+756q^{2}-924q^{3}+996q^{4}-1320q^{5}+1956q^{6}-2652q^{7}+3252q^{8}-3756q^{9}+4212q^{10}+\cdots$
-where the numerator $P(q)=96q^{8}-108q^{7}+48q^{6}+36q^{5}+36q^{3}+48q^{2}-108q+96$ and the denominator $Q(q)=q^{17}+3q^{16}+5q^{15}+7q^{14}+8q^{13}+9q^{12}+11q^{11}+13q^{10}+15q^{9}+15q^{8}+13q^{7}+11q^{6}+9q^{5}+8q^{4}+7q^{3}+5q^{2}+3q+1$.
+where the numerator is $P(q)=96q^{8}-108q^{7}+48q^{6}+36q^{5}+36q^{3}+48q^{2}-108q+96$ and the denominator is $Q(q)=q^{17}+3q^{16}+5q^{15}+7q^{14}+8q^{13}+9q^{12}+11q^{11}+13q^{10}+15q^{9}+15q^{8}+13q^{7}+11q^{6}+9q^{5}+8q^{4}+7q^{3}+5q^{2}+3q+1$.
 Example A.5.
-Consider the bracelet arrangement $\mathcal{A}$ defined by $x_{1}x_{2}x_{3}(x_{1}+x_{4})(x_{2}+x_{4})(x_{3}+x_{4})(x_{1}+x_{2}+x_{4})(x_{1}+x_{3}+x_{4})(x_{2}+x_{3}+x_{4})$ (see Example 1.4 of [7]),
+Consider the bracelet arrangement $\mathcal{A}$ defined by $x_{1}x_{2}x_{3}(x_{1}+x_{4})(x_{2}+x_{4})(x_{3}+x_{4})(x_{1}+x_{2}+x_{4})(x_{1}+x_{3}+x_{4})(x_{2}+x_{3}+x_{4})$ (see Example 1.4 of [7]).
 $\mathrm{Mag}(\mathcal{A})$ $=\frac{6(17q^{8}-4q^{7}-8q^{6}-9q^{5}+44q^{4}-9q^{3}-8q^{2}-4q+17)}{\Phi_{2}(q)^{3}\Phi_{3}(q)^{2}\Phi_{9}(q)\Phi_{10}(q)}$
 $=102-432q+864q^{2}-1176q^{3}+1584q^{4}-2628q^{5}+4344q^{6}-6048q^{7}+7200q^{8}-8130q^{9}+9684q^{10}+\cdots$
-where the numerator $P(q)=102q^{8}-24q^{7}-48q^{6}-54q^{5}+264q^{4}-54q^{3}-48q^{2}-24q+102$ and the denominator $Q(q)=q^{17}+4q^{16}+8q^{15}+11q^{14}+12q^{13}+13q^{12}+16q^{11}+20q^{10}+23q^{9}+23q^{8}+20q^{7}+16q^{6}+13q^{5}+12q^{4}+11q^{3}+8q^{2}+4q+1$.
+where the numerator is $P(q)=102q^{8}-24q^{7}-48q^{6}-54q^{5}+264q^{4}-54q^{3}-48q^{2}-24q+102$ and the denominator is $Q(q)=q^{17}+4q^{16}+8q^{15}+11q^{14}+12q^{13}+13q^{12}+16q^{11}+20q^{10}+23q^{9}+23q^{8}+20q^{7}+16q^{6}+13q^{5}+12q^{4}+11q^{3}+8q^{2}+4q+1$.
 Example A.6.
-We collect here some results of magnitude Betti numbers using SageMath codes of [16].
+We collect some results here on magnitude Betti numbers computed with the SageMath code of [17].
 […]
@@
 8 0 0 0 0 0 0 0 0 36
-Table 1. Magnitude Betti numbers $\beta_{k,l}$ for the Boolean arrangement $Bl_{2}=U_{2,2}$ defined by $xy=0$.
+Table 1. Magnitude Betti numbers $\beta_{k,\ell}$ for the Boolean arrangement $Bl_{2}=U_{2,2}$ defined by $xy=0$.
 […]
@@
 8 0 0 0 0 0 0 0 0 12
-Table 2. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{2}=U_{2,3}$.
+Table 2. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $A_{2}=U_{2,3}$.
 […]
@@
 8 0 0 0 0 0 0 0 0 16
-Table 3. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $B_{2}=U_{2,4}$.
+Table 3. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $B_{2}=U_{2,4}$.
 […]
@@
 8 0 0 0 0 0 0 0 0 240
-Table 4. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{3}$.
+Table 4. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $A_{3}$.
 […]
@@
 6 0 0 0 0 0 0 2280
-Table 5. Magnitude Betti numbers $\beta_{k,l}$ for the Coxeter arrangement $A_{4}$.
+Table 5. Magnitude Betti numbers $\beta_{k,\ell}$ for the Coxeter arrangement $A_{4}$.
 […]
@@
 8 0 0 0 0 0 0 0 0 384
-Table 6. Magnitude Betti numbers $\beta_{k,l}$ for the generic arrangement $U_{3,4}$ defined by $xyz(x+y+z)$ Example A.2.
+Table 6. Magnitude Betti numbers $\beta_{k,\ell}$ for the generic arrangement $U_{3,4}$ defined by $xyz(x+y+z)$; see Example A.2.
 […]
@@
 6 0 0 0 0 0 0 2940
-Table 7. Magnitude Betti numbers $\beta_{k,l}$ for the generic arrangement $U_{4,5}$.
+Table 7. Magnitude Betti numbers $\beta_{k,\ell}$ for the generic arrangement $U_{4,5}$.
 […]
@@
 8 0 0 0 0 0 0 0 0 280
-Table 8. Magnitude Betti numbers $\beta_{k,l}$ for the graphic arrangement of $K_{4}\setminus e$ Example A.3
+Table 8. Magnitude Betti numbers $\beta_{k,\ell}$ for the graphic arrangement of $K_{4}\setminus e$; see Example A.3.
 […]
@@
 8 0 0 0 0 0 0 0 0 264
-Table 9. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(7,1)$
+Table 9. Magnitude Betti numbers $\beta_{k,\ell}$ for $\mathcal{A}(7,1)$
 […]
@@
 8 0 0 0 0 0 0 0 0 344
-Table 10. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(8,1)$
+Table 10. Magnitude Betti numbers $\beta_{k,\ell}$ for $\mathcal{A}(8,1)$
 […]
@@
 8 0 0 0 0 0 0 0 0 480
-Table 11. Magnitude Betti numbers $\beta_{k,l}$ for $\mathcal{A}(9,1)$
-Reference
-  • [1] T.
-Abe, T.
-Horiguchi, M.
-Masuda, S.
-Murai, and T.
-Sato (2020) Hessenberg varieties and hyperplane arrangements.
-J.
-Reine Angew.
-Math.
-764, pp.
-241–286.
-External Links: ISSN 0075-4102,1435-5345, Document, Link, MathReview (Marko Radovanović) Cited by: §2.7.
-  • [2] M.
-Aguiar and S.
-Mahajan (2017) Topics in hyperplane arrangements.
-Mathematical Surveys and Monographs, Vol.
-226, American Mathematical Society, Providence, RI.
-External Links: ISBN 978-1-4704-3711-4, Document, Link, MathReview (Tom C.
-Braden) Cited by: §2.6, Proposition 2.2, Remark 2.25, Theorem 2.33.
-  • [3] A.
-Björner, P.
-H.
-Edelman, and G.
-M.
-Ziegler (1990) Hyperplane arrangements with a lattice of regions.
-Discrete Comput.
-Geom.
-5 (3), pp.
-263–288.
-External Links: ISSN 0179-5376,1432-0444, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.7, Proposition 2.36.
-  • [4] A.
-Björner, M.
-Las Vergnas, B.
-Sturmfels, N.
-White, and G.
-M.
-Ziegler (1999) Oriented matroids.
-Second edition, Encyclopedia of Mathematics and its Applications, Vol.
-46, Cambridge University Press, Cambridge.
-External Links: ISBN 0-521-77750-X, Document, Link, MathReview Entry Cited by: Theorem 1.1, Proposition 2.30, Proposition 2.34, Corollary 2.35.
-  • [5] S.
-Brenner, J.
-Cardinal, T.
-McConville, A.
-Merino, and T.
-Mütze (2026) Combinatorial generation via permutation languages.
-VII.
-Supersolvable hyperplane arrangements.
-European J.
-Combin.
-135, pp.
-Paper No.
-104367.
-External Links: ISSN 0195-6698,1095-9971, Document, Link, MathReview Entry Cited by: §2.5.
-  • [6] M.
-Cuntz, S.
-Elia, and J.
-Labbé (2022) Congruence normality of simplicial hyperplane arrangements via oriented matroids.
-Ann.
-Comb.
-26 (1), pp.
-1–85.
-External Links: ISSN 0218-0006,0219-3094, Document, Link, MathReview (Hery Randriamaro) Cited by: Example A.1.
-  • [7] G. Denham, G. Dorpalen-Barry, and N. Proudfoot (2026) Cleanliness and the varchenko-gelfand algebra. arXiv preprint arXiv:2512.10077v2. Cited by: Example A.5.
-  • [8] G.
-Denham and P.
-Hanlon (1999) Some algebraic properties of the Schechtman-Varchenko bilinear forms.
-In New perspectives in algebraic combinatorics (Berkeley, CA, 1996–97), Math.
-Sci.
-Res.
-Inst.
-Publ., Vol.
-38, pp.
-149–176.
-External Links: ISBN 0-521-77087-4, MathReview (P.
-Orlik) Cited by: §2.6.
-  • [9] R.
-Diestel ([2025] ©2025) Graph theory.
-Sixth edition, Graduate Texts in Mathematics, Vol.
-173, Springer, Berlin.
-External Links: ISBN 978-3-662-70106-5; 978-3-662-70107-2, MathReview Entry Cited by: §2.5.
-  • [10] P.
-H.
-Edelman and J.
-W.
-Walker (1985) The homotopy type of hyperplane posets.
-Proc.
-Amer.
-Math.
-Soc.
-94 (2), pp.
-221–225.
-External Links: ISSN 0002-9939,1088-6826, Document, Link, MathReview (Udo Pachner) Cited by: Proposition 2.34.
-  • [11] P.
-H.
-Edelman (1984) A partial order on the regions of ${\bf R}^{n}$ dissected by hyperplanes.
-Trans.
-Amer.
-Math.
-Soc.
-283 (2), pp.
-617–631.
-External Links: ISSN 0002-9947,1088-6850, Document, Link, MathReview (Robert A.
-Proctor) Cited by: Corollary 2.35.
-  • [12] L. Giordani, T. Möller, P. Mücksch, and G. Roehrle (2025) On connected subgraph arrangements. arXiv preprint arXiv:2502.18144v1. Cited by: §2.7.
-  • [13] K. Gomi (2025) Magnitude homology of geodesic space. arXiv preprint arXiv:1902.07044v3. Cited by: Proposition 2.21.
-  • [14] B.
-Grünbaum (2009) A catalogue of simplicial arrangements in the real projective plane.
-Ars Math.
-Contemp.
-2 (1), pp.
-1–25.
-External Links: ISSN 1855-3966,1855-3974, Document, Link, MathReview Entry Cited by: Example A.1, §1, §3.3.
-  • [15] Y. Gu (2018) Graph magnitude homology via algebraic morse theory. arXiv preprint arXiv:1809.07240v1. Cited by: Example 2.16.
-  • [16] R.
-Hepworth and S.
-Willerton (2017) Categorifying the magnitude of a graph.
-Homology Homotopy Appl.
-19 (2), pp.
-31–60.
-External Links: ISSN 1532-0073,1532-0081, Document, Link, MathReview (Haimiao Chen) Cited by: Example A.6, §1, §2.3, Proposition 2.10, Theorem 2.12, Theorem 2.12, Proposition 2.8, Proposition 2.9, §4.5.
-  • [17] W.
-Hochstättler and V.
-Welker (2019) The Varchenko determinant for oriented matroids.
-Math.
-Z.
-293 (3-4), pp.
-1415–1430.
-External Links: ISSN 0025-5874,1432-1823, Document, Link, MathReview (Daryl Funk) Cited by: §2.6.
-  • [18] M.
-Jambu and L.
-Paris (1995) Combinatorics of inductively factored arrangements.
-European J.
-Combin.
-16 (3), pp.
-267–292.
-External Links: ISSN 0195-6698,1095-9971, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.7.
-  • [19] R.
-Kaneta and M.
-Yoshinaga (2021) Magnitude homology of metric spaces and order complexes.
-Bull.
-Lond.
-Math.
-Soc.
-53 (3), pp.
-893–905.
-External Links: ISSN 0024-6093,1469-2120, Document, Link, MathReview (David Matthew Freeman) Cited by: §2.3, §2.3, Theorem 2.20, Theorem 2.23.
-  • [20] V. Körber, T. Schnieders, J. Stricker, and J. Walizadeh (2025) Hamiltonian cycles in simplicial and supersolvable hyperplane arrangements. arXiv preprint arXiv:2508.14538v2. Cited by: §2.5.
-  • [21] T.
-Leinster and M.
-W.
-Meckes (2017) The magnitude of a metric space: from category theory to geometric measure theory.
-In Measure theory in non-smooth spaces, Partial Differ.
-Equ.
-Meas.
-Theory, pp.
-156–193.
-External Links: ISBN 978-3-11-055082-5; 978-3-11-055083-2, MathReview (Tobias W.
-Fritz) Cited by: §1.
-  • [22] T.
-Leinster and M.
-Meckes (2023) Spaces of extremal magnitude.
-Proc.
-Amer.
-Math.
-Soc.
-151 (9), pp.
-3967–3973.
-External Links: ISSN 0002-9939,1088-6826, Document, Link, MathReview (Shlomo Reisner) Cited by: §3.1.
-  • [23] T.
-Leinster and M.
-Shulman (2021) Magnitude homology of enriched categories and metric spaces.
-Algebr.
-Geom.
-Topol.
-21 (5), pp.
-2175–2221.
-External Links: ISSN 1472-2747,1472-2739, Document, Link, MathReview (Ahmet A.
-Husainov) Cited by: §1.
-  • [24] T. Leinster (2013) The magnitude of metric spaces. Doc. Math. 18, pp. 857–905. External Links: ISSN 1431-0635,1431-0643, MathReview (J. Böhm) Cited by: §1.
-  • [25] T.
-Leinster (2019) The magnitude of a graph.
-Math.
-Proc.
-Cambridge Philos.
-Soc.
-166 (2), pp.
-247–264.
-External Links: ISSN 0305-0041,1469-8064, Document, Link, MathReview (Lorenzo Traldi) Cited by: §1, §2.2, Proposition 2.3, Proposition 2.5, Theorem 2.6, Proposition 2.7.
-  • [26] S.
-Oh, A.
-Postnikov, and H.
-Yoo (2008) Bruhat order, smooth Schubert varieties, and hyperplane arrangements.
-J.
-Combin.
-Theory Ser.
-A 115 (7), pp.
-1156–1166.
-External Links: ISSN 0097-3165,1096-0899, Document, Link, MathReview (David E.
-Anderson) Cited by: §2.7.
-  • [27] P.
-Orlik and H.
-Terao (1992) Arrangements of hyperplanes.
-Grundlehren der mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences], Vol.
-300, Springer-Verlag, Berlin.
-External Links: ISBN 3-540-55259-6, Document, Link, MathReview (Michel Yves Jambu) Cited by: §2.4.
-  • [28] The Sage Developers (2025) Sagemath, the Sage Mathematics Software System (Version 10.8). Note: https://www.sagemath.org Cited by: Appendix A.
-  • [29] L. Solomon (1966) The orders of the finite Chevalley groups. J. Algebra 3, pp. 376–393. External Links: ISSN 0021-8693, Document, Link, MathReview (T. Ono) Cited by: Proposition 2.37.
-  • [30] R.
-P.
-Stanley (2007) An introduction to hyperplane arrangements.
-In Geometric combinatorics, IAS/Park City Math.
-Ser., Vol.
-13, pp.
-389–496.
-External Links: ISBN 978-0-8218-3736-8; 0-8218-3736-2, Document, Link, MathReview Entry Cited by: §2.4, Remark 2.25.
-  • [31] A.
-Varchenko (1993) Bilinear form of real configuration of hyperplanes.
-Adv.
-Math.
-97 (1), pp.
-110–144.
-External Links: ISSN 0001-8708,1090-2082, Document, Link, MathReview (Hiroaki Terao) Cited by: §2.6, Theorem 2.33.
-  • [32] Y. Yagi and M. Yoshinaga (2026) Reconstrucion of oriented matroids from varchenko-gelfand algebras. arXiv preprint arXiv:2509.19905v3. Cited by: §1.
-  • [33] T.
-Zaslavsky (1975) Facing up to arrangements: face-count formulas for partitions of space by hyperplanes.
-Mem.
-Amer.
-Math.
-Soc.
-1, pp.
-vii+102.
-External Links: ISSN 0065-9266,1947-6221, Document, Link, MathReview (E.
-Jucovič) Cited by: §2.4, Remark 2.25.
+Table 11. Magnitude Betti numbers $\beta_{k,\ell}$ for $\mathcal{A}(9,1)$